### v1

# Dimension-independent convergence rates of randomized nets using median-of-means

Abstract
Recent advances in quasi-Monte Carlo integration demonstrate that the median of linearly scrambled digital net estimators achieves near-optimal convergence rates for high-dimensional integrals without requiring a priori knowledge of the integrand’s smoothness. Building on this framework, we prove that the median estimator attains dimension-independent convergence under tractability conditions characterized by low effective dimensionality, a property known as strong tractability in complexity theory. Our analysis strengthens existing guarantees by improving the convergence rates and relaxing the theoretical assumptions previously required for strong tractability.
1 Introduction
The quasi-Monte Carlo (QMC) method is a powerful alternative to conventional Monte Carlo (MC) techniques for high-dimensional integration. While both methods approximate integrals by averaging $n$ function evaluations, QMC replaces MC’s pseudorandom sampling with low-discrepancy point sets engineered to uniformly explore the sampling space. This structured approach allows QMC to avoid the exponential cost growth with dimension of classical quadrature rules while achieving faster error convergence than MC. These advantages have made QMC particularly valuable in fields like computational finance [12, 15] and physics [13], where high-dimensional integrals are pervasive.
Building on these strengths, we focus on randomized quasi-Monte Carlo (RQMC) methods based on linearly scrambled base-2 digital nets. Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [5]. The random linear scrambling [18], a computationally efficient alternative to Owen’s scrambling [21], preserves the structure of digital nets while introducing controlled randomness. This randomization enhances the root mean squared error convergence rate to nearly $O(n^{-3/2})$ under smoothness conditions and enables statistical inference. However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [23].
To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [2, 6, 7, 8, 28]. The median of linearly scrambled digital net estimators was first analyzed in [23], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$. Subsequent work in [24] extended this result to higher dimensions. For finitely differentiable integrands, [22] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
A critical gap in these prior studies is the role of dimensionality. We focus on the regime where the dimension grows indefinitely while the variance remains bounded—a common scenario in path integral approximations [14]. Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [20], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3). This notion of relative variation aligns with the effective dimension framework introduced in [1], which quantifies how high-dimensional integrands can behave as if they reside in lower-dimensional subspaces. The connection to weighted function spaces is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [8]. A key distinction lies in the randomization: their framework omits digital shifts (Subsection 2.1), a critical component of our scrambling method, and employs worst-case integration error as the primary metric. In contrast, we demonstrate that median estimators incorporating digital shifts achieve near-optimal convergence rates with high probability, avoiding worst-case error analysis entirely. This probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases. Within this framework, we improve the convergence rates from nearly $O(n^{-\alpha})$ in [8] to nearly $O(n^{-\alpha-1/2})$, where $\alpha$ denotes the smoothness parameter in their analysis (we note that $\alpha$ is defined differently in this paper). Furthermore, we relax the strong tractability conditions, as detailed in Remarks 3 and 7. Specifically, we show digital nets under the completely random design (Subsection 2.1) achieve strong tractability under conditions similar to those for polynomial lattices in [8], while digital nets under the random linear scrambling attain the same property under strictly weaker conditions than those assumed in [8].
The remainder of this paper is structured as follows. Section 2 reviews essential background on digital nets and establishes notation. Subsection 2.3 in particular details how the median trick enables probabilistic error guarantees and accelerates convergence rates. Section 3 introduces ANOVA decomposition and formally defines the relative variation $\gamma_{u}$, a key measure of effective dimensionality. Section 4 provides a foundational analysis of the median estimator’s error bounds and tractability under the completely random design, a simplified randomization scheme that isolates core theoretical insights. Section 5 generalizes these results to a broader class of randomization schemes, with the random linear scrambling serving as a key example. Section 6 empirically validates our theoretical findings through simulations on high-dimensional integrands. Finally, Section 7 concludes the paper with a discussion of future research directions.
2 Background and notation
Let =$\mathbb{N}=\{1,2,3,\dots\}$ denote the natural numbers and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $s$-dimensional indices, we define $\mathbb{N}^{s}_{*}=\mathbb{N}_{0}^{s}\setminus\{\boldsymbol{0}\}$ to exclude the zero vector. We define =$\mathbb{Z}_{\leqslant\ell}=\{0,1,\dots,\ell\}$ for $\ell\in\mathbb{N}_{0}$ and <=$\mathbb{Z}_{<\ell}=\{0,1,\dots,\ell-1\}$ for $\ell\in\mathbb{N}$. The dimension of the integration domain is denoted by $s$, and we write $1{:}s=\{1,2,\dots,s\}$ for $s\in\mathbb{N}$. For a vector $\boldsymbol{x}\in{}^{s}$ and a subset $u\subseteq 1{:}s$, $\boldsymbol{x}_{u}$ denotes the subvector of $\boldsymbol{x}$ indexed by $u$, while $\boldsymbol{x}_{u^{c}}$ denotes the subvector indexed by $1{:}s\setminus u$. The indicator function $\boldsymbol{1}\{\mathcal{A}\}$ equals $1$ if event $\mathcal{A}$ occurs and 00 otherwise. For $\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{s})\in\mathbb{N}^{s}_{0}$, $\|\boldsymbol{\ell}\|_{1}=\sum_{j=1}^{s}\ell_{j}$. The cardinality of a set $K$ is $|K|$. For positive sequence $a_{n}$ and $b_{n}$, $a_{n}=O(b_{n})$ denotes asymptotic dominance: $a_{n}\leqslant Cb_{n}$ for some constant $C<\infty$ and all sufficiently large $n\in\mathbb{N}$.
We assume the integrand $f:[0,1]^{s}\to\mathbb{R}$ is continuous, denoted as $f\in C([0,1]^{s})$, with $\|f\|_{\infty}=\max_{\boldsymbol{x}\in[0,1]^{s}}|f(\boldsymbol{x})|$. The $L^{p}$-norm of $f$ is $\|f\|_{L^{p}}=\big{(}\int_{[0,1]^{s}}|f(\boldsymbol{x})|^{p}\,\mathrm{d}\boldsymbol{x}\big{)}^{1/p}$, and $L^{p}([0,1]^{s})$ denotes the space of functions with finite $L^{p}$-norm.
Quasi-Monte Carlo (QMC) methods approximate
$\mu=\int_{[0,1]^{s}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}\quad\text{by}\quad\hat{\mu}=\frac{1}{n}\sum_{i=0}^{n-1}f(\boldsymbol{x}_{i})$
for points $\boldsymbol{x}_{i}\in[0,1]^{s}$. In this paper, the points $\boldsymbol{x}_{i}$ are constructed by base-2 digital nets defined in the next subsection.
2.1 Digital net construction and randomization
For $m\in\mathbb{N}$ and $i\in\mathbb{Z}_{2^{m}}$, we express $i$ in its binary form $i=\sum_{\ell=1}^{m}i_{\ell}2^{\ell-1}$, encoded as the column vector $\vec{i}=(i_{1},\dots,i_{m})^{\mathsf{T}}\in\{0,1\}^{m}$. Analogously, for $a\in[0,1)$, we approximate its infinite binary expansion $a=\sum_{\ell=1}^{\infty}a_{\ell}2^{-\ell}$ by truncating to $E\in\mathbb{N}$ digits, yielding $\vec{a}=(a_{1},\dots,a_{E})^{\mathsf{T}}\in\{0,1\}^{E}$. For numbers with dual representations (such as dyadic rationals), we select the finite expansion ending in zeros.
A base-2 digital net in $[0,1]^{s}$ is determined by $s$ matrices $C_{j}\in\{0,1\}^{E\times m}$. The deterministic points $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{is})$ are constructed via:
$\vec{x}_{ij}=C_{j}\vec{i}\ \ \mathsf{mod}\ 2\text{ for }i\in\mathbb{Z}_{2^{m}},j\in 1{:}s,$
where $\vec{x}_{ij}\in\{0,1\}^{E}$ maps to $x_{ij}\in[0,1)$by appending zeros to the truncated $E$-digit representation. Classical constructions typically assume $E\leqslant m$.
To introduce randomness, we modify the construction as:
[…]
where $C_{j}\in\{0,1\}^{E\times m}$ and $\vec{D}_{j}\in\{0,1\}^{E}$ are random with precision $E\geqslant m$. The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli$(0.5)$ entries. A common way to randomize $C_{j}$ is the random linear scrambling [18]:
[…]
where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli$(0.5)$), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [5, Chapter 4.4] for details).
While effective, the reliance of random linear scrambling on carefully designed generating matrices complicates theoretical analysis. We address this in Section 5. For initial insights, Section 4 adopts the complete random design [22], where all entries of $C_{j}$ are drawn independently from Bernoulli$(0.5)$. This approach retains the asymptotic convergence rate of random linear scrambling while circumventing the need for pre-optimized generator matrices.
Let $\boldsymbol{x}_{i}[E]$ denote points generated by equation (1) with $E$-bit precision. The QMC estimator becomes:
$\hat{\mu}_{E}=\frac{1}{n}\sum_{i=0}^{n-1}f(\boldsymbol{x}_{i})\quad\text{for $\boldsymbol{x}_{i}=\boldsymbol{x}_{i}[E]$}.$ (2)
We conveniently assume $E=\infty$ and focus our theoretical analysis on $\hat{\mu}_{\infty}$. In practice, $E$ is constrained by computational precision. Lemma 1 of [24] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$. Thus, $\hat{\mu}_{E}$ closely approximates $\hat{\mu}_{\infty}$ when $\omega_{f}(\sqrt{s}2^{-E})\ll|\hat{\mu}_{E}-\mu|$.
2.2 Walsh functions and error analysis for base-2 digital nets
Walsh functions form the natural orthonormal basis for analyzing base-2 digital nets. For $k\in\mathbb{N}_{0}$ and $x\in[0,1)$, the univariate Walsh function $\mathrm{wal}_{k}(x)$ is defined as
$\mathrm{wal}_{k}(x)=(-1)^{\vec{k}^{\mathsf{T}}\vec{x}},$
where $\vec{k}\in\{0,1\}^{\infty}$ and $\vec{x}\in\{0,1\}^{\infty}$ are the binary expansions of $k$ and $x$, respectively. Since $\vec{k}$ has finitely many nonzero digits, finite-precision truncations of $\vec{x}$ suffice for computation.
For $s$-dimensional functions, the multivariate Walsh function $\mathrm{wal}_{\boldsymbol{k}}:[0,1)^{s}\to\{-1,1\}$ is the tensor product of univariate Walsh functions:
$\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})=\prod_{j=1}^{s}\mathrm{wal}_{k_{j}}(x_{j})=(-1)^{\sum_{j=1}^{s}\vec{k}_{j}^{\mathsf{T}}\vec{x}_{j}},$
[…]
$f(\boldsymbol{x})=\sum_{\boldsymbol{k}\in\mathbb{N}_{0}^{s}}\hat{f}(\boldsymbol{k})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x}),\quad\text{where}\quad\hat{f}(\boldsymbol{k})=\int_{[0,1]^{s}}f(\boldsymbol{x})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}.$ (3)
The equality holds in the $L^{2}$-sense. Using this decomposition, [24] establishes a key error structure for QMC estimators:
Lemma 1.
For $f\in C([0,1]^{s})$, the error of $\hat{\mu}_{\infty}$ defined by equation (2) satisfies
$\hat{\mu}_{\infty}-\mu=\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}}Z(\boldsymbol{k})S(\boldsymbol{k})\hat{f}(\boldsymbol{k}),$ (4)
where
$Z(\boldsymbol{k})=\boldsymbol{1}\Big{\{}\sum_{j=1}^{s}\vec{k}_{j}^{\mathsf{T}}C_{j}=\boldsymbol{0}\ \mathsf{mod}\ 2\Big{\}}\quad\text{and}\quad S(\boldsymbol{k})=(-1)^{\sum_{j=1}^{s}\vec{k}_{j}^{\mathsf{T}}\vec{D}_{j}}.$
Here, $Z(\boldsymbol{k})$ is an indicator for failure to integrate the Walsh mode $\mathrm{wal}_{\boldsymbol{k}}$ exactly, and $S(\boldsymbol{k})$ encodes the random sign induced by the digital shifts $\vec{D}_{j}$.
2.3 Probabilistic error guarantee of the median
The analysis of equation (4) uses two key properties of the coefficients $S(\boldsymbol{k})$: they are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [23, Lemma 4] for the univariate case $s=1$). These properties allow us to compute the variance of $\hat{\mu}_{\infty}$ and construct probabilistic error bounds by Chebyshev’s inequality.
To sharpen these bounds, we introduce a low-probability exclusion event $\mathcal{A}$ satisfying $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and $\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\ll\mathrm{Var}(\hat{\mu}_{\infty})$. This leads to the following median concentration result:
Lemma 2.
Let $\mathcal{A}$ be an event with $\Pr(\mathcal{A})\leqslant\delta<1/8$ and $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$. Then the sample median $\hat{\mu}^{(r)}_{\infty}$ of $12r-$ 1 independently generated replicates of $\hat{\mu}_{\infty}$ satisfies
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\delta^{-1}\Pr(\mathcal{A}^{c})\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\Big{)}\leqslant(8\delta)^{r}.$
Proof.
This is an intermediate result proven in [24, Lemma 6]. ∎
For $f\in C([0,1]^{s})$, the above lemma implies the mean squared error bound
$\mathbb{E}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}\leqslant\delta^{-1}\Pr(\mathcal{A}^{c})\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})+(8\delta)^{r}\|f\|^{2}_{\infty}.$
Because the second term decays exponentially in $r$, it can be made negligible relative to the first term by choosing $r$ sufficiently large.
We choose $\mathcal{A}$ to be the event $\{Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K\}$ for a carefully chosen index set $K\subseteq\mathbb{N}^{s}_{*}$. It is shown in [22, Subsection 2.4] that
${\Pr(\mathcal{A}^{c})}\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\leqslant\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}.$ (5)
The central challenge is to select $K$ such that this sum converges at the desired rate. Our analysis addresses this by constructing $K$ adaptively based on the decay rate of the Walsh coefficients $\hat{f}(\boldsymbol{k})$, which encodes the integrand’s smoothness and effective dimensionality.
2.4 Walsh coefficients and function smoothness
To analyze the decay of Walsh coefficients, we introduce the following notation. For $k=\sum_{\ell=1}^{\infty}a_{\ell}2^{\ell-1}$, let $\kappa=\{\ell\in\mathbb{N}\mid a_{\ell}=1\}$ denote the positions of nonzero bits in the binary expansion of $k$. We use $k$ and $\kappa$ interchangeably because each uniquely determines the other. In this framework, $|\kappa|$ equals the number of nonzero bits of $k$, and the Walsh function simplifies to:
${\mathrm{wal}_{\kappa}}(x)=\prod_{\ell\in\kappa}(-1)^{\vec{x}(\ell)},$
where $\vec{x}(\ell)$ denotes the $\ell$-th digit of $x$ in its binary expansion.
For a finite subset $\kappa\subseteq\mathbb{N}$, we define
  • $\lceil\kappa\rceil_{q}$: The $q$-th largest element of $\kappa$ (or 00 if $|\kappa|<q$).
  • $\lceil\kappa\rceil_{1{:}q}$: The set of the $q$ largest elements of $\kappa$ (or $\varnothing$ if $q=0$).
[…]
$f^{(\boldsymbol{\alpha})}=f^{(\alpha_{1},\dots,\alpha_{s})}=\frac{\partial^{\|\boldsymbol{\alpha}\|_{1}}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{s}^{\alpha_{s}}},$
with the convention $f^{(\boldsymbol{0})}=f$.
Lemma 3.
Let $\boldsymbol{\alpha}\in\mathbb{N}^{s}_{0}$ and $f^{(\boldsymbol{\alpha})}\in C([0,1]^{s})$. If $|\kappa_{j}|\geqslant\alpha_{j}$ for all $j\in 1{:}s$, then
$\hat{f}(\boldsymbol{k})=(-1)^{\|\boldsymbol{\alpha}\|_{1}}\int_{[0,1]^{s}}f^{(\boldsymbol{\alpha})}(\boldsymbol{x})\prod_{j=1}^{s}{\mathrm{wal}_{\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}\alpha_{j}}}}(x_{j})W_{\lceil\kappa_{j}\rceil_{1{:}\alpha_{j}}}(x_{j})\,\mathrm{d}\boldsymbol{x},$ (6)
where $W_{\kappa}:[0,1]\to\mathbb{R}$ for $\kappa\subseteq\mathbb{N}$ is defined recursively by $W_{\varnothing}(x)=1$ and
$W_{\kappa}(x)=\int_{[0,1]}(-1)^{\vec{x}(\lfloor\kappa\rfloor)}W_{\kappa\setminus\lfloor\kappa\rfloor}(x)\,\mathrm{d}x.$
with $\lfloor\kappa\rfloor$ denoting the smallest element of $\kappa$. Furthermore, for nonempty $\kappa$, $W_{\kappa}(x)$ is continuous, nonnegative, periodic with period $2^{-\lfloor\kappa\rfloor+1}$ and
$\int_{[0,1]}W_{\kappa}(x)\,\mathrm{d}x=\prod_{\ell\in\kappa}2^{-\ell-1}\quad\text{and}\quad\max_{x\in[0,1]}W_{\kappa}(x)=2\prod_{\ell\in\kappa}2^{-\ell-1}.$ (7)
Proof.
Equation (6) comes from [26, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [26, Section 3]. ∎
For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $[0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$. In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3. This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (6), as discussed in [22, Remark 1].
3 ANOVA decomposition and effective dimensionality
To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [16]:
$f=\sum_{u\subseteq 1{:}s}f_{u},$
where each component $f_{u}$ depends only on $\boldsymbol{x}_{u}$ and satisfies
$\int_{0}^{1}f_{u}(\boldsymbol{x})\,\mathrm{d}x_{j}=0\quad\forall j\in u.$
This decomposition is unique in the $L^{2}$-sense, with $f_{\varnothing}=\mu$. For $f\in C([0,1]^{s})$, we enforce $f_{u}\in C([0,1]^{s})$ via the recursive definition:
$f_{u}(\boldsymbol{x})=\int_{[0,1]^{s-|u|}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}_{u^{c}}-\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq u\end{subarray}}f_{v}(\boldsymbol{x}).$ (8)
This construction preserves smoothness: $f_{u}\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ for all $u\subseteq 1{:}s$ if $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$. Here, $f_{u}$ captures the variation due to interactions among variables $\boldsymbol{x}_{u}$. By an abuse of notation, we write $f_{u}(\boldsymbol{x})=f_{u}(\boldsymbol{x}_{u})$, where the latter $f_{u}:[0,1]^{|u|}\to\mathbb{R}$ is a $|u|$-dimensional function.
Let $\boldsymbol{s}(\boldsymbol{k})=\{j\in 1{:}s\mid k_{j}\neq 0\}$ denote the support of $\boldsymbol{k}=(k_{1},\dots,k_{s})$. From the Walsh decomposition (3),
$f_{u}(\boldsymbol{x})=\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{N}^{s}_{0}\\ \boldsymbol{s}(\boldsymbol{k})=u\end{subarray}}\hat{f}(\boldsymbol{k})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x}),$ (9)
which follows from the fact that $\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})$ depends only on $\{x_{j}\mid j\in\boldsymbol{s}(\boldsymbol{k})\}$ and
$\int_{0}^{1}\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})\,\mathrm{d}x_{j}=\Big{(}\int_{0}^{1}{\mathrm{wal}_{\kappa_{j}}}(x_{j})\,\mathrm{d}x_{j}\Big{)}\prod_{j^{\prime}\in\boldsymbol{s}(\boldsymbol{k})\setminus\{j\}}{\mathrm{wal}_{\kappa_{j^{\prime}}}}(x_{j^{\prime}})=0$
if $j\in\boldsymbol{s}(\boldsymbol{k})$. Thus, $\hat{f}(\boldsymbol{k})=\hat{f}_{u}(\boldsymbol{k})$ if $\boldsymbol{s}(\boldsymbol{k})=u$.
To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [3]. For a function $f:[0,1]^{d}\to\mathbb{R}$ and a subcube $J=\prod_{j=1}^{d}[a_{j},b_{j})$, we define the alternating sum over $J$:
$\Delta(f,J)=\sum_{v\subseteq 1{:}d}(-1)^{d-|v|}f\Big{(}(\boldsymbol{a}_{v},\boldsymbol{b}_{v^{c}})\Big{)},$
where $(\boldsymbol{a}_{v},\boldsymbol{b}_{v^{c}})$ denotes the vertex of $J$ with $x_{j}=a_{j}$ if $j\in v$ and $x_{j}=b_{j}$ otherwise. For $\lambda\in(0,1]$, the fractional Vitali variation of order $\lambda$ is:
$V^{(d)}_{\lambda}(f)=\sup_{\mathcal{P}}\Bigl{(}\sum_{J\in\mathcal{P}}\mathrm{Vol}(J)\Bigl{|}\frac{\Delta(f,J)}{\mathrm{Vol}(J)^{\lambda}}\Bigr{|}^{2}\Bigr{)}^{1/2}$ (10)
where the supremum is over all partitions $\mathcal{P}$ of $[0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$. This generalizes Sobolev norms: for $f$ with weak mixed derivative $f^{(1,\dots,1)}\in L^{p}([0,1]^{d})$, [17] shows $V_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$. Crucially, $V^{(d)}_{\lambda}(f)$ can be finite even for discontinuous $f$ (see [22, Section 4.1]).
Next, for $v\subseteq 1{:}d$, we define the partial integral operator $I_{v}:L^{1}([0,1]^{d})\to L^{1}([0,1]^{|v|})$ as:
$I_{v}(f)(\boldsymbol{x}_{v})=\int_{[0,1]^{d-|v|}}f(\boldsymbol{x}_{v^{c}};\boldsymbol{x}_{v})\,\mathrm{d}\boldsymbol{x}_{v^{c}},$
where $f(\boldsymbol{x}_{v^{c}};\boldsymbol{x}_{v})$ denotes $f$ with fixed $\boldsymbol{x}_{v}$. Let $D_{v^{c}}$ be the class of functions:
$D_{v^{c}}=\Big{\{}\rho_{v^{c}}\Big{|}\rho_{v^{c}}(\boldsymbol{x})=\prod_{j\in v^{c}}\rho_{j}(x_{j}),\rho_{j}\in C([0,1]),\rho_{j}\geqslant 0,\|\rho_{j}\|_{\infty}\leqslant 1\Big{\}}.$ (11)
The partial fractional Vitali variation of order $\lambda$ with respect to $v\subseteq 1{:}d$ is:
$V^{v}_{\lambda}(f)=\sup_{\rho_{v^{c}}\in D_{v^{c}}}V^{(|v|)}_{\lambda}\Big{(}I_{v}(f\rho_{v^{c}})\Big{)}.$
By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [22, Section 3.2],
$\Bigl{(}V^{v}_{\lambda}(f)\Bigl{)}^{2}\leqslant\int_{[0,1]^{d-|v|}}\Bigl{(}V^{(|v|)}_{\lambda}\Bigl{(}f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr{)}\Bigr{)}^{2}\,\mathrm{d}\boldsymbol{x}_{v^{c}},$ (12)
implying $V^{v}_{\lambda}(f)<\infty$ if almost every section $f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})$ has bounded variation.
The following lemma connects the Walsh coefficients $\hat{f}(\boldsymbol{k})=\hat{f}_{u}(\boldsymbol{k})$ (where $\boldsymbol{s}(\boldsymbol{k})=u$) to the fractional Vitali variation of $f_{u}$:
Lemma 4.
Let $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})<\infty$ for $v\subseteq u\subseteq 1{:}s,u\neq\varnothing$ and $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ with $\alpha_{j}=\alpha\ \forall j\in v$. Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha_{j}$ if $j\in u\setminus v$ and $|\kappa_{j}|=\alpha+1$ if $j\in v$,
$\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}\leqslant 4^{|u|}\Big{(}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})\Big{)}^{2}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1},$
where
$B(\boldsymbol{k}_{u},v)=\{\boldsymbol{k}^{\prime}\in\mathbb{N}^{s}_{*}\mid\boldsymbol{s}(\boldsymbol{k}^{\prime})=u,\kappa^{\prime}_{j}=\kappa_{j}\text{if }j\in u\setminus v,\lceil\kappa^{\prime}_{j}\rceil_{1:(\alpha+1)}=\kappa_{j}\text{ if }j\in v\}.$ (13)
Proof.
Let $\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)$. Because $|\kappa^{\prime}_{j}|=|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$ and $|\kappa^{\prime}_{j}|\geqslant|\kappa_{j}|=\alpha+1$ for $j\in v$, we can apply Lemma 3 to get
$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=$
$\Big{|}\int_{[0,1]^{|u|}}\Big{(}f_{u}^{(\boldsymbol{\alpha}_{u})}(\boldsymbol{x})\prod_{j\in u\setminus v}W_{\kappa_{j}}(x_{j})\Big{)}\Big{(}\prod_{j\in v}{\mathrm{wal}_{\kappa^{\prime}_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}\alpha}}}(x_{j})W_{\lceil\kappa_{j}\rceil_{1{:}\alpha}}(x_{j})\Big{)}\,\mathrm{d}\boldsymbol{x}_{u}\Big{|}^{2},$
Let $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$, $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ and $\kappa^{-}_{j}=\kappa^{\prime}_{j}\setminus\kappa_{j}$ for $j\in v$. We can rewrite the above equation as
$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=\Bigl{|}\int_{[0,1]^{|v|}}g(\boldsymbol{x}_{v})\prod_{j\in v}{\mathrm{wal}_{\kappa^{-}_{j}}}(x_{j})(-1)^{\vec{x}_{j}(\ell_{j})}W_{\kappa^{+}_{j}}(x_{j})\,\mathrm{d}\boldsymbol{x}_{v}\Bigr{|}^{2},$
where
$g(\boldsymbol{x}_{v})=I_{v}\Big{(}f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}W_{\kappa_{j}}\Big{)}.$
From the proof of [22, Theorem 3], we can map $v$ to $1{:}|v|$ and show
$\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}\leqslant 2^{-|v|}\Big{(}V^{(|v|)}_{\lambda}(g)\Big{)}^{2}\Bigl{(}\prod_{j\in v}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}\Bigr{)}.$ (14)
Letting $\rho_{j}(x_{j})=W_{\kappa_{j}}(x_{j})/\|W_{\kappa_{j}}\|_{\infty}$ for $j\in u\setminus v$, equation (7) implies
$V^{(|v|)}_{\lambda}(g)=$ $\Big{(}\prod_{j\in u\setminus v}2\prod_{\ell^{\prime}\in\kappa_{j}}2^{-\ell^{\prime}-1}\Big{)}V^{(|v|)}_{\lambda}\Big{(}I_{u}(f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}\rho_{j})\Big{)}$
$\leqslant$ $2^{|u|-|v|}\Big{(}\prod_{j\in u\setminus v}\prod_{\ell^{\prime}\in\kappa_{j}}2^{-\ell^{\prime}-1}\Big{)}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})$
because $\prod_{j\in u\setminus v}\rho_{j}\in D_{v^{c}}$ from equation (11). Our conclusion follows once we put the above bound into equation (14) and plug in
$\prod_{j\in v}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}=4^{|v|}\prod_{j\in v}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
and $\lceil\kappa_{j}\rceil_{\alpha+1}=0$ for $j\in u\setminus v$. ∎
Motivated by Lemma 4, we define for $\alpha\in\mathbb{N}_{0},\lambda\in(0,1]$ and nonempty $u\subseteq 1{:}s$:
$\|f\|_{u,\alpha,\lambda}=\sup_{\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}}\sup_{v\subseteq\{j\mid\alpha_{j}=\alpha\}}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u}),$ (15)
where $f_{u}$ is the ANOVA component of $f$ defined in equation (8). We set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
$\|f\|_{\alpha,\lambda}=\sup_{u\subseteq 1{:}s}\|f\|_{u,\alpha,\lambda}.$
By Lemma 4, $\|f\|_{\alpha,\lambda}=0$ implies $\hat{f}(\boldsymbol{k})=0$ for all $\boldsymbol{k}\neq\boldsymbol{0}$, forcing $f$ to be a constant. For non-constant $f$, the relative variation of ANOVA component $f_{u}$ is:
$\gamma_{u}=\frac{\|f\|_{u,\alpha,\lambda}}{\|f\|_{\alpha,\lambda}},$ (16)
which satisfies $\sup_{u\subseteq 1{:}s}\gamma_{u}=1$ and $\gamma_{\varnothing}=0$. The decay of $\gamma_{u}$ encodes effective dimensionality, with $\gamma_{u}=1$ for all nonempty $u\subseteq 1{:}s$ representing the worst dimensionality. As shown in Theorem 4, dimension-independent convergence rates emerge under suitable decay conditions on $\gamma_{u}$.
Remark 1.
For positive weights $\{\tilde{\gamma}_{u},u\subseteq 1{:}s\}$, we can define the weighted norm:
$\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}=\sup_{u\subseteq 1{:}s}\tilde{\gamma}_{u}^{-1}\|f\|_{u,\alpha,\lambda}.$
All subsequent results hold equivalently for $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ replacing $\|f\|_{\alpha,\lambda}$ and $\tilde{\gamma}_{u}$ replacing $\gamma_{u}$. However, since $\|f\|_{\alpha,\lambda}$ and $\gamma_{u}$ are intrinsic to $f$, while $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ requires external weights $\tilde{\gamma}_{u}$ unused by our algorithm, we adopt the unweighted framework for simplicity.
Notably, $\|f\|_{\alpha,\lambda}$ equals the infimum of $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ over all weights $\{\tilde{\gamma}_{u},u\subseteq 1{:}s\}$ with $\sup_{u\subseteq 1{:}s}\tilde{\gamma}_{u}=1$. Thus, the median RQMC method implicitly selects optimal weights tailored to $f$, avoiding explicit weight specification.
4 Tractability under the completely random design
In this section, we show the median RQMC estimator $\hat{\mu}^{(r)}_{\infty}$ under the completely random design achieves a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate given $\|f\|_{\alpha,\lambda}<\infty$, with no dimension-dependence if further $\gamma_{u}$ satisfies equation (34).
First, we define $\|\kappa\|=\sum_{\ell\in\kappa}\ell$ for $\kappa\subseteq\mathbb{N}$, with $\|\varnothing\|=0$. Motivated by Lemma 4, we choose $K$ in equation (5) for given $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$ to be:
$K=\bigcup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}K_{u}(T_{u,m})$ (17)
with thresholds $T_{u,m}$ and
$K_{u}(T)=\bigcup_{v\subseteq u}\bigcup_{\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)}B(\boldsymbol{k}_{u},v)$
for $B(\boldsymbol{k}_{u},v)$ defined as in equation (13) and
$\mathcal{K}_{u,v}(T)=\Big{\{}(k_{j},j\in u)\in\mathbb{N}^{|u|}\Big{|}\begin{matrix}|\kappa_{j}|\leqslant\alpha\ \forall j\in u\setminus v,|\kappa_{j}|=\alpha+1\ \forall j\in v,\\ (\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|\leqslant T\end{matrix}.\Big{\}}$
We note that $K_{u}(T)$ implicitly depends on $\alpha$ and $\lambda$.
The next lemma quantifies how $T$ controls the size of $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$.
Lemma 5.
Suppose $\|f\|_{u,\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0},\lambda\in(0,1]$ and nonempty $u\subseteq 1{:}s$. Then for $T\in\mathbb{R}$ and $\theta\in(0,1)$,
$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T)}|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T}}C_{\alpha,\lambda,\theta}^{|u|},$
where
$C_{\alpha,\lambda,\theta}=\frac{4^{-\alpha-1}}{\alpha!(4^{(\alpha+\lambda)(1-\theta)}-1)(4^{1-\theta}-1)^{\alpha}}+\sum_{\alpha^{\prime}=0}^{\alpha}\frac{4^{-\alpha^{\prime}}}{(\alpha^{\prime})!(4^{1-\theta}-1)^{\alpha^{\prime}}}.$ (18)
Proof.
By equation (9), $\hat{f}_{u}(\boldsymbol{k})=0$ if $\boldsymbol{s}(\boldsymbol{k})\neq u$. Letting
$\mathbb{N}^{u}_{\alpha}=\{(k_{j},j\in u)\in\mathbb{N}^{|u|}\mid|\kappa_{j}|\leqslant\alpha+1\text{ for }j\in u\},$
we can apply Lemma 4 and equation (15) to get
$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T)}|\hat{f}_{u}(\boldsymbol{k})|^{2}$ (19)
$\leqslant$ $4^{|u|}\|f\|^{2}_{u,\alpha,\lambda}\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\boldsymbol{1}\Big{\{}\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)\Big{\}}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}.$
Because $\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)$ implies
$(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|>T,$
we have for any $\theta\in(0,1)$ and $\rho_{\theta}=4^{-1+\theta}$
$\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\boldsymbol{1}\Big{\{}\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)\Big{\}}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
$\leqslant$ $\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}4^{-\theta T+\theta(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\theta\sum_{j\in u}\|\kappa_{j}\|}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
$=$ $4^{-\theta T}\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\prod_{j\in u}4^{-(1-\theta)(\lambda-1)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-(1-\theta)\ell^{\prime}-1}$
$=$ $4^{-\theta T}\Big{(}\sum_{\alpha^{\prime}=0}^{\alpha}4^{-\alpha^{\prime}}\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\quad+\quad 4^{-\alpha-1}\sum_{\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})}\rho_{\theta}^{(\lambda-1)\lceil\kappa\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\Big{)}^{|u|},$
where $\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})=\{\kappa\subseteq\mathbb{N}\mid|\kappa|=\alpha^{\prime}\}$. To bound the above sum, consider the generating function
$g(t)=\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha^{\prime}},$
and let $[t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$. It follows that $[t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$. Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
$\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\leqslant\frac{1}{(\alpha^{\prime})!}\Big{(}\sum_{N=1}^{\infty}\rho_{\theta}^{N}\Big{)}^{\alpha^{\prime}}=\frac{1}{(\alpha^{\prime})!(\rho_{\theta}^{-1}-1)^{\alpha^{\prime}}}.$ (20)
For $\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})$, we let $\kappa^{+}=\lceil\kappa\rceil_{1{:}\alpha}$ and $\ell=\lceil\kappa\rceil_{\alpha+1}$. By the one-to-one correspondence between $\kappa^{+}$ and $\kappa^{+}-\ell=\{\ell^{\prime}-\ell\mid\ell^{\prime}\in\kappa\}\in\mathcal{S}_{\alpha}(\mathbb{N})$,
$\sum_{\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})}\rho_{\theta}^{(\lambda-1)\lceil\kappa\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}=$ $\sum_{\ell=1}^{\infty}\rho_{\theta}^{(\alpha+\lambda)\ell}\sum_{\kappa^{+}-\ell\in\mathcal{S}_{\alpha}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa^{+}-\ell}\rho_{\theta}^{\ell^{\prime}}$
$=$ $\sum_{\ell=1}^{\infty}\rho_{\theta}^{(\alpha+\lambda)\ell}\frac{1}{\alpha!(\rho_{\theta}^{-1}-1)^{\alpha}}$
$=$ $\frac{1}{\alpha!(\rho^{-\alpha-\lambda}_{\theta}-1)(\rho_{\theta}^{-1}-1)^{\alpha}}.$
Our conclusion follows once we put the above bounds into equation (19). ∎
While larger $T$ implies a tighter bound on $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$, it also increases the probability $Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K_{u}(T)$. The next lemma provides a bound on the cardinality of $K_{u}(T)$.
Lemma 6.
For $\alpha\in\mathbb{N}_{0},\lambda\in(0,1],T\in\mathbb{R}$ and nonempty $u\subseteq 1{:}s$,
$|K_{u}(T)|\leqslant 2^{T/(\alpha+\lambda)}\Big{(}\max\big{(}T/(\alpha+\lambda),1\big{)}A_{\alpha,\lambda}+B_{\alpha,\lambda}\Big{)}^{|u|},$
where
$A_{\alpha,\lambda}=\frac{1}{\alpha!(2^{1/(\alpha+\lambda)}-1)^{\alpha}}\quad\text{and}\quad B_{\alpha,\lambda}=\sum_{\alpha^{\prime}=0}^{\alpha}\frac{1}{(\alpha^{\prime})!(2^{1/(\alpha+\lambda)}-1)^{\alpha^{\prime}}}.$ (21)
Proof.
The conclusion is trivial when $T<0$, so we assume $T\geqslant 0$. As in the proof Lemma 4, for each $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ we write $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$ and $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ for $j\in v$. By equation (13), each $\boldsymbol{k}^{\prime}=(k^{\prime}_{1},\dots,k^{\prime}_{s})\in B(\boldsymbol{k}_{u},v)$ must have the same $\kappa^{\prime}_{j}$ for $j\in 1{:}s\setminus v$. For $j\in v$, however, $\kappa^{\prime}_{j}\setminus\kappa_{j}$ can be any subset of $1{:}(\ell_{j}-1)$ if $\ell_{j}\geqslant 2$ and $\kappa^{\prime}_{j}=\kappa_{j}$ if $\ell_{j}=1$. Thus, $|B(\boldsymbol{k}_{u},v)|=\prod_{j\in v}2^{\ell_{j}-1}$. Because $B(\boldsymbol{k}_{u},v)$ for distinct $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ are mutually disjoint,
$|K_{u}(T)|=\sum_{v\subseteq u}\sum_{\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)}\prod_{j\in v}2^{\ell_{j}-1}.$ (22)
Next, we notice
$(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|=(\alpha+\lambda)\sum_{j\in v}\ell_{j}+\sum_{j\in v}\sum_{\ell^{\prime}\in\kappa^{+}_{j}}(\ell^{\prime}-\ell_{j})+\sum_{j\in u\setminus v}\|\kappa_{j}\|$
and there is a one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}=\{\ell^{\prime}-\ell_{j}\mid\ell^{\prime}\in\kappa^{+}_{j}\}$ for $j\in v$. Furthermore, $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ implies
$\sum_{j\in v}\sum_{\ell^{\prime}\in\kappa^{+}_{j}}(\ell^{\prime}-\ell_{j})+\sum_{j\in u\setminus v}\|\kappa_{j}\|=\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|+\sum_{j\in u\setminus v}\|\kappa_{j}\|\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}.$
For $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$, let
$\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathbb{N}_{0}^{|u|}\Big{|}|\kappa_{j}|=\alpha_{j}\text{ for }j\in u,\sum_{j\in u}\|\kappa_{j}\|=N\Big{\}}.$
We can decompose $\mathcal{K}_{u,v}(T)$ into subsets with the same $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ and $(\kappa^{+}_{j}-\ell_{j},j\in v;\kappa_{j},j\in u\setminus v)\in\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})$ for $N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}$, and rewrite equation (22) as
$|K_{u}(T)|=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$ (23)
$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|+$
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
Let $\overline{N}=\lfloor(T-N)/(\alpha+\lambda)\rfloor$. For nonempty $v\subseteq u$ and integer $n\geqslant|v|$, there are (${n-1\choose|v|-1}$ number of $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfying $\|\boldsymbol{\ell}_{v}\|_{1}=n$. Hence, when $\overline{N}\geqslant|v|$,
$\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}=$ $\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}$ (24)
$=$ $\sum_{N=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n+1}-\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}$
$=$ (${\overline{N}\choose|v|-1}2^{\overline{N}+1}-\sum_{n=|v|}^{\overline{N}}\Big{\{}{n-1\choose|v|-1}-{n-2\choose|v|-1}\Big{\}}2^{n}$
$\leqslant$ (${\overline{N}\choose|v|-1}2^{\overline{N}+1}.$
When $\overline{N}<|v|$, no $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfies $\|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)$ and the above sum is trivially 00. Putting the above bound into equation (23), we get
$|K_{u}(T)|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|$ (25)
$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
$\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}.$
To bound the sum over $N$, consider the generating function
$g(t)=\prod_{j\in u}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}.$
Because each element of $\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})$ is counted $\prod_{j\in u}(\alpha_{j})!$ times by $t^{N}[g(t)]$ , an argument similar to equation (20) gives
$\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|t^{N}\leqslant\prod_{j\in u}\frac{1}{(\alpha_{j})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}=\prod_{j\in u}\frac{1}{(\alpha_{j})!(t^{-1}-1)^{\alpha_{j}}}$ (26)
for $t=2^{-1/(\alpha+\lambda)}$ and
$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\prod_{j\in u}\frac{1}{(\alpha_{j})!(2^{1/(\alpha+\lambda)}-1)^{\alpha_{j}}}$
$=$ $A_{\alpha,\lambda}^{|v|}B^{|u|-|v|}_{\alpha,\lambda},$
where $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are defined in equation (21). By a similar calculation,
$2^{-T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant B^{|u|}_{\alpha,\lambda}.$
Putting the above bounds into equation (25), we finally get
$|K_{u}(T)|\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}\Big{(}1+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}A_{\alpha,\lambda}^{|v|}B^{-|v|}_{\alpha,\lambda}\Big{)}$ (27)
$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}\Big{(}1+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}\max\big{(}T/(\alpha+\lambda),1\big{)}^{|v|}A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{-|v|}\Big{)}$
$=$ $2^{T/(\alpha+\lambda)}\Big{(}B_{\alpha,\lambda}+\max\big{(}T/(\alpha+\lambda),1\big{)}A_{\alpha,\lambda}\Big{)}^{|u|}.\qed$ (28)
Corollary 1.
For $m\in\mathbb{N}$ and $\gamma_{u}$ defined by equation (16), let
$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}.$ (29)
Then for any $\delta\in(0,1)$,
$\Pr\Big{(}Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in\bigcup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}K_{u}(T_{u,m})\Big{)}\leqslant\delta$
under the completely random design, where
$T_{u,m}=(\alpha+\lambda)m-(\alpha+\lambda)\Big{(}\log_{2}\big{(}\Gamma_{\alpha,\lambda,\gamma_{u}}(m)\big{)}-\log_{2}(\delta)-\frac{\log_{2}(\gamma_{u})}{\alpha+\lambda+1/2}\Big{)}.$ (30)
Proof.
Because $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)\geqslant 1,\delta<1$ and $\sup_{u\subseteq 1{:}s}\gamma_{u}=1$, we have $T_{u,m}/(\alpha+\lambda)\leqslant m$ for all nonempty $u\subseteq 1{:}s$. Lemma 6 then implies
$\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}|K_{u}(T_{u,m})|\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{T_{u,m}/(\alpha+\lambda)}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
$=$ $\frac{\delta 2^{m}}{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
$\leqslant$ $\delta 2^{m}.$
Under the completely random design, $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for any $\boldsymbol{k}\neq\boldsymbol{0}$, so our conclusion follows by taking a union bound. ∎
Now we are ready to show the following error bounds on $\hat{\mu}^{(r)}_{\infty}$:
Theorem 1.
Suppose $\|f\|_{\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Then for $m\in\mathbb{N},r\in\mathbb{N}$ and $\theta\in(0,1)$, the completely random design ensures
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\Gamma_{\alpha,\lambda,\gamma_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{(\Gamma_{\alpha,\lambda,\gamma_{u}}(m))^{2\theta(\alpha+\lambda)}}{2^{(2\theta(\alpha+\lambda)+1)m}}\Big{)}\leqslant 2^{-r},$ (31)
where $\gamma_{u}$ is given by equation (16), $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is given by equation (29) and
$\Gamma_{\alpha,\lambda,\gamma_{u},\theta}=16^{2\theta(\alpha+\lambda)+1}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1+2(1-\theta)(\alpha+\lambda)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|},$
with $C_{\alpha,\lambda,\theta}$ given by equation (18).
Proof.
Let $\delta=1/16$ in Lemma 2 and Corollary 1. By Lemma 2, equation (5) and $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for $\boldsymbol{k}\neq\boldsymbol{0}$, it suffices to show
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}\leqslant\Gamma_{\alpha,\lambda,\gamma_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2\theta(\alpha+\lambda)}}{2^{2\theta(\alpha+\lambda)(m-4)}}$
for $K$ defined by equation (17) with $T_{u,m}$ given by equation (30). From equation (9) and Lemma 5,
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}=\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$ (32)
Our conclusion follows after plugging in $\|f\|^{2}_{u,\alpha,\lambda}=\gamma^{2}_{u}\|f\|^{2}_{\alpha,\lambda}$ and equation (30). ∎
Remark 2.
Given a failure probability tolerance $\varepsilon>0$, we can set $r=\lceil\log_{2}(\varepsilon^{-1})\rceil$ to ensure $2^{-r}\leqslant\varepsilon$. For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, $N=r2^{m}$ is the total number of function evaluations.
To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$, so $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$. Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}\leqslant 1$ that
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big{)}\leqslant 2^{-r}\leqslant\varepsilon,$ (33)
where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is almost optimal [22, Subsection 3.2].
Remark 3.
In settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$, dimension-independent error bounds hold if there exist weights $\{\gamma_{j},j\in 1{:}s\}$ satisfying
$\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}\quad\forall u\subseteq 1{:}s\quad\text{ and }\quad\sum_{j=1}^{s}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\leqslant C_{\infty}$ (34)
for a constant $C_{\infty}$ independent of $s$. To see this, we first bound
$\frac{\Gamma_{\alpha,\lambda,\gamma_{u},\theta}}{16^{2\theta(\alpha+\lambda)+1}}\leqslant\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}\leqslant\prod_{j=1}^{s}(1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}C_{\alpha,\lambda,\theta})\leqslant\exp(C_{\infty}C_{\alpha,\lambda,\theta}),$
where we have used $1+x\leqslant\exp(x)$. Similarly,
$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=$ $1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
$\leqslant$ $\prod_{j=1}^{s}\Big{(}1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})\Big{)},$
which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ and a constant $C^{\prime}_{\eta}$ independent of $s$ [10, Lemma 3]. Substituting into equation (31), equation (33) holds with a constant $C_{\eta}$ independent of $s$, fulfilling the definition of strong tractability [20]. Notably, equation (34) resembles the tractability conditions assumed by [8, Corollary 3.13] for polynomial lattices to achieve a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate.
5 Tractability under more general randomization
While the completely random design offers theoretical advantages in near-optimal convergence rates and tractability, the random linear scrambling often outperforms it in practical settings due to tighter error bounds on low-dimensional ANOVA components. For example, [11] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets. In this section, we extend the analysis of Section 4 to a broader class of randomization schemes, with the random linear scrambling as a key example.
To state our assumption, we first define, for $d\in\mathbb{N}$ and finite subset $\kappa\subseteq\mathbb{N}$:
$\|\kappa\|_{(d)}=\begin{cases}\|\kappa\|,&\text{ if }|\kappa|\leqslant d\\ \|\lceil\kappa\rceil_{1{:}d}\|,&\text{ if }|\kappa|>d\end{cases}.$
Assumption 1.
Under the randomization scheme for $C_{j}\in\{0,1\}^{E\times m}$, $j\in 1{:}s$, there exists $d\in\mathbb{N}$ and a set of nonnegative integers $\tilde{t}_{u}$ for each nonempty $u\subseteq 1{:}s$ such that for $\boldsymbol{k}\in\mathbb{N}^{s}_{*}$ with $\boldsymbol{s}(\boldsymbol{k})=u$,
$\Pr(Z(\boldsymbol{k})=1)\leqslant 2^{-m+\tilde{t}_{u}}\boldsymbol{1}\Big{\{}\sum_{j\in u}\|\kappa_{j}\|_{(d)}>d(m-\tilde{t}_{u})\Big{\}}.$
From [8, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [8, Section 2]. We frame this assumption to align with broader applications in higher-order digital nets [4], where the parameter $d$ (the order of digital nets) may exceed $1$. This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
We first consider the $d\geqslant\alpha+\lambda$ case. The analysis is easy thanks to the good properties of our randomization.
Theorem 2.
Suppose $\|f\|_{\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Then for $m\in\mathbb{N},r\in\mathbb{N}$ and $\theta\in(0,1)$, a randomization scheme satisfying Assumption 1 with $d\geqslant\alpha+\lambda$ ensures
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\Gamma_{\alpha,\lambda,\gamma_{u},\tilde{t}_{u},\theta}\|f\|^{2}_{\alpha,\lambda}2^{-(2\theta(\alpha+\lambda)+1)m}\Big{)}\leqslant 2^{-r},$
where $\gamma_{u}$ is given by equation (16), $\tilde{t}_{u}$ is given in Assumption 1 and
$\Gamma_{\alpha,\lambda,\gamma_{u},\tilde{t}_{u},\theta}=16\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{2}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}(2\theta(\alpha+\lambda)+1)}$
with $C_{\alpha,\lambda,\theta}$ given by equation (18).
Proof.
Let $K$ be defined as in equation (17) with $T_{u,m}=(\alpha+\lambda)(m-\tilde{t}_{u})$. For any $\boldsymbol{k}^{\prime}=(k^{\prime}_{1},\dots,k^{\prime}_{s})\in B(\boldsymbol{k}_{u},v)$ with $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathcal{K}_{u,v}(T_{u,m})$, $\|\kappa^{\prime}_{j}\|_{(d)}=\|\kappa_{j}\|$ for $j\in u\setminus v$ because $|\kappa^{\prime}_{j}|=|\kappa_{j}|\leqslant d$ and
$\|\kappa^{\prime}_{j}\|_{(d)}\leqslant\frac{d}{\alpha+\lambda}\Big{(}(\lambda-1)\lceil\kappa_{j}\rceil_{\alpha+1}+\|\kappa_{j}\|\Big{)}$
for $j\in v$ because $\lceil\kappa^{\prime}_{j}\rceil_{1{:}(\alpha+1)}=\kappa_{j}$. Hence,
$\sum_{j\in u}\|\kappa^{\prime}_{j}\|_{(d)}\leqslant\frac{d}{\alpha+\lambda}\Big{(}(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|\Big{)}\leqslant\frac{d}{\alpha+\lambda}T_{u,m}=d(m-\tilde{t}_{u}).$
It follows from Assumption 1 that $\Pr(Z(\boldsymbol{k})=1)=0$ for all $\boldsymbol{k}\in K$ and we can apply Lemma 2 with $\delta=1/16$ to get
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>16\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}\Big{)}\leqslant 2^{-r}.$
Assumption 1 also implies $\Pr(Z(\boldsymbol{k})=1)\leqslant 2^{-m+\tilde{t}_{u}}$ if $\boldsymbol{s}(\boldsymbol{k})=u$. An argument similar to equation (32) gives
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}|\hat{f}_{u}(\boldsymbol{k})|^{2}$
$\leqslant$ $2^{-m}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{\tilde{t}_{u}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}$
Our conclusion follows after plugging in $\|f\|^{2}_{u,\alpha,\lambda}=\gamma^{2}_{u}\|f\|^{2}_{\alpha,\lambda}$ and $T_{u,m}=(\alpha+\lambda)(m-\tilde{t}_{u})$. ∎
Remark 4.
Analogous to Corollary 1, the bound in Theorem 2 can be refined by selecting thresholds $T_{u,m}\geqslant(\alpha+\lambda)(m-\tilde{t}_{u})$ while ensuring
$\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\Big{(}|K_{u}(T_{u,m})|-|K_{u}\big{(}(\alpha+\lambda)(m-\tilde{t}_{u})\big{)}|\Big{)}\leqslant\delta.$
However, this refinement offers limited practical improvement: Lemma 6 suggests $|K_{u}(T)|$ grows exponentially in $T$, so the margin $T_{u,m}-(\alpha+\lambda)(m-\tilde{t}_{u})$ is at best $O(\alpha+\lambda)$. Consequently, Consequently, the error bound improves by only a factor of $2^{O(\alpha+\lambda)}$, leaving the asymptotic convergence rate unchanged.
Next, we consider the $d<\alpha+\lambda$ case. We first establish the counterpart of Lemma 6.
Lemma 7.
For $d\in\mathbb{N},T^{\prime}\geqslant 0$ and nonempty $u\subseteq 1{:}s$, let
$K^{\prime}_{u}(T^{\prime})=\Big{\{}\boldsymbol{k}\in\mathbb{N}^{s}_{*}\mid\boldsymbol{s}(\boldsymbol{k})=u,\sum_{j\in u}\|\kappa_{j}\|_{(d)}>T^{\prime}\Big{\}}.$
Then for $\alpha\in\mathbb{N}_{0},\lambda\in(0,1]$ satisfying $\alpha+\lambda>d$, $\theta^{\prime}\in(0,1)$ and $T\in[T^{\prime},(\alpha+\lambda)T^{\prime}/d]$,
$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$
$\leqslant$ $2^{(T-T^{\prime})/\beta}\Big{(}\max\big{(}(T-T^{\prime})/\beta,1\big{)}A_{\alpha,\lambda}+B_{\alpha,\lambda}\Big{)}^{|u|}+2^{(T-\theta^{\prime}T^{\prime})/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}},$
where $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are given by equation (21), $\beta=\alpha+\lambda-d$ and
$D_{\alpha,\lambda,d,\theta^{\prime}}=\frac{1}{d!(2^{(1-\theta^{\prime})/\beta}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(2^{1/\beta}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(2^{(1-\theta^{\prime})/\beta}-1)^{\alpha^{\prime}}}.$
Proof.
We adopt the notation used in the proof of Lemma 6. For $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$, we again consider the one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}$ for $j\in v$. Because $d\in\mathbb{N}$ and $d<\alpha+\lambda$, we must have $d\leqslant\alpha$. By $|\kappa^{+}_{j}|=\alpha\geqslant d$ for $j\in v$,
$\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|_{(d)}=\sum_{j\in v}\|\kappa^{+}_{j}\|_{(d)}-d\ell_{j}=\Big{(}\sum_{j\in v}\|\kappa_{j}\|_{(d)}\Big{)}-d\|\boldsymbol{\ell}_{v}\|_{1}$
and
$\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|_{(d)}+\sum_{j\in u\setminus v}\|\kappa_{j}\|_{(d)}=\Big{(}\sum_{j\in u}\|\kappa_{j}\|_{(d)}\Big{)}-d\|\boldsymbol{\ell}_{v}\|_{1}.$
Let
$\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})\Big{|}\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N^{\prime}\Big{\}}.$
Similar to equation (23), we can write
$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$ (35)
$=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha,\forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}-d\|\boldsymbol{\ell}_{v}\|_{1}<N^{\prime}\leqslant N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$
$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|+$
$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
The sum over $\boldsymbol{\ell}_{v}$ is empty if $(T^{\prime}-N^{\prime})/d\geqslant(T-N)/(\alpha+\lambda)$. Because $T^{\prime}/d\geqslant T/(\alpha+\lambda)$ and $N\geqslant N^{\prime}$, we further infer the sum is empty unless $N>N_{*}$ for
$N_{*}=\frac{(\alpha+\lambda)T^{\prime}-dT}{\alpha+\lambda-d}.$
Therefore,
$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}\sum_{N^{\prime}=0}^{N}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1},$
where we have applied equation (24) in the last step. Letting $T_{*}=(T-N_{*})/(\alpha+\lambda)$, we put the above bound into equation (35) and get
$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (36)
$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
$\leqslant$ $2^{T_{*}}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T_{*}\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
$\leqslant$ $2^{T_{*}}\Big{(}B_{\alpha,\lambda}+\max(T_{*},1)A_{\alpha,\lambda}\Big{)}^{|u|},$
where the last step follows from a calculation similar to equation (27).
Next, we notice
$t^{T}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}t^{N^{\prime}}\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (37)
for $t=2^{-1/(\alpha+\lambda-d)}$. To bound the sum over $N$, we let $w=\{j\in u\mid\alpha_{j}>d\}$ for $\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and define $\boldsymbol{\alpha}^{-}_{u}=(\min(\alpha_{j},d),j\in u)$ and $\boldsymbol{\alpha}^{+}_{u}=(\alpha_{j}-d,j\in w)$ if $w\neq\varnothing$. For $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$, we further let $\kappa^{-}_{j,d}=\lceil\kappa_{j}\rceil_{1{:}\min(\alpha_{j},d)}$ for $j\in u$ and $\kappa^{+}_{j,d}=\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}d}$ for $j\in w$. Because for $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$,
$\sum_{j\in w}\|\kappa^{+}_{j,d}\|=\sum_{j\in w}\|\kappa_{j}\|-\sum_{j\in w}\|\kappa^{-}_{j,d}\|=\sum_{j\in u}\|\kappa_{j}\|-\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N-N^{\prime},$
the number of possible $(\kappa^{+}_{j,d},j\in w)$ for $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$ with a given set of $(\kappa^{-}_{j,d},j\in u)\in\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$ is bounded by the number of $(\kappa^{+}_{j,d},j\in w)$ with $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=N-N^{\prime}$, which is $|\mathcal{K}^{\prime}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ by definition. By summing this bound over all elements of $\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$, we get
$\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=N^{\prime}}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|t^{N-N^{\prime}}$
$\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{w}(N,\boldsymbol{\alpha}^{+}_{u})|t^{N}$
$\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}$
where we have applied equation (26) in the last step. Putting the above bound into equation (37), we get for any $\theta^{\prime}\in(0,1)$,
$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$
$\leqslant$ $t^{-T}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{N^{\prime}}$
$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{(1-\theta^{\prime})N^{\prime}}$
$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
where we have again applied equation (26). Our conclusion finally follows from equation (36), $t=2^{-1/(\alpha+\lambda-d)}$ and
$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
$=$ $\sum_{w\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{-}_{u}\in\mathbb{Z}^{|u|}_{\leqslant d}\\ w=\{j\mid\alpha^{-}_{j}=d\}\end{subarray}}\Big{(}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}^{|u|}$
$=$ $\sum_{w\subseteq u}\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|-|w|}$
$=$ $\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|}.\qed$
Corollary 2.
For $m\in\mathbb{N}$ and $\gamma_{u}$ defined by equation (16), let
$\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$ (38)
Then for any $\delta\in(0,1)$,
$\Pr\Big{(}Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in\bigcup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}K_{u}(T_{u,m})\Big{)}\leqslant\delta$
under a randomization scheme satisfying Assumption 1 with $d<\alpha+\lambda$, where
$T_{u,m}=(\alpha+\lambda)(m-\tilde{t}_{u})-(\alpha+\lambda-d)\Big{(}\log_{2}\big{(}\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\big{)}-\log_{2}(\delta)-\frac{\log_{2}(\gamma_{u})}{\alpha+\lambda+1/2}\Big{)}.$ (39)
Proof.
Let $T^{\prime}_{u,m}=d(m-\tilde{t}_{u})$. By Assumption 1, for any $\boldsymbol{k}\in\mathbb{N}^{s}_{*}$ with $\boldsymbol{s}(\boldsymbol{k})=u$, $\Pr(Z(\boldsymbol{k})=1)=0$ if $\boldsymbol{k}\notin K^{\prime}_{u}(T^{\prime}_{u,m})$ and $\Pr(Z(\boldsymbol{k})=1)\leqslant 2^{-m+\tilde{t}_{u}}$ if $\boldsymbol{k}\in K^{\prime}_{u}(T^{\prime}_{u,m})$. Because $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\geqslant 1,\delta<1$ and $\sup_{u\subseteq 1{:}s}\gamma_{u}=1$, we have $T_{u,m}\leqslant(\alpha+\lambda)(m-\tilde{t}_{u})=(\alpha+\lambda)T^{\prime}_{u,m}/d$. If further $T_{u,m}\geqslant T^{\prime}_{u,m}$, we can apply Lemma 7 to get
$|K_{u}(T_{u,m})\cap K^{\prime}_{u}(T^{\prime}_{u,m})|$
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}2^{m-\tilde{t}_{u}}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{(1-\theta^{\prime})dm/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
A union bound over all $\boldsymbol{k}\in K_{u}(T_{u,m})\cap K^{\prime}_{u}(T^{\prime}_{u,m})$ implies
$\Pr(Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in K_{u}(T_{u,m}))$
$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{(1-\theta^{\prime})dm/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
On the other hand, when $T_{u,m}<T^{\prime}_{u,m}$, every $\boldsymbol{k}\in K_{u}(T_{u,m})$ satisfies
$\sum_{j\in u}\|\kappa_{j}\|_{(d)}\leqslant\sum_{j\in u}\|\kappa_{j}\|\leqslant T_{u,m}<T^{\prime}_{u,m}.$
Thus, $K_{u}(T_{u,m})\cap K^{\prime}_{u}(T^{\prime}_{u,m})=\varnothing$ and $\Pr(Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in K_{u}(T_{u,m}))=0$. The conclusion follows by taking an union bound over all nonempty $u\subseteq 1{:}s$. ∎
The following theorem is the counterpart of Theorem 1 when $d<\alpha+\lambda$.
Theorem 3.
Suppose $\|f\|_{\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Then for $m\in\mathbb{N},r\in\mathbb{N},\theta\in(0,1)$ and $\theta^{\prime}\in(0,1)$, a randomization scheme satisfying Assumption 1 with $d<\alpha+\lambda$ ensures
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{(\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m))^{2\theta(\alpha+\lambda-d)}}{2^{(2\theta(\alpha+\lambda)+1)m}}\Big{)}\leqslant 2^{-r},$
where $\gamma_{u}$ is given by equation (16), $\tilde{t}_{u}$ is given in Assumption 1 and
$\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}=16^{2\theta(\alpha+\lambda-d)+1}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1+2\theta d+2(1-\theta)(\alpha+\lambda)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}(2\theta(\alpha+\lambda)+1)}$
with $C_{\alpha,\lambda,\theta}$ given by equation (18).
Proof.
Let $\delta=1/16$ in Lemma 2 and Corollary 2. By Lemma 5 and Assumption 1,
$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}\Pr(Z(\boldsymbol{k})=1)|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant 2^{-m+\tilde{t}_{u}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}$
for $T_{u,m}$ defined by equation (39). From Equation (9) and (16),
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}\leqslant\|f\|^{2}_{\alpha,\lambda}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\frac{\gamma^{2}_{u}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$ (40)
Our conclusion finally follows from Lemma 2 and equation (5). ∎
Remark 5.
By selecting $\theta$ and $\theta^{\prime}$ sufficiently close to $1$, an argument analogous to Remark 2 shows $\hat{\mu}^{(r)}_{\infty}$ achieves an asymptotic convergence rate of $O(2^{-(\alpha+\lambda+1/2-\eta)m})$ for any $\eta>0$.
Remark 6.
Consider again settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$. We assume equation (34) holds for weights $\{\gamma_{j},j\in 1{:}s\}$. Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for $C_{\eta}$ independent of $s$. Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [27] that the base-2 Niederreiter sequence [19] satisfies
$t_{u}\leqslant\sum_{j\in u}\log_{2}(j)+\log_{2}\log_{2}(j+2)+2,$ (41)
while the Sobol’ sequence [25] satisfies
$t_{u}\leqslant\sum_{j\in u}\log_{2}(j)+\log_{2}\log_{2}(j+1)+\log_{2}\log_{2}\log_{2}(j+3)+C$
for an absolute constant $C$. For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for a nonnegative sequence $\{\tilde{t}_{j},j\in 1{:}s\}$. We can then bound
$\frac{\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}}{16^{2\theta(\alpha+\lambda-d)+1}}\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\prod_{j\in u}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}C_{\alpha,\lambda,\theta}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}$
$\leqslant$ $\prod_{j=1}^{s}\Big{(}1+\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}C_{\alpha,\lambda,\theta}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}\Big{)}$
$\leqslant$ $\exp\Big{(}C_{\alpha,\lambda,\theta}\sum_{j=1}^{s}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}\Big{)}.$
Therefore, $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ has a dimension-independent bound if
$\lim_{s\to\infty}\sum_{j=1}^{s}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
Remark 7.
Viewing $\Gamma_{\alpha,\lambda,\gamma_{u},\tilde{t}_{u},\theta}$ from Theorem 2 as a special case of $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ with $d=\alpha+\lambda$, the argument in Remark 6 yields dimension-independent error bounds for the $d\geqslant\alpha+\lambda$ case if $\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}$, $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ and
$\lim_{s\to\infty}\sum_{j=1}^{s}\gamma_{j}^{2}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [4], and unbounded increases in $d$ ultimately violate the tractability condition above. Thus, prudent selection of $d$ is critical for practical implementations.
When $d<\alpha+\lambda$, we can refine our argument to establish strong tractability under a weaker condition analogous to that outlined in Remark 7.
Theorem 4.
Suppose $\|f\|_{\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$. Given a randomization scheme satisfying Assumption 1 with $d<\alpha+\lambda$, we further assume $\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}$ for $\{\gamma_{j},j\in 1{:}s\}$, $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for $\{\tilde{t}_{j},j\in 1{:}s\}$ and
$\lim_{s\to\infty}\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big{)}^{\phi}<\infty\quad\text{for}\quad\phi=\frac{\alpha+\lambda+1/2}{\alpha+\lambda-d+1/2}.$ (42)
Then under the given randomization scheme, we can find a constant $C_{\eta}$ independent of $s$ for any $\eta>0$ such that for all $m\in\mathbb{N}$ and $r\in\mathbb{N}$,
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big{)}\leqslant 2^{-r}.$
Proof.
By equation (42), only finitely many $j\in\mathbb{N}$ satisfy $\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\geqslant 1$. Hence,
$\sup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\leqslant\sup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\prod_{j\in u}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\leqslant\Lambda$
for a constant $\Lambda$ independent of $s$. Let $\delta^{\prime}=\min(\Lambda^{-\phi},1/16)$, $\theta^{\prime}=1-\eta d,$
$\Gamma(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}$
and
$T_{u,m}=$ $(\alpha+\lambda)(m-\tilde{t}_{u})-$
$(\alpha+\lambda-d)\Big{(}\log_{2}(\Gamma(m))-\log_{2}(\delta^{\prime})-\frac{\phi\log_{2}(\gamma_{u})}{\alpha+\lambda+1/2}-\phi\tilde{t}_{u}\Big{)}.$
By our choice of $\delta^{\prime}$,
$\log_{2}(\delta^{\prime})+\frac{\phi\log_{2}(\gamma_{u})}{\alpha+\lambda+1/2}+\phi\tilde{t}_{u}=\log_{2}(\delta^{\prime})+\phi\log_{2}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}\leqslant 0.$
Therefore, $T_{u,m}\leqslant(\alpha+\lambda)(m-\tilde{t}_{u})$ and an argument similar to Corollary 2 shows
$\Pr\Big{(}Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in\bigcup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}K_{u}(T_{u,m})\Big{)}$
$\leqslant$ $\frac{\delta^{\prime}}{\Gamma(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)},$
which is bounded by $\delta^{\prime}$ due to the definition of $\Gamma(m)$. Equation (40) with the above choice of $T_{u,m}$ then implies
$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}\leqslant$ $\|f\|^{2}_{\alpha,\lambda}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\frac{\gamma^{2}_{u}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}$
$=$ $\Gamma^{\prime}\|f\|^{2}_{\alpha,\lambda}\frac{(\Gamma(m)/\delta^{\prime})^{2\theta(\alpha+\lambda-d)}}{2^{(2\theta(\alpha+\lambda)+1)m}},$
where
$\Gamma^{\prime}=\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{2-\frac{2\phi\theta(\alpha+\lambda-d)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}\big{(}1+2\theta(\alpha+\lambda)-2\theta\phi(\alpha+\lambda-d)\big{)}}.$
Like in Remark 6, we can show $\Gamma(m)\leqslant C^{\prime}_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C^{\prime}_{\eta}$ independent of $s$ when equation (42) holds. Furthermore, we notice that $\alpha+\lambda>\phi(\alpha+\lambda-d)$ and
$\Gamma^{\prime}\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{2-\frac{2\phi(\alpha+\lambda-d)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}\big{(}1+2(\alpha+\lambda)-2\phi(\alpha+\lambda-d)\big{)}}$
$=$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}C^{|u|}_{\alpha,\lambda,\theta},$
which is bounded as $s\to\infty$ given equation (42). Our conclusion then follows from Lemma 2 with $\delta=1/16$ and equation (5). ∎
Remark 8.
By Hölder’s inequality,
$\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big{)}^{\phi}$
$\leqslant$ $\Big{(}\sum_{j=1}^{s}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\Big{)}^{1-\frac{\phi}{1+2(\alpha+\lambda)}}\Big{(}\sum_{j=1}^{s}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}\Big{)}^{\frac{\phi}{1+2(\alpha+\lambda)}}.$
Therefore, equation (42) holds under the assumptions in Remark 6.
Remark 9.
For random linear scrambling using base-2 Niederreiter sequences, substituting the bound (41) into equation (42) yields the simplified tractability condition:
$\lim_{s\to\infty}\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}j\log_{2}(j)\Big{)}^{\phi}<\infty.$
Since $\phi>1$, this condition is strictly weaker than the requirements in [8, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, strong tractability holds if
$\phi(\frac{p}{\alpha+\lambda+1/2}-q)=\frac{p-(\alpha+\lambda+1/2)q}{\alpha+\lambda-d+1/2}>1.$ (43)
This criterion interpolates between the tractability condition for the completely random design (by setting $d=0$ and $q=0$) and that for randomization schemes with $d\geqslant\alpha+\lambda$ (by setting $d=\alpha+\lambda$).
6 Numerical experiments
In this section, we test our theory across integrands with varying dimension and decay on $\gamma_{u}$. The test integrands have the common form
$f_{s,\gamma,\alpha}(\boldsymbol{x})=\prod_{j=1}^{s}\Big{(}1+\frac{c_{\alpha}}{j^{\gamma}}(x_{j}\exp(x_{j})-1)\Big{)},$
where $s=10,100,1000$ controls the dimension of integration domain, $\gamma=2,3,4$ controls the decay of $\gamma_{u}$ and $c_{\alpha}$ for $\alpha=0,1$ is a normalizing constant to ensure $\|f\|_{\alpha,1}=1$. To determine $c_{\alpha}$, we start by computing the ANOVA component given by equation (8):
$f_{u}(\boldsymbol{x})=\frac{c^{|u|}_{\alpha}}{(\prod_{j\in u}j)^{\gamma}}\prod_{j\in u}(x_{j}\exp(x_{j})-1).$ (44)
For $\lambda=1$ and smooth function $f$, the Vitali variation defined in equation (10) can be viewed as a Riemann sum for
$\Big{(}\int_{[0,1]^{d}}\Big{(}f^{(1,\dots,1)}(\boldsymbol{x})\Big{)}^{2}\,\mathrm{d}\boldsymbol{x}\Big{)}^{1/2}.$
Hence,
$V^{(|u|)}_{1}(f_{u})=\|f^{(1,\dots,1)}_{u}\|_{L^{2}}=\frac{c^{|u|}_{\alpha}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}.$
By equation (12), for $v\subseteq u$,
$\Big{(}V^{v}_{\lambda}(f_{u})\Big{)}^{2}$
$\leqslant$ $\frac{c^{2|u|}_{\alpha}}{(\prod_{j\in u}j)^{2\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|v|}\Big{(}\int_{0}^{1}(x\exp(x)-1)^{2}\,\mathrm{d}x\Big{)}^{|u|-|v|}$
$\leqslant$ $\Big{(}V^{|u|}_{\lambda}(f_{u})\Big{)}^{2},$
while $V^{u}_{\lambda}(f_{u})=V^{|u|}_{\lambda}(f_{u})$. Therefore,
$\|f\|_{u,0,1}=\sup_{v\subseteq u}V^{v}_{1}(f_{u})=V^{u}_{1}(f_{u})=\frac{c^{|u|}_{\alpha}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}$
and we can make $\|f_{s,\gamma,0}\|_{0,1}=1$ by setting
$c_{0}=\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{-1/2}.$
A similar calculation shows $\|f_{s,\gamma,1}\|_{1,1}=1$ if we set
$c_{1}=\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{-1/2}.$
With the above choice of $c_{\alpha}$, it also follows that $\gamma_{u}=\prod_{j\in}\gamma_{j}$ for $\gamma_{j}=j^{-\gamma}$.
For each integrand $f_{s,\gamma,\alpha}$, we evaluate the performance of the median estimator $\hat{\mu}^{(r)}_{E}$ for $m=1,2,\dots,16$ under two randomization schemes: the completely random design (CRD) and the random linear scrambling (RLS). The estimator $\hat{\mu}^{(r)}_{E}$ is computed as the sample median of $12r-$ 1 independent realizations of $\hat{\mu}_{E}$, where $\hat{\mu}_{E}$ is defined as in equation (2) with precision $E=64$. The observed integration errors in our tested range of $m$ is much larger than $2^{-E}$, ensuring that $\hat{\mu}^{(r)}_{E}$ closely approximates $\hat{\mu}^{(r)}_{\infty}$. Consequently, the error bounds in Theorem 1, 2 and 3 apply to $\hat{\mu}^{(r)}_{E}$ with probability at least $1-2^{-r}$. We set $r=10$, limiting the failure probability per data point to below $0.1\%$. As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the average of the $12r-$ 1 $\hat{\mu}_{E}$ realizations used by RLS. The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [11].
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Integration error for $f_{s,\gamma,\alpha}$ with $\alpha=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Integration error for $f_{s,\gamma,\alpha}$ with $\alpha=1$.
Figure 1 and Figure 2 summarize the simulation results. First, we observe that errors for $\alpha=0$ and $\alpha=1$ exhibit nearly identical convergence rates. Surprisingly, for $\alpha=0$ and $\gamma=3$ or $4$, the empirical rate approaches $O(2^{-2m})$, despite the theoretical bound $O(2^{-(\alpha+\lambda+1/2)m})=O(2^{-(3/2)m})$. This phenomenon arises because $f_{s,\gamma,0}$ also has a finite $\|\cdot\|_{1,1}$ norm. By equation (44), the ANOVA component $f_{u}$ of $f_{s,\gamma,0}$ equals $(c_{0}/c_{1})^{|u|}$ times that of $f_{s,\gamma,1}$. Consequently,
$\|f_{s,\gamma,0}\|_{u,1,1}=\Big{(}\frac{c_{0}}{c_{1}}\Big{)}^{|u|}\|f_{s,\gamma,1}\|_{u,1,1}=\Big{(}\frac{c_{0}}{c_{1}}\Big{)}^{|u|}\prod_{j\in u}j^{-\gamma}$
and
$\|f_{s,\gamma,0}\|_{1,1}=\sup_{u\subseteq 1{:}s}\|f_{s,\gamma,0}\|_{u,1,1}=\sup_{u\subseteq 1{:}s}\prod_{j\in u}\Big{(}\frac{c_{0}}{c_{1}}j^{-\gamma}\Big{)}<\infty$
as $j^{-\gamma}>c_{1}/c_{0}$ holds for only finitely many $j\in\mathbb{N}$. Furthermore, because $\|f_{s,\gamma,0}\|_{1,1}\geqslant 1$, the relative variation $\gamma^{\prime}_{u}$ of $f_{s,\gamma,0}$ under the $\|\cdot\|_{1,1}$ norm satisfies
$\gamma^{\prime}_{u}=\frac{\|f_{s,\gamma,0}\|_{1,1,u}}{\|f_{s,\gamma,0}\|_{1,1}}\leqslant\prod_{j\in u}\Big{(}\frac{c_{0}}{c_{1}}j^{-\gamma}\Big{)}=\prod_{j\in u}\gamma^{\prime}_{j}$
for $\gamma^{\prime}_{j}=(c_{0}/c_{1})j^{-\gamma}$. In general, this analysis extends to arbitrary norms $\|\cdot\|_{\alpha^{\prime},\lambda^{\prime}}$, $\alpha^{\prime}\in\mathbb{N},\lambda^{\prime}\in(0,1]$, and the corresponding $\gamma_{j}$ is always proportional to $j^{-\gamma}$. Thus, the decay rate $\gamma$ dominates convergence behavior as $s\to\infty$, irrespective of $\alpha$.
Next, we observe that for a fixed value of $\gamma$, the integration errors exhibit similar behavior as $s$ varies, aligning with our analysis of dimension-independent convergence. Applying equation (43) with parameters $d=1$, $p=\gamma$ and $q=1+\eta$ for any $\eta>0$, our theory predicts that for large $s$, $\hat{\mu}^{(r)}_{\infty}$ under RLS achieve a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate, provided $\alpha+\lambda+1/2<(\gamma+1)/2$. Empirically, the observed rates approximate $O(2^{-(3/2)m})$ when $\gamma=2$ and $O(2^{-2m})$ when $\gamma=3$. For $\gamma=4$, the convergence rates slightly exceed $O(2^{-2m})$ but remain significantly slower than $O(2^{-(5/2)m})$. This discrepancy is likely due to the limited range of $m$ tested. Notably, while our theory suggests that $\hat{\mu}^{(r)}_{\infty}$ under CRD can attain a near $O(2^{-(\alpha+\lambda+1/2)m})$ rate for $\alpha+\lambda+1/2<\gamma$, practical results show that CRD consistently underperforms RLS. We hypothesize that the superior theoretical rates for CRD may only manifest at substantially larger $m$.
Finally, we compare the integration errors under RLS and STD. Across the vast majority of test cases, RLS demonstrates superior performance over STD, reinforcing that the median RQMC approach (via RLS) is a more robust alternative to traditional RQMC (via STD) in high-dimensional settings. We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m}$ rate derived from the root mean squared error when $\gamma=3$ or $4$. This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 2 and 3. However, averaging multiple replicates in STD compromises these bounds, as even a single outlier among the replicates can dominate the error and inflate its magnitude. This is the key mechanism that motivates our shift toward median RQMC approaches.
7 Discussion
Our current analysis presumes fixed constants $\alpha$ and $\lambda$ for which $\|f\|_{\alpha,\lambda}<\infty$. In practice, integrands are expected to satisfy $\|f\|_{\alpha,\lambda}<\infty$ across a spectrum of $\alpha$ and $\lambda$ values, as demonstrated by $f_{s,\gamma,\alpha}$ in Section 6. Crucially, since median RQMC methods do not explicitly incorporate smoothness information, the resulting error bound achieves the minimum over all viable $\alpha$ and $\lambda$. A more nuanced analysis should consequently incorporate variable smoothness parameters and derive error bounds under the optimal ones. Developing a framework for such an analysis is an open problem we leave for future research.
Another unresolved challenge is how to select $d$ for our randomization. As noted in Remark 7, $\tilde{t}_{u}$ generally scales with $d$. Theorem 2 further implies that increasing $d$ beyond $\alpha+\lambda$ does not enhance the convergence rate, suggesting the upper bound $d\leqslant\lceil\alpha+\lambda\rceil$. This constraint, however, offers little insights for high-dimensional integration of smooth integrands, where dimensionality—rather than smoothness—governs the convergence rate. A critical open question is how to rigorously define an effective smoothness criterion that jointly incorporates dimensionality and traditional smoothness parameters. Addressing this could yield a principled framework for selecting $d$, thereby optimizing integration performance.
A natural extension of our analysis is to integrands with discontinuities in their derivatives or function values. As discussed in Subsection 2.4, the requirement $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ can be relaxed provided that the weak derivatives satisfy equation (6). Notably, the fractional Vitali variation framework inherently accommodates discontinuities. By developing a theory for non-smooth integrands, our results could be extended to real-world applications in finance and related fields. For instance, [9] showcases an infinite-dimensional integration problem where all ANOVA components are smooth despite discontinuities in the integrand’s derivatives, suggesting that non-smoothness may not fundamentally compromise convergence rates in high-dimensional settings.
Acknowledgments
The author acknowledges the support of the Austrian Science Fund (FWF) Project DOI 10.55776/P34808. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v2

@@
 Building on this framework, we prove that the median estimator attains dimension-independent convergence under tractability conditions characterized by low effective dimensionality, a property known as strong tractability in complexity theory.
-Our analysis strengthens existing guarantees by improving the convergence rates and relaxing the theoretical assumptions previously required for strong tractability.
+Our analysis strengthens existing guarantees by improving the convergence rates and relaxing the theoretical assumptions previously required for dimension-independent convergence.
 1 Introduction
@@
 A critical gap in these prior studies is the role of dimensionality.
-We focus on the regime where the dimension grows indefinitely while the variance remains bounded—a common scenario in path integral approximations [14].
+We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [17].
 Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [23], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3).
-This notion of relative variation aligns with the effective dimension framework introduced in [1], which quantifies how high-dimensional integrands can behave as if they reside in lower-dimensional subspaces.
-The connection to weighted function spaces is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
+This notion of relative variation aligns with the effective dimension framework introduced in [1], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions.
+The connection to weighted function spaces [28] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
 Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [9].
-A key distinction lies in the randomization: their framework omits digital shifts (Subsection 2.1), a critical component of our scrambling method, and employs worst-case integration error as the primary metric.
-In contrast, we demonstrate that median estimators incorporating digital shifts achieve near-optimal convergence rates with high probability, avoiding worst-case error analysis entirely.
-This probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases.
+A key distinction lies in the randomization: their framework omits digital shifts (Subsection 2.1), a critical component of our scrambling method, and employs worst-case integration error over the unit ball in a weighted Sobolev space as the primary metric.
+In contrast, we demonstrate that median estimators incorporating digital shifts achieve near-optimal convergence rates on a fixed integrand with high probability, avoiding worst-case error analysis entirely.
+This probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [14, 15].
 Within this framework, we improve the convergence rates from nearly $O(n^{-\alpha})$ in [9] to nearly $O(n^{-\alpha-1/2})$, where $\alpha$ denotes the smoothness parameter in their analysis (we note that $\alpha$ is defined differently in this paper).
-Furthermore, we relax the strong tractability conditions, as detailed in Remarks 3 and 7.
-Specifically, we show digital nets under the completely random design (Subsection 2.1) achieve strong tractability under conditions similar to those for polynomial lattices in [8], while digital nets under the random linear scrambling attain the same property under strictly weaker conditions than those assumed in [8].
+Furthermore, we relax the conditions for dimension-independent convergence (Remarks 3 and 9).
+Specifically, we show digital nets under the completely random design (Subsection 2.1) achieve dimension-independent convergence under conditions comparable to those for polynomial lattices in [9], while digital nets under the random linear scrambling attain the same property under strictly weaker conditions than those assumed in [9].
 The remainder of this paper is structured as follows.
@@
 $f(\boldsymbol{x})=\sum_{\boldsymbol{k}\in\mathbb{N}_{0}^{s}}\hat{f}(\boldsymbol{k})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x}),\quad\text{where}\quad\hat{f}(\boldsymbol{k})=\int_{[0,1]^{s}}f(\boldsymbol{x})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}.$ (3)
-The equality holds in the $L^{2}$-sense. Using this decomposition, [24] establishes a key error structure for QMC estimators:
+The equality holds in the $L^{2}$-sense. Using this decomposition, [25, Theorem 1] establishes a key error structure for QMC estimators:
 Lemma 1.
@@
 2.3 Probabilistic error guarantee of the median
-The analysis of equation (4) uses two key properties of the coefficients $S(\boldsymbol{k})$: they are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [23, Lemma 4] for the univariate case $s=1$).
+The analysis of equation (4) uses two key properties of the coefficients $S(\boldsymbol{k})$: they are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [26, Lemma 4] for the univariate case $s=1$).
 These properties allow us to compute the variance of $\hat{\mu}_{\infty}$ and construct probabilistic error bounds by Chebyshev’s inequality.
@@
 Proof.
-This is an intermediate result proven in [24, Lemma 6]. ∎
+This is an intermediate result proven in [27, Lemma 6]. ∎
 For $f\in C([0,1]^{s})$, the above lemma implies the mean squared error bound
@@
 We choose $\mathcal{A}$ to be the event $\{Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K\}$ for a carefully chosen index set $K\subseteq\mathbb{N}^{s}_{*}$.
-It is shown in [22, Subsection 2.4] that
+It is shown in [25, Subsection 2.4] that
 ${\Pr(\mathcal{A}^{c})}\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\leqslant\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}.$ (5)
@@
 Proof.
-Equation (6) comes from [26, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [26, Section 3]. ∎
+Equation (6) comes from [30, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [30, Section 3]. ∎
 For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $[0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$.
 In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3.
-This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (6), as discussed in [22, Remark 1].
+This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (6), as discussed in [25, Remark 1].
 3 ANOVA decomposition and effective dimensionality
@@
 By an abuse of notation, we write $f_{u}(\boldsymbol{x})=f_{u}(\boldsymbol{x}_{u})$, where the latter $f_{u}:[0,1]^{|u|}\to\mathbb{R}$ is a $|u|$-dimensional function.
-Let $\boldsymbol{s}(\boldsymbol{k})=\{j\in 1{:}s\mid k_{j}\neq 0\}$ denote the support of $\boldsymbol{k}=(k_{1},\dots,k_{s})$. From the Walsh decomposition (3),
+Let $\boldsymbol{s}(\boldsymbol{k})=\{j\in 1{:}s\mid k_{j}\neq 0\}$ denote the support of $\boldsymbol{k}=(k_{1},\dots,k_{s})$. From the Walsh decomposition (3), we know that
 $f_{u}(\boldsymbol{x})=\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{N}^{s}_{0}\\ \boldsymbol{s}(\boldsymbol{k})=u\end{subarray}}\hat{f}(\boldsymbol{k})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x}),$ (9)
-which follows from the fact that $\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})$ depends only on $\{x_{j}\mid j\in\boldsymbol{s}(\boldsymbol{k})\}$ and
+because $\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})$ depends only on $\{x_{j}\mid j\in\boldsymbol{s}(\boldsymbol{k})\}$ and
 $\int_{0}^{1}\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})\,\mathrm{d}x_{j}=\Big{(}\int_{0}^{1}{\mathrm{wal}_{\kappa_{j}}}(x_{j})\,\mathrm{d}x_{j}\Big{)}\prod_{j^{\prime}\in\boldsymbol{s}(\boldsymbol{k})\setminus\{j\}}{\mathrm{wal}_{\kappa_{j^{\prime}}}}(x_{j^{\prime}})=0$
@@
 where the supremum is over all partitions $\mathcal{P}$ of $[0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$.
-This generalizes Sobolev norms: for $f$ with weak mixed derivative $f^{(1,\dots,1)}\in L^{p}([0,1]^{d})$, [17] shows $V_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$.
-Crucially, $V^{(d)}_{\lambda}(f)$ can be finite even for discontinuous $f$ (see [22, Section 4.1]).
+When $d=0$ and $f\in\mathbb{R}$, $V^{(0)}_{\lambda}(f)=|f|$ by convention.
+The Vitali variation generalizes Sobolev norms: for $f$ with weak mixed derivative $f^{(1,\dots,1)}\in L^{p}([0,1]^{d})$, [20] shows $V_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$.
+Crucially, $V^{(d)}_{\lambda}(f)$ can be finite even for discontinuous $f$ (see [25, Section 4.1] for an example).
 Next, for $v\subseteq 1{:}d$, we define the partial integral operator $I_{v}:L^{1}([0,1]^{d})\to L^{1}([0,1]^{|v|})$ as:
@@
 The partial fractional Vitali variation of order $\lambda$ with respect to $v\subseteq 1{:}d$ is:
-$V^{v}_{\lambda}(f)=\sup_{\rho_{v^{c}}\in D_{v^{c}}}V^{(|v|)}_{\lambda}\Big{(}I_{v}(f\rho_{v^{c}})\Big{)}.$
-By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [22, Section 3.2],
-$\Bigl{(}V^{v}_{\lambda}(f)\Bigl{)}^{2}\leqslant\int_{[0,1]^{d-|v|}}\Bigl{(}V^{(|v|)}_{\lambda}\Bigl{(}f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr{)}\Bigr{)}^{2}\,\mathrm{d}\boldsymbol{x}_{v^{c}},$ (12)
+$V^{v}_{\lambda}(f)=\sup_{\rho_{v^{c}}\in D_{v^{c}}}V^{(|v|)}_{\lambda}\Big{(}I_{v}(f\rho_{v^{c}})\Big{)}.$ (12)
+By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [25, Section 3.2],
+$\Bigl{(}V^{v}_{\lambda}(f)\Bigl{)}^{2}\leqslant\int_{[0,1]^{d-|v|}}\Bigl{(}V^{(|v|)}_{\lambda}\Bigl{(}f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr{)}\Bigr{)}^{2}\,\mathrm{d}\boldsymbol{x}_{v^{c}},$
 implying $V^{v}_{\lambda}(f)<\infty$ if almost every section $f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})$ has bounded variation.
@@
 Let $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$.
-Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})<\infty$ for $v\subseteq u\subseteq 1{:}s,u\neq\varnothing$ and $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ with $\alpha_{j}=\alpha\ \forall j\in v$.
-Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha_{j}$ if $j\in u\setminus v$ and $|\kappa_{j}|=\alpha+1$ if $j\in v$,
+Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})<\infty$ for nonempty $u\subseteq 1{:}s$, $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and a subset $v\subseteq\{j\mid\alpha_{j}=\alpha\}$.
+Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$,
 $\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}\leqslant 4^{|u|}\Big{(}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})\Big{)}^{2}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1},$
 where
-$B(\boldsymbol{k}_{u},v)=\{\boldsymbol{k}^{\prime}\in\mathbb{N}^{s}_{*}\mid\boldsymbol{s}(\boldsymbol{k}^{\prime})=u,\kappa^{\prime}_{j}=\kappa_{j}\text{if }j\in u\setminus v,\lceil\kappa^{\prime}_{j}\rceil_{1:(\alpha+1)}=\kappa_{j}\text{ if }j\in v\}.$ (13)
+$B(\boldsymbol{k}_{u},v)=\{\boldsymbol{k}^{\prime}\in\mathbb{N}^{s}_{*}\mid\boldsymbol{s}(\boldsymbol{k}^{\prime})=u,\ \lceil\kappa^{\prime}_{j}\rceil_{1:(\alpha+1)}=\kappa_{j}\ \forall j\in v,\ \kappa^{\prime}_{j}=\kappa_{j}\ \forall j\in u\setminus v\}.$ (13)
 Proof.
 Let $\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)$.
-Because $|\kappa^{\prime}_{j}|=|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$ and $|\kappa^{\prime}_{j}|\geqslant|\kappa_{j}|=\alpha+1$ for $j\in v$, we can apply Lemma 3 to get
+Because $|\kappa^{\prime}_{j}|\geqslant|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa^{\prime}_{j}|=|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$, we can apply Lemma 3 to get
 $|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=$
@@
 $g(\boldsymbol{x}_{v})=I_{v}\Big{(}f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}W_{\kappa_{j}}\Big{)}.$
-From the proof of [22, Theorem 3], we can map $v$ to $1{:}|v|$ and show
+From the proof of [25, Theorem 3], we can map $v$ to $1{:}|v|$ and show
 $\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}\leqslant 2^{-|v|}\Big{(}V^{(|v|)}_{\lambda}(g)\Big{)}^{2}\Bigl{(}\prod_{j\in v}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}\Bigr{)}.$ (14)
@@
 Motivated by Lemma 4, we define for $\alpha\in\mathbb{N}_{0},\lambda\in(0,1]$ and nonempty $u\subseteq 1{:}s$:
-$\|f\|_{u,\alpha,\lambda}=\sup_{\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}}\sup_{v\subseteq\{j\mid\alpha_{j}=\alpha\}}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u}),$ (15)
-where $f_{u}$ is the ANOVA component of $f$ defined in equation (8). We set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
+$\|f\|_{u,\alpha,\lambda}=\sup_{v\subseteq u}\sup_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u}),$ (15)
+where $f_{u}$ is the ANOVA component of $f$ defined in equation (8).
+Notice that we add the restriction $\alpha_{j}>0$ for $j\in u\setminus v$ because $k_{j}\in\mathbb{N}$ implies $\alpha_{j}=|\kappa_{j}|>0$ in the statement of Lemma 4.
+In particular, $\|f\|_{u,\alpha,\lambda}=V^{u}_{\lambda}(f_{u})$ when $\alpha=0$.
+We further set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
 $\|f\|_{\alpha,\lambda}=\sup_{u\subseteq 1{:}s}\|f\|_{u,\alpha,\lambda}.$
@@
 $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}=\sup_{u\subseteq 1{:}s}\tilde{\gamma}_{u}^{-1}\|f\|_{u,\alpha,\lambda}.$
-All subsequent results hold equivalently for $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ replacing $\|f\|_{\alpha,\lambda}$ and $\tilde{\gamma}_{u}$ replacing $\gamma_{u}$.
+All subsequent results hold equivalently with $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ replacing $\|f\|_{\alpha,\lambda}$ and $\tilde{\gamma}_{u}$ replacing $\gamma_{u}$.
 However, since $\|f\|_{\alpha,\lambda}$ and $\gamma_{u}$ are intrinsic to $f$, while $\|f\|_{\alpha,\lambda,\tilde{\gamma}_{u}}$ requires external weights $\tilde{\gamma}_{u}$ unused by our algorithm, we adopt the unweighted framework for simplicity.
@@
 4 Tractability under the completely random design
-In this section, we show the median RQMC estimator $\hat{\mu}^{(r)}_{\infty}$ under the completely random design achieves a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate given $\|f\|_{\alpha,\lambda}<\infty$, with no dimension-dependence if further $\gamma_{u}$ satisfies equation (34).
+In this section, we show the median RQMC estimator $\hat{\mu}^{(r)}_{\infty}$ under the completely random design achieves near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates given $\|f\|_{\alpha,\lambda}<\infty$, with no dimension-dependence if further $\gamma_{u}$ satisfies equation (34).
 First, we define $\|\kappa\|=\sum_{\ell\in\kappa}\ell$ for $\kappa\subseteq\mathbb{N}$, with $\|\varnothing\|=0$.
@@
 for $B(\boldsymbol{k}_{u},v)$ defined as in equation (13) and
-$\mathcal{K}_{u,v}(T)=\Big{\{}(k_{j},j\in u)\in\mathbb{N}^{|u|}\Big{|}\begin{matrix}|\kappa_{j}|\leqslant\alpha\ \forall j\in u\setminus v,|\kappa_{j}|=\alpha+1\ \forall j\in v,\\ (\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|\leqslant T\end{matrix}.\Big{\}}$
+$\mathcal{K}_{u,v}(T)=\Big{\{}(k_{j},j\in u)\in\mathbb{N}^{|u|}\Big{|}\begin{matrix}|\kappa_{j}|=\alpha+1\ \forall j\in v,\ |\kappa_{j}|\leqslant\alpha\ \forall j\in u\setminus v,\\ (\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|\leqslant T\end{matrix}\Big{\}}.$
 We note that $K_{u}(T)$ implicitly depends on $\alpha$ and $\lambda$.
-The next lemma quantifies how $T$ controls the size of $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$.
+The next lemma quantifies how $T$ controls the size of $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ for $\boldsymbol{k}\notin K_{u}(T)$.
 Lemma 5.
@@
 where
-$C_{\alpha,\lambda,\theta}=\frac{4^{-\alpha-1}}{\alpha!(4^{(\alpha+\lambda)(1-\theta)}-1)(4^{1-\theta}-1)^{\alpha}}+\sum_{\alpha^{\prime}=0}^{\alpha}\frac{4^{-\alpha^{\prime}}}{(\alpha^{\prime})!(4^{1-\theta}-1)^{\alpha^{\prime}}}.$ (18)
+$C_{\alpha,\lambda,\theta}=\frac{4^{-\alpha}}{\alpha!(4^{(\alpha+\lambda)(1-\theta)}-1)(4^{1-\theta}-1)^{\alpha}}+\sum_{\alpha^{\prime}=0}^{\alpha}\frac{4^{-\alpha^{\prime}+1}}{(\alpha^{\prime})!(4^{1-\theta}-1)^{\alpha^{\prime}}}.$ (18)
 Proof.
@@
 $g(t)=\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha^{\prime}},$
-and let $[t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$.
+and let $[t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [6] for background knowledge on generating functions).
 It follows that $[t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$.
 Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
-$\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\leqslant\frac{1}{(\alpha^{\prime})!}\Big{(}\sum_{N=1}^{\infty}\rho_{\theta}^{N}\Big{)}^{\alpha^{\prime}}=\frac{1}{(\alpha^{\prime})!(\rho_{\theta}^{-1}-1)^{\alpha^{\prime}}}.$ (20)
+$\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}t^{\ell^{\prime}}\leqslant\frac{1}{(\alpha^{\prime})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha^{\prime}}=\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}.$ (20)
+with $t=\rho_{\theta}$.
 For $\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})$, we let $\kappa^{+}=\lceil\kappa\rceil_{1{:}\alpha}$ and $\ell=\lceil\kappa\rceil_{\alpha+1}$.
@@
 Our conclusion follows once we put the above bounds into equation (19). ∎
-While larger $T$ implies a tighter bound on $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$, it also increases the probability $Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K_{u}(T)$.
+While larger $T$ implies a tighter bound on $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$, it also increases the probability that $Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K_{u}(T)$.
 The next lemma provides a bound on the cardinality of $K_{u}(T)$.
@@
 For $\alpha\in\mathbb{N}_{0},\lambda\in(0,1],T\in\mathbb{R}$ and nonempty $u\subseteq 1{:}s$,
-$|K_{u}(T)|\leqslant 2^{T/(\alpha+\lambda)}\Big{(}\max\big{(}T/(\alpha+\lambda),1\big{)}A_{\alpha,\lambda}+B_{\alpha,\lambda}\Big{)}^{|u|},$
+$|K_{u}(T)|\leqslant 2^{T/(\alpha+\lambda)}\Big{(}A_{\alpha,\lambda}\max\big{(}T/(\alpha+\lambda),1\big{)}+B_{\alpha,\lambda}\Big{)}^{|u|},$
 where
-$A_{\alpha,\lambda}=\frac{1}{\alpha!(2^{1/(\alpha+\lambda)}-1)^{\alpha}}\quad\text{and}\quad B_{\alpha,\lambda}=\sum_{\alpha^{\prime}=0}^{\alpha}\frac{1}{(\alpha^{\prime})!(2^{1/(\alpha+\lambda)}-1)^{\alpha^{\prime}}}.$ (21)
+$A_{\alpha,\lambda}=\frac{1}{\alpha!(2^{1/(\alpha+\lambda)}-1)^{\alpha}}\quad\text{and}\quad B_{\alpha,\lambda}=\sum_{\alpha^{\prime}=1}^{\alpha}\frac{1}{(\alpha^{\prime})!(2^{1/(\alpha+\lambda)}-1)^{\alpha^{\prime}}},$ (21)
+with the sum from $1$ to $\alpha$ interpreted as 00 when $\alpha=0$.
 Proof.
@@
 Thus, $|B(\boldsymbol{k}_{u},v)|=\prod_{j\in v}2^{\ell_{j}-1}$.
-Because $B(\boldsymbol{k}_{u},v)$ for distinct $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ are mutually disjoint,
+Because $B(\boldsymbol{k}_{u},v)$ are mutually disjoint for distinct $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$,
 $|K_{u}(T)|=\sum_{v\subseteq u}\sum_{\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)}\prod_{j\in v}2^{\ell_{j}-1}.$ (22)
@@
 For $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$, let
-$\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathbb{N}_{0}^{|u|}\Big{|}|\kappa_{j}|=\alpha_{j}\text{ for }j\in u,\sum_{j\in u}\|\kappa_{j}\|=N\Big{\}}.$
-We can decompose $\mathcal{K}_{u,v}(T)$ into subsets with the same $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ and $(\kappa^{+}_{j}-\ell_{j},j\in v;\kappa_{j},j\in u\setminus v)\in\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})$ for $N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}$, and rewrite equation (22) as
-$|K_{u}(T)|=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$ (23)
-$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|+$
-$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
+$\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathbb{N}_{0}^{|u|}\Big{|}|\kappa_{j}|=\alpha_{j}\ \forall j\in u,\sum_{j\in u}\|\kappa_{j}\|=N\Big{\}}.$
+We can decompose $\mathcal{K}_{u,v}(T)$ into subsets with the same $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ and $(\kappa^{+}_{j}-\ell_{j},j\in v;\kappa_{j},j\in u\setminus v)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ for $N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}$, and rewrite equation (22) as
+$|K_{u}(T)|=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$ (23)
+$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|+$
+$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
 Let $\overline{N}=\lfloor(T-N)/(\alpha+\lambda)\rfloor$.
@@
 Putting the above bound into equation (23), we get
-$|K_{u}(T)|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|$ (25)
-$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
-$\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}.$
+$|K_{u}(T)|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|$ (25)
+$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
+$\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}.$
 To bound the sum over $N$, consider the generating function
 $g(t)=\prod_{j\in u}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}.$
-Because each element of $\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})$ is counted $\prod_{j\in u}(\alpha_{j})!$ times by $t^{N}[g(t)]$ , an argument similar to equation (20) gives
-$\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|t^{N}\leqslant\prod_{j\in u}\frac{1}{(\alpha_{j})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}=\prod_{j\in u}\frac{1}{(\alpha_{j})!(t^{-1}-1)^{\alpha_{j}}}$ (26)
+Because each element of $\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ is counted $\prod_{j\in u}(\alpha_{j})!$ times by $t^{N}[g(t)]$ , an argument similar to equation (20) gives
+$\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|t^{N}\leqslant\prod_{j\in u}\frac{1}{(\alpha_{j})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}=\prod_{j\in u}\frac{1}{(\alpha_{j})!(t^{-1}-1)^{\alpha_{j}}}$ (26)
 for $t=2^{-1/(\alpha+\lambda)}$ and
-$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\prod_{j\in u}\frac{1}{(\alpha_{j})!(2^{1/(\alpha+\lambda)}-1)^{\alpha_{j}}}$
-$=$ $A_{\alpha,\lambda}^{|v|}B^{|u|-|v|}_{\alpha,\lambda},$
+$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\prod_{j\in u}\frac{1}{(\alpha_{j})!(2^{1/(\alpha+\lambda)}-1)^{\alpha_{j}}}$
+$=$ $A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|},$
 where $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are defined in equation (21). By a similar calculation,
-$2^{-T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant B^{|u|}_{\alpha,\lambda}.$
+$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
+$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}.$
 Putting the above bounds into equation (25), we finally get
-$|K_{u}(T)|\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}\Big{(}1+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}A_{\alpha,\lambda}^{|v|}B^{-|v|}_{\alpha,\lambda}\Big{)}$ (27)
-$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}\Big{(}1+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}\max\big{(}T/(\alpha+\lambda),1\big{)}^{|v|}A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{-|v|}\Big{)}$
+$|K_{u}(T)|\leqslant$ $2^{T/(\alpha+\lambda)}\Big{(}B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}A_{\alpha,\lambda}^{|v|}B^{|u|-|v|}_{\alpha,\lambda}\Big{)}$ (27)
+$\leqslant$ $2^{T/(\alpha+\lambda)}\Big{(}B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}\max\big{(}T/(\alpha+\lambda),1\big{)}^{|v|}A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|}\Big{)}$
 $=$ $2^{T/(\alpha+\lambda)}\Big{(}B_{\alpha,\lambda}+\max\big{(}T/(\alpha+\lambda),1\big{)}A_{\alpha,\lambda}\Big{)}^{|u|}.\qed$ (28)
@@
 For $m\in\mathbb{N}$ and $\gamma_{u}$ defined by equation (16), let
-$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}.$ (29)
+$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}.$ (29)
 Then for any $\delta\in(0,1)$,
@@
 Lemma 6 then implies
-$\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}|K_{u}(T_{u,m})|\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{T_{u,m}/(\alpha+\lambda)}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
-$=$ $\frac{\delta 2^{m}}{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
+$\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}|K_{u}(T_{u,m})|\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{T_{u,m}/(\alpha+\lambda)}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}$
+$=$ $\frac{\delta 2^{m}}{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}$
 $\leqslant$ $\delta 2^{m}.$
 Under the completely random design, $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for any $\boldsymbol{k}\neq\boldsymbol{0}$, so our conclusion follows by taking a union bound. ∎
-Now we are ready to show the following error bounds on $\hat{\mu}^{(r)}_{\infty}$:
+Now we are ready to prove the following error bounds on $\hat{\mu}^{(r)}_{\infty}$:
 Theorem 1.
@@
 Let $\delta=1/16$ in Lemma 2 and Corollary 1. By Lemma 2, equation (5) and $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for $\boldsymbol{k}\neq\boldsymbol{0}$, it suffices to show
-$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}\leqslant\Gamma_{\alpha,\lambda,\gamma_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2\theta(\alpha+\lambda)}}{2^{2\theta(\alpha+\lambda)(m-4)}}$
+$2^{-m+4}\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}\leqslant\Gamma_{\alpha,\lambda,\gamma_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2\theta(\alpha+\lambda)}}{2^{(2\theta(\alpha+\lambda)+1)m}}$
 for $K$ defined by equation (17) with $T_{u,m}$ given by equation (30). From equation (9) and Lemma 5,
 $\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}=\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$ (32)
-Our conclusion follows after plugging in $\|f\|^{2}_{u,\alpha,\lambda}=\gamma^{2}_{u}\|f\|^{2}_{\alpha,\lambda}$ and equation (30). ∎
+Our conclusion follows by plugging in $\|f\|^{2}_{u,\alpha,\lambda}=\gamma^{2}_{u}\|f\|^{2}_{\alpha,\lambda}$ and equation (30). ∎
 Remark 2.
 Given a failure probability tolerance $\varepsilon>0$, we can set $r=\lceil\log_{2}(\varepsilon^{-1})\rceil$ to ensure $2^{-r}\leqslant\varepsilon$.
-For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, $N=r2^{m}$ is the total number of function evaluations.
+For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, where $N=r2^{m}$ is the total number of function evaluations.
 To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$, so $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$.
-Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}\leqslant 1$ that
+Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}\leqslant 1$,
 $\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big{)}\leqslant 2^{-r}\leqslant\varepsilon,$ (33)
-where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is almost optimal [22, Subsection 3.2].
+where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is almost optimal [25, Subsection 3.2].
 Remark 3.
@@
 where we have used $1+x\leqslant\exp(x)$. Similarly,
-$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=$ $1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}$
-$\leqslant$ $\prod_{j=1}^{s}\Big{(}1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})\Big{)},$
-which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ and a constant $C^{\prime}_{\eta}$ independent of $s$ [10, Lemma 3].
-Substituting into equation (31), equation (33) holds with a constant $C_{\eta}$ independent of $s$, fulfilling the definition of strong tractability [20].
-Notably, equation (34) resembles the tractability conditions assumed by [8, Corollary 3.13] for polynomial lattices to achieve a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate.
+$\Gamma_{\alpha,\lambda,\gamma_{u}}(m)=$ $1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}$
+$\leqslant$ $\prod_{j=1}^{s}\Big{(}1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})\Big{)},$
+which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [11, Lemma 3].
+Putting the above bounds into equation (31), we conclude that equation (33) holds with a constant $C_{\eta}$ independent of $s$.
+Notably, equation (34) resembles the tractability conditions assumed by [9, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
 5 Tractability under more general randomization
@@
 $\Pr(Z(\boldsymbol{k})=1)\leqslant 2^{-m+\tilde{t}_{u}}\boldsymbol{1}\Big{\{}\sum_{j\in u}\|\kappa_{j}\|_{(d)}>d(m-\tilde{t}_{u})\Big{\}}.$
-From [8, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [8, Section 2].
+From [9, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [9, Section 2].
 We frame this assumption to align with broader applications in higher-order digital nets [4], where the parameter $d$ (the order of digital nets) may exceed $1$.
@@
 However, this refinement offers limited practical improvement: Lemma 6 suggests $|K_{u}(T)|$ grows exponentially in $T$, so the margin $T_{u,m}-(\alpha+\lambda)(m-\tilde{t}_{u})$ is at best $O(\alpha+\lambda)$.
-Consequently, Consequently, the error bound improves by only a factor of $2^{O(\alpha+\lambda)}$, leaving the asymptotic convergence rate unchanged.
+Consequently, the error bound improves by only a factor of $2^{O(\alpha+\lambda)}$, leaving the asymptotic convergence rate unchanged.
 Next, we consider the $d<\alpha+\lambda$ case. We first establish the counterpart of Lemma 6.
@@
 $|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$
-$\leqslant$ $2^{(T-T^{\prime})/\beta}\Big{(}\max\big{(}(T-T^{\prime})/\beta,1\big{)}A_{\alpha,\lambda}+B_{\alpha,\lambda}\Big{)}^{|u|}+2^{(T-\theta^{\prime}T^{\prime})/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}},$
-where $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are given by equation (21), $\beta=\alpha+\lambda-d$ and
-$D_{\alpha,\lambda,d,\theta^{\prime}}=\frac{1}{d!(2^{(1-\theta^{\prime})/\beta}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(2^{1/\beta}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(2^{(1-\theta^{\prime})/\beta}-1)^{\alpha^{\prime}}}.$
+$\leqslant$ $2^{(T-T^{\prime})/\beta}\Big{(}A_{\alpha,\lambda}\max\big{(}(T-T^{\prime})/\beta,1\big{)}+B_{\alpha,\lambda}\Big{)}^{|u|}+2^{(T-\theta^{\prime}T^{\prime})/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}},$
+where $\beta=\alpha+\lambda-d$, $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are given by equation (21), and
+$D_{\alpha,\lambda,d,\theta^{\prime}}=\frac{1}{d!(2^{(1-\theta^{\prime})/\beta}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(2^{1/\beta}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(2^{(1-\theta^{\prime})/\beta}-1)^{\alpha^{\prime}}},$ (35)
+with the sum from $1$ to $\alpha-d$ interpreted as 00 when $\alpha=d$.
 Proof.
@@
 Let
-$\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})\Big{|}\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N^{\prime}\Big{\}}.$
+$\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})\Big{|}\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N^{\prime}\Big{\}}.$
 Similar to equation (23), we can write
 $|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$ (36)
-$=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha,\forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}-d\|\boldsymbol{\ell}_{v}\|_{1}<N^{\prime}\leqslant N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$
-$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|+$
-$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
+$=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}-d\|\boldsymbol{\ell}_{v}\|_{1}<N^{\prime}\leqslant N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$
+$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|+$
+$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
 The sum over $\boldsymbol{\ell}_{v}$ is empty if $(T^{\prime}-N^{\prime})/d\geqslant(T-N)/(\alpha+\lambda)$.
@@
 Therefore,
-$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
-$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}\sum_{N^{\prime}=0}^{N}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
-$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1},$
+$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
+$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}\sum_{N^{\prime}=0}^{N}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
+$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1},$
 where we have applied equation (24) in the last step. Letting $T_{*}=(T-N_{*})/(\alpha+\lambda)$, we put the above bound into equation (36) and get
-$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (36)
-$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
-$\leqslant$ $2^{T_{*}}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T_{*}\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v\end{subarray}}\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
+$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (37)
+$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
+$\leqslant$ $2^{T_{*}}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T_{*}\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
 $\leqslant$ $2^{T_{*}}\Big{(}B_{\alpha,\lambda}+\max(T_{*},1)A_{\alpha,\lambda}\Big{)}^{|u|},$
-where the last step follows from a calculation similar to equation (27).
+where the last step follows from a calculation similar to equation (25) and (27).
 Next, we notice
-$t^{T}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}t^{N^{\prime}}\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (37)
+$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant t^{-T}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}t^{N^{\prime}}\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (38)
 for $t=2^{-1/(\alpha+\lambda-d)}$.
 To bound the sum over $N$, we let $w=\{j\in u\mid\alpha_{j}>d\}$ for $\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and define $\boldsymbol{\alpha}^{-}_{u}=(\min(\alpha_{j},d),j\in u)$ and $\boldsymbol{\alpha}^{+}_{u}=(\alpha_{j}-d,j\in w)$ if $w\neq\varnothing$.
-For $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$, we further let $\kappa^{-}_{j,d}=\lceil\kappa_{j}\rceil_{1{:}\min(\alpha_{j},d)}$ for $j\in u$ and $\kappa^{+}_{j,d}=\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}d}$ for $j\in w$.
-Because for $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$,
+For $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$, we further let $\kappa^{-}_{j,d}=\lceil\kappa_{j}\rceil_{1{:}\min(\alpha_{j},d)}$ for $j\in u$ and $\kappa^{+}_{j,d}=\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}d}$ for $j\in w$.
+Because for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$,
 $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=\sum_{j\in w}\|\kappa_{j}\|-\sum_{j\in w}\|\kappa^{-}_{j,d}\|=\sum_{j\in u}\|\kappa_{j}\|-\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N-N^{\prime},$
-the number of possible $(\kappa^{+}_{j,d},j\in w)$ for $\boldsymbol{k}_{u}\in\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$ with a given set of $(\kappa^{-}_{j,d},j\in u)\in\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$ is bounded by the number of $(\kappa^{+}_{j,d},j\in w)$ with $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=N-N^{\prime}$, which is $|\mathcal{K}^{\prime}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ by definition.
-By summing this bound over all elements of $\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$, we get
-$\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=N^{\prime}}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|t^{N-N^{\prime}}$
-$\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=0}^{\infty}|\mathcal{K}^{\prime}_{w}(N,\boldsymbol{\alpha}^{+}_{u})|t^{N}$
-$\leqslant$ $|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}$
+the number of possible $(\kappa^{+}_{j,d},j\in w)$ for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$ with a given set of $(\kappa^{-}_{j,d},j\in u)\in\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$ is bounded by the number of $(\kappa^{+}_{j,d},j\in w)$ with $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=N-N^{\prime}$, which is $|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ by definition.
+This implies $|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\times|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ and
+$\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=N^{\prime}}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|t^{N-N^{\prime}}$
+$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{w}(N,\boldsymbol{\alpha}^{+}_{u})|t^{N}$
+$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}},$
 where we have applied equation (26) in the last step. Putting the above bound into equation (38), we get for any $\theta^{\prime}\in(0,1)$,
-$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\mathcal{K}^{\prime}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$
-$\leqslant$ $t^{-T}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{N^{\prime}}$
-$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=0}^{\infty}|\mathcal{K}^{\prime}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{(1-\theta^{\prime})N^{\prime}}$
+$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$
+$\leqslant$ $t^{-T}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{N^{\prime}}$
+$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{(1-\theta^{\prime})N^{\prime}}$
 $\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
-where we have again applied equation (26). Our conclusion finally follows from equation (36), $t=2^{-1/(\alpha+\lambda-d)}$ and
-$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\end{subarray}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
-$=$ $\sum_{w\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{-}_{u}\in\mathbb{Z}^{|u|}_{\leqslant d}\\ w=\{j\mid\alpha^{-}_{j}=d\}\end{subarray}}\Big{(}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}^{|u|}$
-$=$ $\sum_{w\subseteq u}\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|-|w|}$
-$=$ $\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=0}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=0}^{d-1}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|}.\qed$
+where we have applied equation (26) once more. Our conclusion finally follows from equation (37), $t=2^{-1/(\alpha+\lambda-d)}$ and
+$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
+$=$ $\sum_{w\subseteq u}\Big{(}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{-}_{u}\in\mathbb{Z}^{|u|}_{\leqslant d}\\ \alpha_{j}=d\ \forall j\in w\\ \alpha^{-}_{j}>0\ \forall j\in u\setminus w\end{subarray}}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}^{|u|}$
+$=$ $\sum_{w\subseteq u}\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|-|w|}$
+$=$ $D_{\alpha,\lambda,d,\theta^{\prime}}^{|u|},$
+where $D_{\alpha,\lambda,d,\theta^{\prime}}$ is defined in equation (35). ∎
 Corollary 2.
 For $m\in\mathbb{N}$ and $\gamma_{u}$ defined by equation (16), let
-$\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$ (38)
+$\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$ (39)
 Then for any $\delta\in(0,1)$,
@@
 $|K_{u}(T_{u,m})\cap K^{\prime}_{u}(T^{\prime}_{u,m})|$
-$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}2^{m-\tilde{t}_{u}}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{(1-\theta^{\prime})dm/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
+$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}2^{m-\tilde{t}_{u}}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
 A union bound over all $\boldsymbol{k}\in K_{u}(T_{u,m})\cap K^{\prime}_{u}(T^{\prime}_{u,m})$ implies
 $\Pr(Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in K_{u}(T_{u,m}))$
-$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{(1-\theta^{\prime})dm/\beta}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
+$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
 On the other hand, when $T_{u,m}<T^{\prime}_{u,m}$, every $\boldsymbol{k}\in K_{u}(T_{u,m})$ satisfies
@@
 Let $\delta=1/16$ in Lemma 2 and Corollary 2. By Lemma 5 and Assumption 1,
-$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}\Pr(Z(\boldsymbol{k})=1)|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant 2^{-m+\tilde{t}_{u}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}$
-for $T_{u,m}$ defined by equation (39). From Equation (9) and (16),
-$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}\leqslant\|f\|^{2}_{\alpha,\lambda}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\frac{\gamma^{2}_{u}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$ (40)
-Our conclusion finally follows from Lemma 2 and equation (5). ∎
+$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}\Pr(Z(\boldsymbol{k})=1)|\hat{f}_{u}(\boldsymbol{k})|^{2}\leqslant 2^{-m+\tilde{t}_{u}}\frac{\|f\|^{2}_{u,\alpha,\lambda}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$
+From Equation (9) and (16),
+$\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}=$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T_{u,m})}\Pr(Z(\boldsymbol{k})=1)|\hat{f}_{u}(\boldsymbol{k})|^{2}$ (41)
+$\leqslant$ $\|f\|^{2}_{\alpha,\lambda}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\frac{\gamma^{2}_{u}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$
+Our conclusion follows from Lemma 2 and equation (5) after plugging in $T_{u,m}$ from equation (40). ∎
 Remark 5.
@@
 We assume equation (34) holds for weights $\{\gamma_{j},j\in 1{:}s\}$.
-Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for $C_{\eta}$ independent of $s$.
+Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C_{\eta}$ independent of $s$.
 Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
@@
 for an absolute constant $C$.
-For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for a nonnegative sequence $\{\tilde{t}_{j},j\in 1{:}s\}$.
-We can then bound
+For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for nonnegative numbers $\{\tilde{t}_{j},j\in 1{:}s\}$. We can then bound
 $\frac{\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}}{16^{2\theta(\alpha+\lambda-d)+1}}\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\prod_{j\in u}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}C_{\alpha,\lambda,\theta}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}$
@@
 Thus, prudent selection of $d$ is critical for practical implementations.
-When $d<\alpha+\lambda$, we can refine our argument to establish strong tractability under a weaker condition analogous to that outlined in Remark 7.
+When $d<\alpha+\lambda$, we can refine our argument to establish dimension-independent convergence under a weaker condition analogous to that outlined in Remark 7.
 Theorem 4.
@@
 $\sup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\leqslant\sup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\prod_{j\in u}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\leqslant\Lambda$
-for a constant $\Lambda$ independent of $s$. Let $\delta^{\prime}=\min(\Lambda^{-\phi},1/16)$, $\theta^{\prime}=1-\eta d,$
-$\Gamma(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}$
+for a constant $\Lambda$ independent of $s$. Let $\delta^{\prime}=\min(\Lambda^{-\phi},1/16)$, $\theta^{\prime}=1-\eta/d,$
+$\Gamma(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}$
 and
@@
 $\log_{2}(\delta^{\prime})+\frac{\phi\log_{2}(\gamma_{u})}{\alpha+\lambda+1/2}+\phi\tilde{t}_{u}=\log_{2}(\delta^{\prime})+\phi\log_{2}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}\leqslant 0.$
-Therefore, $T_{u,m}\leqslant(\alpha+\lambda)(m-\tilde{t}_{u})$ and an argument similar to Corollary 2 shows
+Therefore, $T_{u,m}\leqslant(\alpha+\lambda)(m-\tilde{t}_{u})$ and an argument similar to that used in Corollary 2 shows
 $\Pr\Big{(}Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in\bigcup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}K_{u}(T_{u,m})\Big{)}$
-$\leqslant$ $\frac{\delta^{\prime}}{\Gamma(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(mA_{\alpha,\lambda}+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)},$
+$\leqslant$ $\frac{\delta^{\prime}}{\Gamma(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)},$
 which is bounded by $\delta^{\prime}$ due to the definition of $\Gamma(m)$. Equation (41) with the above choice of $T_{u,m}$ then implies
@@
 where
-$\Gamma^{\prime}=\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{2-\frac{2\phi\theta(\alpha+\lambda-d)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}\big{(}1+2\theta(\alpha+\lambda)-2\theta\phi(\alpha+\lambda-d)\big{)}}.$
-Like in Remark 6, we can show $\Gamma(m)\leqslant C^{\prime}_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C^{\prime}_{\eta}$ independent of $s$ when equation (42) holds.
+$\Gamma^{\prime}=\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{2-\frac{2\theta\phi(\alpha+\lambda-d)}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}2^{\tilde{t}_{u}\big{(}1+2\theta(\alpha+\lambda)-2\theta\phi(\alpha+\lambda-d)\big{)}}.$
+Analogous to Remark 6, we can show $\Gamma(m)\leqslant C^{\prime}_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C^{\prime}_{\eta}$ independent of $s$ when equation (43) holds.
 Furthermore, we notice that $\alpha+\lambda>\phi(\alpha+\lambda-d)$ and
@@
 $=$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}C^{|u|}_{\alpha,\lambda,\theta},$
-which is bounded as $s\to\infty$ given equation (42). Our conclusion then follows from Lemma 2 with $\delta=1/16$ and equation (5). ∎
+which is bounded as $s\to\infty$ because of equation (43). Our conclusion then follows from Lemma 2 with $\delta=1/16$ and equation (5). ∎
 Remark 8.
@@
 Remark 9.
-For random linear scrambling using base-2 Niederreiter sequences, substituting the bound (41) into equation (42) yields the simplified tractability condition:
+For the random linear scrambling using base-2 Niederreiter sequences, substituting the bound (42) into equation (43) yields the simplified tractability condition:
 $\lim_{s\to\infty}\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}j\log_{2}(j)\Big{)}^{\phi}<\infty.$
-Since $\phi>1$, this condition is strictly weaker than the requirements in [8, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
-In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, strong tractability holds if
+Since $\phi>1$, this condition is strictly weaker than the requirements in [9, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
+In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, $\mu^{(r)}_{\infty}$ attains dimension-independent convergence if
 $\phi(\frac{p}{\alpha+\lambda+1/2}-q)=\frac{p-(\alpha+\lambda+1/2)q}{\alpha+\lambda-d+1/2}>1.$ (44)
@@
 6 Numerical experiments
-In this section, we test our theory across integrands with varying dimension and decay on $\gamma_{u}$. The test integrands have the common form
+In this section, we test our theory across integrands with varying dimensions and decay rates on $\gamma_{u}$. The test integrands have the common form
 $f_{s,\gamma,\alpha}(\boldsymbol{x})=\prod_{j=1}^{s}\Big{(}1+\frac{c_{\alpha}}{j^{\gamma}}(x_{j}\exp(x_{j})-1)\Big{)},$
-where $s=10,100,1000$ controls the dimension of integration domain, $\gamma=2,3,4$ controls the decay of $\gamma_{u}$ and $c_{\alpha}$ for $\alpha=0,1$ is a normalizing constant to ensure $\|f\|_{\alpha,1}=1$.
+where $s=10,100,1000$ controls the dimension of integration domain, $\gamma=2,3,4$ controls the decay rate of $\gamma_{u}$ and $c_{\alpha}$ for $\alpha=0,1$ is a normalizing constant to ensure $\|f\|_{\alpha,1}=1$.
 To determine $c_{\alpha}$, we start by computing the ANOVA component given by equation (8):
@@
 $\Big{(}\int_{[0,1]^{d}}\Big{(}f^{(1,\dots,1)}(\boldsymbol{x})\Big{)}^{2}\,\mathrm{d}\boldsymbol{x}\Big{)}^{1/2}.$
-Hence,
-$V^{(|u|)}_{1}(f_{u})=\|f^{(1,\dots,1)}_{u}\|_{L^{2}}=\frac{c^{|u|}_{\alpha}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}.$
-By equation (12), for $v\subseteq u$,
-$\Big{(}V^{v}_{\lambda}(f_{u})\Big{)}^{2}$
-$\leqslant$ $\frac{c^{2|u|}_{\alpha}}{(\prod_{j\in u}j)^{2\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|v|}\Big{(}\int_{0}^{1}(x\exp(x)-1)^{2}\,\mathrm{d}x\Big{)}^{|u|-|v|}$
-$\leqslant$ $\Big{(}V^{|u|}_{\lambda}(f_{u})\Big{)}^{2},$
-while $V^{u}_{\lambda}(f_{u})=V^{|u|}_{\lambda}(f_{u})$. Therefore,
-$\|f\|_{u,0,1}=\sup_{v\subseteq u}V^{v}_{1}(f_{u})=V^{u}_{1}(f_{u})=\frac{c^{|u|}_{\alpha}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}$
+Hence, $V^{(|u|)}_{1}(f_{u})=\|f^{(1,\dots,1)}_{u}\|_{L^{2}}$. As discussed in Section 3, $\|f\|_{u,\alpha,\lambda}=V^{u}_{\lambda}(f_{u})$ when $\alpha=0$. Therefore,
+$\|f\|_{u,0,1}=\|f^{(1,\dots,1)}_{u}\|_{L^{2}}=\frac{c^{|u|}_{0}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}$
 and we can make $\|f_{s,\gamma,0}\|_{0,1}=1$ by setting
 $c_{0}=\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{-1/2}.$
-A similar calculation shows $\|f_{s,\gamma,1}\|_{1,1}=1$ if we set
+Next, when $\alpha=1$, it follows from the definition (16) that
+$\|f\|_{u,1,1}=\sup_{v\subseteq u}\sup_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant 1}^{|u|}\\ \alpha_{j}=1\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}V^{v}_{1}(f^{(\boldsymbol{\alpha}_{u})}_{u})=\sup_{v\subseteq u}V^{v}_{1}(f^{(1,\dots,1)}_{u}).$
+Let $g=f^{(1,\dots,1)}_{u}.$ For each $v\subseteq u$, because $((x+1)\exp(x))^{\prime}=(x+2)\exp(x)$,
+$V^{(|v|)}_{1}\Bigl{(}g(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr{)}=\frac{c^{|u|}_{1}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|v|/2}\prod_{j\in u\setminus v}(x_{j}+1)\exp(x_{j}).$
+Therefore, by equation (12),
+$\Big{(}V^{v}_{1}(g)\Big{)}^{2}$
+$\leqslant$ $\frac{c^{2|u|}_{1}}{(\prod_{j\in u}j)^{2\gamma}}\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|v|}\Big{(}\int_{0}^{1}(x+1)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|-|v|}$
+$\leqslant$ $\|f^{(2,\dots,2)}_{u}\|_{L^{2}}^{2},$
+while $V^{u}_{1}(g)=\|g^{(1,\dots,1)}\|_{L^{2}}=\|f^{(2,\dots,2)}_{u}\|_{L^{2}}$. Therefore,
+$\|f\|_{u,1,1}=\|f^{(2,\dots,2)}_{u}\|_{L^{2}}=\frac{c^{|u|}_{1}}{(\prod_{j\in u}j)^{\gamma}}\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{|u|/2}$
+and we can make $\|f_{s,\gamma,1}\|_{1,1}=1$ by setting
 $c_{1}=\Big{(}\int_{0}^{1}(x+2)^{2}\exp(2x)\,\mathrm{d}x\Big{)}^{-1/2}.$
-With the above choice of $c_{\alpha}$, it also follows that $\gamma_{u}=\prod_{j\in}\gamma_{j}$ for $\gamma_{j}=j^{-\gamma}$.
+With the above choice of $c_{\alpha}$ for $\alpha=0,1$, we also have $\gamma_{u}=\prod_{j\in}\gamma_{j}$ for $\gamma_{j}=j^{-\gamma}$.
 For each integrand $f_{s,\gamma,\alpha}$, we evaluate the performance of the median estimator $\hat{\mu}^{(r)}_{E}$ for $m=1,2,\dots,16$ under two randomization schemes: the completely random design (CRD) and the random linear scrambling (RLS).
@@
 We set $r=10$, limiting the failure probability per data point to below $0.1\%$.
-As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the average of the $12r-$ 1 $\hat{\mu}_{E}$ realizations used by RLS.
+As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the $12r-$ 1 $\hat{\mu}_{E}$ realizations used by RLS.
 The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [12].
@@
 Refer to caption Refer to caption Refer to caption
-Figure 1: Integration error for $f_{s,\gamma,\alpha}$ with $\alpha=0$.
+Figure 1: Integration errors for $f_{s,\gamma,\alpha}$ with $\alpha=0$.
 Refer to caption Refer to caption Refer to caption
@@
 Refer to caption Refer to caption Refer to caption
-Figure 2: Integration error for $f_{s,\gamma,\alpha}$ with $\alpha=1$.
+Figure 2: Integration errors for $f_{s,\gamma,\alpha}$ with $\alpha=1$.
 Figure 1 and Figure 2 summarize the simulation results.
@@
 Furthermore, because $\|f_{s,\gamma,0}\|_{1,1}\geqslant 1$, the relative variation $\gamma^{\prime}_{u}$ of $f_{s,\gamma,0}$ under the $\|\cdot\|_{1,1}$ norm satisfies
-$\gamma^{\prime}_{u}=\frac{\|f_{s,\gamma,0}\|_{1,1,u}}{\|f_{s,\gamma,0}\|_{1,1}}\leqslant\prod_{j\in u}\Big{(}\frac{c_{0}}{c_{1}}j^{-\gamma}\Big{)}=\prod_{j\in u}\gamma^{\prime}_{j}$
+$\gamma^{\prime}_{u}=\frac{\|f_{s,\gamma,0}\|_{u,1,1}}{\|f_{s,\gamma,0}\|_{1,1}}\leqslant\prod_{j\in u}\Big{(}\frac{c_{0}}{c_{1}}j^{-\gamma}\Big{)}=\prod_{j\in u}\gamma^{\prime}_{j}$
 for $\gamma^{\prime}_{j}=(c_{0}/c_{1})j^{-\gamma}$.
-In general, this analysis extends to arbitrary norms $\|\cdot\|_{\alpha^{\prime},\lambda^{\prime}}$, $\alpha^{\prime}\in\mathbb{N},\lambda^{\prime}\in(0,1]$, and the corresponding $\gamma_{j}$ is always proportional to $j^{-\gamma}$.
+In general, this analysis extends to arbitrary norms $\|\cdot\|_{\alpha^{\prime},\lambda^{\prime}}$ with $\alpha^{\prime}\in\mathbb{N}_{0},\lambda^{\prime}\in(0,1]$, and the corresponding $\gamma_{j}$ is always proportional to $j^{-\gamma}$.
 Thus, the decay rate $\gamma$ dominates convergence behavior as $s\to\infty$, irrespective of $\alpha$.
 Next, we observe that for a fixed value of $\gamma$, the integration errors exhibit similar behavior as $s$ varies, aligning with our analysis of dimension-independent convergence.
-Applying equation (43) with parameters $d=1$, $p=\gamma$ and $q=1+\eta$ for any $\eta>0$, our theory predicts that for large $s$, $\hat{\mu}^{(r)}_{\infty}$ under RLS achieve a near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rate, provided $\alpha+\lambda+1/2<(\gamma+1)/2$.
+Applying equation (44) with parameters $d=1$, $p=\gamma$ and $q=1+\eta$ for an arbitrarily small $\eta>0$, our theory predicts that for large $s$, $\hat{\mu}^{(r)}_{\infty}$ under RLS achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates, provided $\alpha+\lambda+1/2<(\gamma+1)/2$.
 Empirically, the observed rates approximate $O(2^{-(3/2)m})$ when $\gamma=2$ and $O(2^{-2m})$ when $\gamma=3$.
@@
 This discrepancy is likely due to the limited range of $m$ tested.
-Notably, while our theory suggests that $\hat{\mu}^{(r)}_{\infty}$ under CRD can attain a near $O(2^{-(\alpha+\lambda+1/2)m})$ rate for $\alpha+\lambda+1/2<\gamma$, practical results show that CRD consistently underperforms RLS.
+Notably, while our theory suggests that $\hat{\mu}^{(r)}_{\infty}$ under CRD can attain near $O(2^{-(\alpha+\lambda+1/2)m})$ rates for $\alpha+\lambda+1/2<\gamma$, practical results show that CRD consistently underperforms RLS.
 We hypothesize that the superior theoretical rates for CRD may only manifest at substantially larger $m$.
 Finally, we compare the integration errors under RLS and STD.
-Across the vast majority of test cases, RLS demonstrates superior performance over STD, reinforcing that the median RQMC approach (via RLS) is a more robust alternative to traditional RQMC (via STD) in high-dimensional settings.
-We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m}$ rate derived from the root mean squared error when $\gamma=3$ or $4$.
+Across the vast majority of test cases, RLS demonstrates superior performance over STD, reinforcing that the median RQMC approach (via RLS) is a more robust alternative to traditional RQMC (via STD).
+We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m})$ rate derived from the root mean squared error when $\gamma=3$ or $4$.
 This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 2 and 3.

### v3

@@
 Recent advances in quasi-Monte Carlo integration demonstrate that the median of linearly scrambled digital net estimators achieves near-optimal convergence rates for high-dimensional integrals without requiring a priori knowledge of the integrand’s smoothness.
-Building on this framework, we prove that the median estimator attains dimension-independent convergence under tractability conditions characterized by low effective dimensionality, a property known as strong tractability in complexity theory.
-Our analysis strengthens existing guarantees by improving the convergence rates and relaxing the theoretical assumptions previously required for dimension-independent convergence.
+Building on this framework, we prove that the median estimator attains dimension-independent convergence, a property known as strong tractability in complexity theory, under tractability conditions characterized by low effective dimensionality.
+Using a probabilistic, integrand-specific error criterion, our analysis establishes both faster and dimension-independent convergence under weaker assumptions than previously possible in the worst-case setting.
 1 Introduction
@@
 This structured approach allows QMC to avoid the exponential cost growth with dimension of classical quadrature rules while achieving faster error convergence than MC.
-These advantages have made QMC particularly valuable in fields like computational finance [13, 18] and physics [16], where high-dimensional integrals are pervasive.
+These advantages have made QMC particularly valuable in fields like computational finance [joy1996quasi, l2009quasi] and physics [kuo:nuye:2016], where high-dimensional integrals are pervasive.
 Building on these strengths, we focus on randomized quasi-Monte Carlo (RQMC) methods based on linearly scrambled base-2 digital nets.
-Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [5].
-The random linear scrambling [21], a computationally efficient alternative to Owen’s scrambling [24], preserves the structure of digital nets while introducing controlled randomness.
+Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [dick:pill:2010].
+The random linear scrambling [mato:1998:2], a computationally efficient alternative to Owen’s scrambling [rtms], preserves the structure of digital nets while introducing controlled randomness.
 This randomization enhances the root mean squared error convergence rate to nearly $O(n^{-3/2})$ under smoothness conditions and enables statistical inference.
-However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [26].
-To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [2, 7, 8, 9, 32].
-The median of linearly scrambled digital net estimators was first analyzed in [26], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$.
-Subsequent work in [27] extended this result to higher dimensions.
-For finitely differentiable integrands, [25] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
+However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [superpolyone].
+To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [chen2025randomintegrationalgorithmhighdimensional, goda2024simpleuniversalalgorithmhighdimensional, goda:lecu:2022, Goda2024, ye2025medianqmcmethodunbounded].
+The median of linearly scrambled digital net estimators was first analyzed in [superpolyone], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$.
+Subsequent work in [superpolymulti] extended this result to higher dimensions.
+For finitely differentiable integrands, [pan2024automatic] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
 A critical gap in these prior studies is the role of dimensionality.
-We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [17].
-Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [23], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3).
-This notion of relative variation aligns with the effective dimension framework introduced in [1], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions.
-The connection to weighted function spaces [28] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
-Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [9].
-A key distinction lies in the randomization: their framework omits digital shifts (Subsection 2.1), a critical component of our scrambling method, and employs worst-case integration error over the unit ball in a weighted Sobolev space as the primary metric.
-In contrast, we demonstrate that median estimators incorporating digital shifts achieve near-optimal convergence rates on a fixed integrand with high probability, avoiding worst-case error analysis entirely.
-This probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [14, 15].
-Within this framework, we improve the convergence rates from nearly $O(n^{-\alpha})$ in [9] to nearly $O(n^{-\alpha-1/2})$, where $\alpha$ denotes the smoothness parameter in their analysis (we note that $\alpha$ is defined differently in this paper).
-Furthermore, we relax the conditions for dimension-independent convergence (Remarks 3 and 9).
-Specifically, we show digital nets under the completely random design (Subsection 2.1) achieve dimension-independent convergence under conditions comparable to those for polynomial lattices in [9], while digital nets under the random linear scrambling attain the same property under strictly weaker conditions than those assumed in [9].
+We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [kuo:sloa:wasi:wozn:2010].
+Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [nova:wozn:2008], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3).
+This notion of relative variation aligns with the effective dimension framework introduced in [cafmowen], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions.
+The connection to weighted function spaces [sloa:wozn:1998] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
+Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [Goda2024].
+A key distinction lies in the error criterion: their analysis employs the worst-case integration error over the unit ball in a weighted Sobolev space, whereas we demonstrate that median estimators achieve near-optimal convergence rates on a fixed integrand with high probability.
+Our probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [kunsch2019solvable, kunsch2019optimal].
+Within this framework, we prove, informally stated, the following result:
+Theorem 1.
+Given $\eta>0$, $k\in\mathbb{N}$ and $f\in C^{(k,\dots,k)}([0,1]^{s})$, it holds with probability $1-2^{-r}$ that the median estimator $\hat{\mu}^{(r)}_{\infty}$ from Subsection 2.3 satisfies
+$|\hat{\mu}^{(r)}_{\infty}-\mu|\leqslant C_{\eta,k,f}n^{-k-1/2+\eta}.$
+The constant $C_{\eta,k,f}$ is independent of $s$ under the random linear scrambling if the relative variation $\gamma_{u}$ satisfies the tractability conditions in Theorem 5.
+These nearly $O(n^{-k-1/2})$ integrand-specific error rates complement the nearly $O(n^{-k})$ worst-case rates established in [Goda2024].
+Furthermore, Theorem 5 requires strictly weaker assumptions than those assumed in [Goda2024] for the random linear scrambling, demonstrating that dimension-independent convergence is more easily achieved under our error criterion.
+As a final remark, we note that while several algorithms match lower complexity bounds for integration in specific spaces—such as higher-order digital nets [goda2016explicit, goda2017optimal, goda2018optimal] and Frolov lattice rules [frolov1976upper, krieg2017universal, ullrich2016role]—our method offers distinct practical advantages.
+It requires no smoothness parameter specification, yet automatically matches theoretical lower bounds up to an arbitrarily small $\eta>0$ in the convergence rate, making it both efficient and immune to parameter misspecification.
+Combined with its straightforward implementation and the dimension-independent convergence established in this paper, the median estimator presents a highly practical and robust alternative for applied problems.
+Furthermore, recent work has successfully extended median RQMC to $L^{2}$-approximation [pan2025universal, pan2025l_2], demonstrating its broader utility beyond numerical integration.
 The remainder of this paper is structured as follows.
@@
 2 Background and notation
-Let =$\mathbb{N}=\{1,2,3,\dots\}$ denote the natural numbers and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$.
+Let $\mathbb{N}=\{1,2,3,\dots\}$ denote the natural numbers and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$.
 For $s$-dimensional indices, we define $\mathbb{N}^{s}_{*}=\mathbb{N}_{0}^{s}\setminus\{\boldsymbol{0}\}$ to exclude the zero vector.
-We define =$\mathbb{Z}_{\leqslant\ell}=\{0,1,\dots,\ell\}$ for $\ell\in\mathbb{N}_{0}$ and <=$\mathbb{Z}_{<\ell}=\{0,1,\dots,\ell-1\}$ for $\ell\in\mathbb{N}$.
+We define $\mathbb{Z}_{\leqslant\ell}=\{0,1,\dots,\ell\}$ for $\ell\in\mathbb{N}_{0}$ and $\mathbb{Z}_{<\ell}=\{0,1,\dots,\ell-1\}$ for $\ell\in\mathbb{N}$.
 The dimension of the integration domain is denoted by $s$, and we write $1{:}s=\{1,2,\dots,s\}$ for $s\in\mathbb{N}$.
-For a vector $\boldsymbol{x}\in{}^{s}$ and a subset $u\subseteq 1{:}s$, $\boldsymbol{x}_{u}$ denotes the subvector of $\boldsymbol{x}$ indexed by $u$, while $\boldsymbol{x}_{u^{c}}$ denotes the subvector indexed by $1{:}s\setminus u$.
-The indicator function $\boldsymbol{1}\{\mathcal{A}\}$ equals $1$ if event $\mathcal{A}$ occurs and 00 otherwise.
+For a vector $\boldsymbol{x}\in\real^{s}$ and a subset $u\subseteq 1{:}s$, $\boldsymbol{x}_{u}$ denotes the subvector of $\boldsymbol{x}$ indexed by $u$, while $\boldsymbol{x}_{u^{c}}$ denotes the subvector indexed by $1{:}s\setminus u$.
+The indicator function $\boldsymbol{1}\{\mathcal{A}\}$ equals $1$ if event $\mathcal{A}$ occurs and 0 otherwise.
 For $\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{s})\in\mathbb{N}^{s}_{0}$, $\|\boldsymbol{\ell}\|_{1}=\sum_{j=1}^{s}\ell_{j}$.
 The cardinality of a set $K$ is $|K|$.
-For positive sequence $a_{n}$ and $b_{n}$, $a_{n}=O(b_{n})$ denotes asymptotic dominance: $a_{n}\leqslant Cb_{n}$ for some constant $C<\infty$ and all sufficiently large $n\in\mathbb{N}$.
+For positive sequences $a_{n}$ and $b_{n}$, $a_{n}=O(b_{n})$ denotes asymptotic dominance: $a_{n}\leqslant Cb_{n}$ for some constant $C<\infty$ and all sufficiently large $n\in\mathbb{N}$.
 We assume the integrand $f:[0,1]^{s}\to\mathbb{R}$ is continuous, denoted as $f\in C([0,1]^{s})$, with $\|f\|_{\infty}=\max_{\boldsymbol{x}\in[0,1]^{s}}|f(\boldsymbol{x})|$.
@@
 2.1 Digital net construction and randomization
-For $m\in\mathbb{N}$ and $i\in\mathbb{Z}_{2^{m}}$, we express $i$ in its binary form $i=\sum_{\ell=1}^{m}i_{\ell}2^{\ell-1}$, encoded as the column vector $\vec{i}=(i_{1},\dots,i_{m})^{\mathsf{T}}\in\{0,1\}^{m}$.
+For $m\in\mathbb{N}$ and $i\in\mathbb{Z}_{<2^{m}}$, we express $i$ in its binary form $i=\sum_{\ell=1}^{m}i_{\ell}2^{\ell-1}$, encoded as the column vector $\vec{i}=(i_{1},\dots,i_{m})^{\mathsf{T}}\in\{0,1\}^{m}$.
 Analogously, for $a\in[0,1)$, we approximate its infinite binary expansion $a=\sum_{\ell=1}^{\infty}a_{\ell}2^{-\ell}$ by truncating to $E\in\mathbb{N}$ digits, yielding $\vec{a}=(a_{1},\dots,a_{E})^{\mathsf{T}}\in\{0,1\}^{E}$.
 For numbers with dual representations (such as dyadic rationals), we select the finite expansion ending in zeros.
-A base-2 digital net in $[0,1]^{s}$ is determined by $s$ matrices $C_{j}\in\{0,1\}^{E\times m}$. The deterministic points $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{is})$ are constructed via:
-$\vec{x}_{ij}=C_{j}\vec{i}\ \ \mathsf{mod}\ 2\text{ for }i\in\mathbb{Z}_{2^{m}},j\in 1{:}s,$
+A base-2 digital net in $0,1]^{s}$ is determined by $s$ matrices $C_{j}\in\{0,1\}^{E\times m}$. The deterministic points $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{is})$ are constructed via:
+$\vec{x}_{ij}=C_{j}\vec{i}\ \ \mathsf{mod}\ 2\text{ for }i\in\mathbb{Z}_{<2^{m}},j\in 1{:}s,$
 where $\vec{x}_{ij}\in\{0,1\}^{E}$ maps to $x_{ij}\in[0,1)$ by appending zeros to the truncated $E$-digit representation. Classical constructions typically assume $E\leqslant m$.
 To introduce randomness, we modify the construction as:
-[…]
+$\vec{x}_{ij}=C_{j}\vec{i}+\vec{D}_{j}\ \ \mathsf{mod}\ 2,$ (1)
 where $C_{j}\in\{0,1\}^{E\times m}$ and $\vec{D}_{j}\in\{0,1\}^{E}$ are random with precision $E\geqslant m$.
-The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli$(0.5)$ entries.
-A common way to randomize $C_{j}$ is the random linear scrambling [21]:
-[…]
-where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli$(0.5)$), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [5, Chapter 4.4] for details).
+The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli(0.5)(0.5) entries.
+A common way to randomize $C_{j}$ is the random linear scrambling [mato:1998:2]:
+$C_{j}=M_{j}\mathcal{C}_{j}\ \ \mathsf{mod}\ 2,$
+where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli(0.5)(0.5)), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [dick:pill:2010, Chapter 4.4] for details).
 While effective, the reliance of random linear scrambling on carefully designed generating matrices complicates theoretical analysis.
 We address this in Section 5.
-For initial insights, Section 4 adopts the complete random design [25], where all entries of $C_{j}$ are drawn independently from Bernoulli$(0.5)$.
+For initial insights, Section 4 adopts the complete random design [pan2024automatic], where all entries of $C_{j}$ are drawn independently from Bernoulli(0.5)(0.5).
 This approach retains the asymptotic convergence rate of random linear scrambling while circumventing the need for pre-optimized generator matrices.
@@
 In practice, $E$ is constrained by computational precision.
-Lemma 1 of [27] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$.
+Lemma 1 of [superpolymulti] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$.
 Thus, $\hat{\mu}_{E}$ closely approximates $\hat{\mu}_{\infty}$ when $\omega_{f}(\sqrt{s}2^{-E})\ll|\hat{\mu}_{E}-\mu|$.
@@
 $\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})=\prod_{j=1}^{s}\mathrm{wal}_{k_{j}}(x_{j})=(-1)^{\sum_{j=1}^{s}\vec{k}_{j}^{\mathsf{T}}\vec{x}_{j}},$
-[…]
+where $\boldsymbol{k}=(k_{1},\dots,k_{s})\in\mathbb{N}^{s}_{0}$.
+These functions constitute a complete orthonormal basis for $L^{2}([0,1]^{s})$ [dick:pill:2010], enabling the Walsh series decomposition:
 $f(\boldsymbol{x})=\sum_{\boldsymbol{k}\in\mathbb{N}_{0}^{s}}\hat{f}(\boldsymbol{k})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x}),\quad\text{where}\quad\hat{f}(\boldsymbol{k})=\int_{[0,1]^{s}}f(\boldsymbol{x})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}.$ (3)
-The equality holds in the $L^{2}$-sense. Using this decomposition, [25, Theorem 1] establishes a key error structure for QMC estimators:
+The equality holds in the $L^{2}$-sense. Using this decomposition, [pan2024automatic, Theorem 1] establishes a key error structure for QMC estimators:
 Lemma 1.
@@
 Here, $Z(\boldsymbol{k})$ is an indicator for failure to integrate the Walsh mode $\mathrm{wal}_{\boldsymbol{k}}$ exactly, and $S(\boldsymbol{k})$ encodes the random sign induced by the digital shifts $\vec{D}_{j}$.
-2.3 Probabilistic error guarantee of the median
-The analysis of equation (4) uses two key properties of the coefficients $S(\boldsymbol{k})$: they are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [26, Lemma 4] for the univariate case $s=1$).
-These properties allow us to compute the variance of $\hat{\mu}_{\infty}$ and construct probabilistic error bounds by Chebyshev’s inequality.
-To sharpen these bounds, we introduce a low-probability exclusion event $\mathcal{A}$ satisfying $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and $\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\ll\mathrm{Var}(\hat{\mu}_{\infty})$.
-This leads to the following median concentration result:
+2.3 Probabilistic error guarantee of the median estimator
+Given $r\in\mathbb{N}$, we consider 2r12r-1 independently generated replicates of $\hat{\mu}_{\infty}$ ordered from the smallest to the largest:
+$\hat{\mu}^{(1)}_{\infty}\leqslant\dots\leqslant\hat{\mu}^{(r)}_{\infty}\leqslant\dots\leqslant\hat{\mu}^{(2r-1)}_{\infty}.$
+A key property of the sample median $\hat{\mu}^{(r)}_{\infty}$ is the following probabilistic error bound:
 Lemma 2.
-Let $\mathcal{A}$ be an event with $\Pr(\mathcal{A})\leqslant\delta<1/8$ and $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$.
-Then the sample median $\hat{\mu}^{(r)}_{\infty}$ of $12r-$ 1 independently generated replicates of $\hat{\mu}_{\infty}$ satisfies
-$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\delta^{-1}\Pr(\mathcal{A}^{c})\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\Big{)}\leqslant(8\delta)^{r}.$
+Let $K\subseteq\mathbb{N}^{s}_{*}$ and $\mathcal{A}$ be the event $\{Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in K\}$. If $\Pr(\mathcal{A})\leqslant\delta<1/8$, then
+$\Pr(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta)\leqslant(8\delta)^{r}\quad\text{for}\quad\varepsilon_{K}=\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}.$
 Proof.
-This is an intermediate result proven in [27, Lemma 6]. ∎
-For $f\in C([0,1]^{s})$, the above lemma implies the mean squared error bound
-$\mathbb{E}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}\leqslant\delta^{-1}\Pr(\mathcal{A}^{c})\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})+(8\delta)^{r}\|f\|^{2}_{\infty}.$
+First note that the random signs $S(\boldsymbol{k})$ in equation (4) are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [superpolyone, Lemma 4] for the one-dimensional case).
+Since $\mathcal{A}$ is independent of $S(\boldsymbol{k})$, we have $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and
+$\mathrm{Var}(\hat{\mu}_{\infty}\mid\mathcal{A}^{c})=\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1\mid\mathcal{A}^{c})|\hat{f}(\boldsymbol{k})|^{2}\leqslant\frac{\varepsilon_{K}}{\Pr(\mathcal{A}^{c})}.$
+Applying Chebyshev’s inequality yields
+$\Pr(|\hat{\mu}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta)\leqslant$ $\Pr(\mathcal{A})+\Pr(\mathcal{A}^{c})\Pr(|\hat{\mu}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta\mid\mathcal{A}^{c})\leqslant 2\delta.$
+Finally, $|\hat{\mu}^{(r)}_{\infty}-\mu|>\varepsilon_{K}/\delta$ only if the event $|\hat{\mu}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta$ occurs for at least $r$ out of the 2r12r-1 independent replicates.
+A union bound then implies
+$\Pr(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta)\leqslant{2r-1\choose r}(2\delta)^{r}\leqslant 2^{2r-1}(2\delta)^{r}\leqslant(8\delta)^{r}.\qed$
+This lemma justifies the use of the sample median $\hat{\mu}^{(r)}_{\infty}$ as an estimator for $\mu$. For $f\in C([0,1]^{s})$, the mean squared error satisfies
+$\mathbb{E}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}\leqslant\varepsilon_{K}/\delta+(8\delta)^{r}\|f\|^{2}_{\infty}.$
 Because the second term decays exponentially in $r$, it can be made negligible relative to the first term by choosing $r$ sufficiently large.
-We choose $\mathcal{A}$ to be the event $\{Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K\}$ for a carefully chosen index set $K\subseteq\mathbb{N}^{s}_{*}$.
-It is shown in [25, Subsection 2.4] that
-${\Pr(\mathcal{A}^{c})}\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\leqslant\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}.$ (5)
-The central challenge is to select $K$ such that this sum converges at the desired rate.
-Our analysis addresses this by constructing $K$ adaptively based on the decay rate of the Walsh coefficients $\hat{f}(\boldsymbol{k})$, which encodes the integrand’s smoothness and effective dimensionality.
+The main challenge is therefore to select a set $K$ so that $\varepsilon_{K}$ converges at the desired rate.
+We address this by constructing $K$ adaptively based on the decay rate of the Walsh coefficients $\hat{f}(\boldsymbol{k})$, which encodes the integrand’s smoothness and effective dimensionality.
+Notably, the median estimator $\hat{\mu}^{(r)}_{\infty}$ does not require explicit knowledge of $K$, making it applicable without prior information.
 2.4 Walsh coefficients and function smoothness
@@
 For a finite subset $\kappa\subseteq\mathbb{N}$, we define
-  • $\lceil\kappa\rceil_{q}$: The $q$-th largest element of $\kappa$ (or 00 if $|\kappa|<q$).
+  • $\lceil\kappa\rceil_{q}$: The $q$-th largest element of $\kappa$ (or 0 if $|\kappa|<q$).
   • $\lceil\kappa\rceil_{1{:}q}$: The set of the $q$ largest elements of $\kappa$ (or $\varnothing$ if $q=0$).
-[…]
+The following lemma connects the Walsh coefficients $\hat{f}(\boldsymbol{k})$ to the smoothness of $f$.
+For $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{s})\in\mathbb{N}^{s}_{0}$, we define the mixed partial derivative:
 $f^{(\boldsymbol{\alpha})}=f^{(\alpha_{1},\dots,\alpha_{s})}=\frac{\partial^{\|\boldsymbol{\alpha}\|_{1}}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{s}^{\alpha_{s}}},$
@@
 where $W_{\kappa}:[0,1]\to\mathbb{R}$ for $\kappa\subseteq\mathbb{N}$ is defined recursively by $W_{\varnothing}(x)=1$ and
-$W_{\kappa}(x)=\int_{[0,1]}(-1)^{\vec{x}(\lfloor\kappa\rfloor)}W_{\kappa\setminus\lfloor\kappa\rfloor}(x)\,\mathrm{d}x.$
+$W_{\kappa}(x)=\int_{[0,1]}(-1)^{\vec{x}(\lfloor\kappa\rfloor)}W_{\kappa\setminus\lfloor\kappa\rfloor}(x)\,\mathrm{d}x$
 with $\lfloor\kappa\rfloor$ denoting the smallest element of $\kappa$.
@@
 Proof.
-Equation (6) comes from [30, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [30, Section 3]. ∎
-For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $[0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$.
+Equation (5) comes from [SUZUKI20161, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [SUZUKI20161, Section 3]. ∎
+For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$.
 In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3.
-This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (6), as discussed in [25, Remark 1].
+This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (5), as discussed in [pan2024automatic, Remark 1].
 3 ANOVA decomposition and effective dimensionality
-To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [19]:
+To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [meandim]:
 $f=\sum_{u\subseteq 1{:}s}f_{u},$
@@
 if $j\in\boldsymbol{s}(\boldsymbol{k})$. Thus, $\hat{f}(\boldsymbol{k})=\hat{f}_{u}(\boldsymbol{k})$ if $\boldsymbol{s}(\boldsymbol{k})=u$.
-To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [3].
+To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [dick:2008], which generalizes the usual Sobolev norms.
 For a function $f:[0,1]^{d}\to\mathbb{R}$ and a subcube $J=\prod_{j=1}^{d}[a_{j},b_{j})$, we define the alternating sum over $J$:
@@
 $V^{(d)}_{\lambda}(f)=\sup_{\mathcal{P}}\Bigl(\sum_{J\in\mathcal{P}}\mathrm{Vol}(J)\Bigl|\frac{\Delta(f,J)}{\mathrm{Vol}(J)^{\lambda}}\Bigr|^{2}\Bigr)^{1/2}$ (9)
-where the supremum is over all partitions $\mathcal{P}$ of $[0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$.
+where the supremum is over all partitions $\mathcal{P}$ of $0,1)^{d}$ into subcubes and $\mathrm{Vol}(J)$ is the volume of subcube $J$.
 When $d=0$ and $f\in\mathbb{R}$, $V^{(0)}_{\lambda}(f)=|f|$ by convention.
-The Vitali variation generalizes Sobolev norms: for $f$ with weak mixed derivative $f^{(1,\dots,1)}\in L^{p}([0,1]^{d})$, [20] shows $V_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$.
-Crucially, $V^{(d)}_{\lambda}(f)$ can be finite even for discontinuous $f$ (see [25, Section 4.1] for an example).
+For $f$ with weak mixed derivative $f^{(1,\dots,1)}$, we have the inequality
+$V^{(d)}_{1}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{2}},$ (10)
+with equality attained when $f\in C^{(1,\dots,1)}([0,1]^{s})$ [dick:2008, equation (3.4)].
+More generally, [LIU2025101935] shows $V^{(d)}_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$.
+Discontinuous $f$ can also have finite $V^{(d)}_{\lambda}(f)$ (see [pan2024automatic, Section 4.1] for an example).
 Next, for $v\subseteq 1{:}d$, we define the partial integral operator $I_{v}:L^{1}([0,1]^{d})\to L^{1}([0,1]^{|v|})$ as:
@@
 $V^{v}_{\lambda}(f)=\sup_{\rho_{v^{c}}\in D_{v^{c}}}V^{(|v|)}_{\lambda}\Big(I_{v}(f\rho_{v^{c}})\Big).$ (12)
-By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [25, Section 3.2],
+By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [pan2024automatic, Section 3.2],
 $\Bigl(V^{v}_{\lambda}(f)\Bigl)^{2}\leqslant\int_{[0,1]^{d-|v|}}\Bigl(V^{(|v|)}_{\lambda}\Bigl(f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr)\Bigr)^{2}\,\mathrm{d}\boldsymbol{x}_{v^{c}},$
-implying $V^{v}_{\lambda}(f)<\infty$ if almost every section $f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})$ has bounded variation.
+which, together with equation (10), implies
+$V^{v}_{1}(f)\leqslant\|f^{(\boldsymbol{1}_{v},\boldsymbol{0}_{v^{c}})}\|_{L^{2}}$ (13)
+if the weak derivative $f^{(\boldsymbol{1}_{v},\boldsymbol{0}_{v^{c}})}$ exists.
+Here $(\boldsymbol{1}_{v},\boldsymbol{0}_{v^{c}})$ denotes the vector whose $j$-th entry equals $1$ if $j\in v$ and equals 0 otherwise.
 The following lemma connects the Walsh coefficients $\hat{f}(\boldsymbol{k})=\hat{f}_{u}(\boldsymbol{k})$ (where $\boldsymbol{s}(\boldsymbol{k})=u$) to the fractional Vitali variation of $f_{u}$:
@@
 Let $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$.
-Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})<\infty$ for nonempty $u\subseteq 1{:}s$, $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and a subset $v\subseteq\{j\mid\alpha_{j}=\alpha\}$.
+Suppose $V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})$$\infty$ for nonempty $u\subseteq 1{:}s$, $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and a subset $v\subseteq\{j\mid\alpha_{j}=\alpha\}$.
 Then for $\boldsymbol{k}_{u}=(k_{j},j\in u)\in\mathbb{N}^{|u|}$ satisfying $|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$,
@@
 Proof.
-Let $\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)$.
-Because $|\kappa^{\prime}_{j}|\geqslant|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa^{\prime}_{j}|=|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$, we can apply Lemma 3 to get
-$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=$
-$\Big{|}\int_{[0,1]^{|u|}}\Big{(}f_{u}^{(\boldsymbol{\alpha}_{u})}(\boldsymbol{x})\prod_{j\in u\setminus v}W_{\kappa_{j}}(x_{j})\Big{)}\Big{(}\prod_{j\in v}{\mathrm{wal}_{\kappa^{\prime}_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}\alpha}}}(x_{j})W_{\lceil\kappa_{j}\rceil_{1{:}\alpha}}(x_{j})\Big{)}\,\mathrm{d}\boldsymbol{x}_{u}\Big{|}^{2},$
-Let $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$, $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ and $\kappa^{-}_{j}=\kappa^{\prime}_{j}\setminus\kappa_{j}$ for $j\in v$.
-We can rewrite the above equation as
-$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=\Bigl{|}\int_{[0,1]^{|v|}}g(\boldsymbol{x}_{v})\prod_{j\in v}{\mathrm{wal}_{\kappa^{-}_{j}}}(x_{j})(-1)^{\vec{x}_{j}(\ell_{j})}W_{\kappa^{+}_{j}}(x_{j})\,\mathrm{d}\boldsymbol{x}_{v}\Bigr{|}^{2},$
-where
-$g(\boldsymbol{x}_{v})=I_{v}\Big{(}f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}W_{\kappa_{j}}\Big{)}.$
-From the proof of [25, Theorem 3], we can map $v$ to $1{:}|v|$ and show
-$\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}\leqslant 2^{-|v|}\Big{(}V^{(|v|)}_{\lambda}(g)\Big{)}^{2}\Bigl{(}\prod_{j\in v}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}\Bigr{)}.$ (14)
-Letting $\rho_{j}(x_{j})=W_{\kappa_{j}}(x_{j})/\|W_{\kappa_{j}}\|_{\infty}$ for $j\in u\setminus v$, equation (7) implies
-$V^{(|v|)}_{\lambda}(g)=$ $\Big{(}\prod_{j\in u\setminus v}2\prod_{\ell^{\prime}\in\kappa_{j}}2^{-\ell^{\prime}-1}\Big{)}V^{(|v|)}_{\lambda}\Big{(}I_{u}(f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}\rho_{j})\Big{)}$
-$\leqslant$ $2^{|u|-|v|}\Big{(}\prod_{j\in u\setminus v}\prod_{\ell^{\prime}\in\kappa_{j}}2^{-\ell^{\prime}-1}\Big{)}V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})$
-because $\prod_{j\in u\setminus v}\rho_{j}\in D_{v^{c}}$ from equation (11). Our conclusion follows once we put the above bound into equation (14) and plug in
-$\prod_{j\in v}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}=4^{|v|}\prod_{j\in v}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
-and $\lceil\kappa_{j}\rceil_{\alpha+1}=0$ for $j\in u\setminus v$. ∎
+See Appendix A. ∎
 Motivated by Lemma 4, we define for $\alpha\in\mathbb{N}_{0},\lambda\in(0,1]$ and nonempty $u\subseteq 1{:}s$:
@@
 Notice that we add the restriction $\alpha_{j}>0$ for $j\in u\setminus v$ because $k_{j}\in\mathbb{N}$ implies $\alpha_{j}=|\kappa_{j}|>0$ in the statement of Lemma 4.
-In particular, $\|f\|_{u,\alpha,\lambda}=V^{u}_{\lambda}(f_{u})$ when $\alpha=0$.
+In particular, $\|f\|_{u,0,\lambda}=V^{u}_{\lambda}(f_{u})$.
+When $\lambda=1$ and the weak mixed derivative $f^{(\alpha+1,\dots,\alpha+1)}$ exists, equation (13) implies the bound
+$\|f\|_{u,\alpha,1}\leqslant\sup_{\boldsymbol{\alpha}_{u}\in\{1:(\alpha+1)\}^{|u|}}\|f_{u}^{(\boldsymbol{\alpha}_{u})}\|_{L^{2}}.$
 We further set $\|f\|_{u,\alpha,\lambda}=0$ if $u=\varnothing$, and define the global variation norm:
@@
 For non-constant $f$, the relative variation of ANOVA component $f_{u}$ is:
-$\gamma_{u}=\frac{\|f\|_{u,\alpha,\lambda}}{\|f\|_{\alpha,\lambda}},$ (16)
+$\gamma_{u}=\gamma_{u}(f,\alpha,\lambda)=\frac{\|f\|_{u,\alpha,\lambda}}{\|f\|_{\alpha,\lambda}},$ (16)
 which satisfies $\sup_{u\subseteq 1{:}s}\gamma_{u}=1$ and $\gamma_{\varnothing}=0$.
+The relative variation $\gamma_{u}$ depends on $f,\alpha$ and $\lambda$, but we suppress this dependence in the notation as these variables are fixed throughout our theoretical analysis.
+Section 6 later exemplifies how $\gamma_{u}$ changes under different choices of $\alpha$.
 The decay of $\gamma_{u}$ encodes effective dimensionality, with $\gamma_{u}=1$ for all nonempty $u\subseteq 1{:}s$ representing the worst dimensionality.
@@
 First, we define $\|\kappa\|=\sum_{\ell\in\kappa}\ell$ for $\kappa\subseteq\mathbb{N}$, with $\|\varnothing\|=0$.
-Motivated by Lemma 4, we choose $K$ in equation (5) for given $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$ to be:
+Motivated by Lemma 4, we choose $K$ in Lemma 2 for given $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$ to be:
 $K=\bigcup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}K_{u}(T_{u,m})$ (17)
@@
 Proof.
-By equation (9), $\hat{f}_{u}(\boldsymbol{k})=0$ if $\boldsymbol{s}(\boldsymbol{k})\neq u$. Letting
-$\mathbb{N}^{u}_{\alpha}=\{(k_{j},j\in u)\in\mathbb{N}^{|u|}\mid|\kappa_{j}|\leqslant\alpha+1\text{ for }j\in u\},$
-we can apply Lemma 4 and equation (15) to get
-$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T)}|\hat{f}_{u}(\boldsymbol{k})|^{2}$ (19)
-$\leqslant$ $4^{|u|}\|f\|^{2}_{u,\alpha,\lambda}\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\boldsymbol{1}\Big{\{}\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)\Big{\}}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}.$
-Because $\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)$ implies
-$(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|>T,$
-we have for any $\theta\in(0,1)$ and $\rho_{\theta}=4^{-1+\theta}$
-$\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\boldsymbol{1}\Big{\{}\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)\Big{\}}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
-$\leqslant$ $\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}4^{-\theta T+\theta(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\theta\sum_{j\in u}\|\kappa_{j}\|}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
-$=$ $4^{-\theta T}\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\prod_{j\in u}4^{-(1-\theta)(\lambda-1)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-(1-\theta)\ell^{\prime}-1}$
-$=$ $4^{-\theta T}\Big{(}\sum_{\alpha^{\prime}=0}^{\alpha}4^{-\alpha^{\prime}}\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\quad+\quad 4^{-\alpha-1}\sum_{\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})}\rho_{\theta}^{(\lambda-1)\lceil\kappa\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\Big{)}^{|u|},$
-where $\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})=\{\kappa\subseteq\mathbb{N}\mid|\kappa|=\alpha^{\prime}\}$. To bound the above sum, consider the generating function
-$g(t)=\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha^{\prime}},$
-and let $[t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [6] for background knowledge on generating functions).
-It follows that $[t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$.
-Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
-$\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}t^{\ell^{\prime}}\leqslant\frac{1}{(\alpha^{\prime})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha^{\prime}}=\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}.$ (20)
-with $t=\rho_{\theta}$.
-For $\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})$, we let $\kappa^{+}=\lceil\kappa\rceil_{1{:}\alpha}$ and $\ell=\lceil\kappa\rceil_{\alpha+1}$.
-By the one-to-one correspondence between $\kappa^{+}$ and $\kappa^{+}-\ell=\{\ell^{\prime}-\ell\mid\ell^{\prime}\in\kappa\}\in\mathcal{S}_{\alpha}(\mathbb{N})$,
-$\sum_{\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})}\rho_{\theta}^{(\lambda-1)\lceil\kappa\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}=$ $\sum_{\ell=1}^{\infty}\rho_{\theta}^{(\alpha+\lambda)\ell}\sum_{\kappa^{+}-\ell\in\mathcal{S}_{\alpha}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa^{+}-\ell}\rho_{\theta}^{\ell^{\prime}}$
-$=$ $\sum_{\ell=1}^{\infty}\rho_{\theta}^{(\alpha+\lambda)\ell}\frac{1}{\alpha!(\rho_{\theta}^{-1}-1)^{\alpha}}$
-$=$ $\frac{1}{\alpha!(\rho^{-\alpha-\lambda}_{\theta}-1)(\rho_{\theta}^{-1}-1)^{\alpha}}.$
-Our conclusion follows once we put the above bounds into equation (19). ∎
+See Appendix B. ∎
 While larger $T$ implies a tighter bound on $|\hat{f}_{u}(\boldsymbol{k})|^{2}$ with $\boldsymbol{k}\notin K_{u}(T)$, it also increases the probability that $Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K_{u}(T)$.
@@
 $A_{\alpha,\lambda}=\frac{1}{\alpha!(2^{1/(\alpha+\lambda)}-1)^{\alpha}}\quad\text{and}\quad B_{\alpha,\lambda}=\sum_{\alpha^{\prime}=1}^{\alpha}\frac{1}{(\alpha^{\prime})!(2^{1/(\alpha+\lambda)}-1)^{\alpha^{\prime}}},$ (19)
-with the sum from $1$ to $\alpha$ interpreted as 00 when $\alpha=0$.
+with the sum from 11 to $\alpha$ interpreted as 0 when $\alpha=0$.
 Proof.
-The conclusion is trivial when $T<0$, so we assume $T\geqslant 0$.
-As in the proof Lemma 4, for each $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ we write $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$ and $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ for $j\in v$.
-By equation (13), each $\boldsymbol{k}^{\prime}=(k^{\prime}_{1},\dots,k^{\prime}_{s})\in B(\boldsymbol{k}_{u},v)$ must have the same $\kappa^{\prime}_{j}$ for $j\in 1{:}s\setminus v$.
-For $j\in v$, however, $\kappa^{\prime}_{j}\setminus\kappa_{j}$ can be any subset of $1{:}(\ell_{j}-1)$ if $\ell_{j}\geqslant 2$ and $\kappa^{\prime}_{j}=\kappa_{j}$ if $\ell_{j}=1$.
-Thus, $|B(\boldsymbol{k}_{u},v)|=\prod_{j\in v}2^{\ell_{j}-1}$.
-Because $B(\boldsymbol{k}_{u},v)$ are mutually disjoint for distinct $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$,
-$|K_{u}(T)|=\sum_{v\subseteq u}\sum_{\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)}\prod_{j\in v}2^{\ell_{j}-1}.$ (22)
-Next, we notice
-$(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|=(\alpha+\lambda)\sum_{j\in v}\ell_{j}+\sum_{j\in v}\sum_{\ell^{\prime}\in\kappa^{+}_{j}}(\ell^{\prime}-\ell_{j})+\sum_{j\in u\setminus v}\|\kappa_{j}\|$
-and there is a one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}=\{\ell^{\prime}-\ell_{j}\mid\ell^{\prime}\in\kappa^{+}_{j}\}$ for $j\in v$.
-Furthermore, $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ implies
-$\sum_{j\in v}\sum_{\ell^{\prime}\in\kappa^{+}_{j}}(\ell^{\prime}-\ell_{j})+\sum_{j\in u\setminus v}\|\kappa_{j}\|=\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|+\sum_{j\in u\setminus v}\|\kappa_{j}\|\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}.$
-For $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$, let
-$\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\mathbb{N}_{0}^{|u|}\Big{|}|\kappa_{j}|=\alpha_{j}\ \forall j\in u,\sum_{j\in u}\|\kappa_{j}\|=N\Big{\}}.$
-We can decompose $\mathcal{K}_{u,v}(T)$ into subsets with the same $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ and $(\kappa^{+}_{j}-\ell_{j},j\in v;\kappa_{j},j\in u\setminus v)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ for $N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}$, and rewrite equation (22) as
-$|K_{u}(T)|=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$ (23)
-$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|+$
-$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
-Let $\overline{N}=\lfloor(T-N)/(\alpha+\lambda)\rfloor$.
-For nonempty $v\subseteq u$ and integer $n\geqslant|v|$, there are (${n-1\choose|v|-1}$ number of $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfying $\|\boldsymbol{\ell}_{v}\|_{1}=n$.
-Hence, when $\overline{N}\geqslant|v|$,
-$\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}=$ $\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}$ (24)
-$=$ $\sum_{N=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n+1}-\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}$
-$=$ (${\overline{N}\choose|v|-1}2^{\overline{N}+1}-\sum_{n=|v|}^{\overline{N}}\Big{\{}{n-1\choose|v|-1}-{n-2\choose|v|-1}\Big{\}}2^{n}$
-$\leqslant$ (${\overline{N}\choose|v|-1}2^{\overline{N}+1}.$
-When $\overline{N}<|v|$, no $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfies $\|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)$ and the above sum is trivially 00.
-Putting the above bound into equation (23), we get
-$|K_{u}(T)|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|$ (25)
-$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
-$\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}.$
-To bound the sum over $N$, consider the generating function
-$g(t)=\prod_{j\in u}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}.$
-Because each element of $\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ is counted $\prod_{j\in u}(\alpha_{j})!$ times by $t^{N}[g(t)]$ , an argument similar to equation (20) gives
-$\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|t^{N}\leqslant\prod_{j\in u}\frac{1}{(\alpha_{j})!}\Big{(}\sum_{N=1}^{\infty}t^{N}\Big{)}^{\alpha_{j}}=\prod_{j\in u}\frac{1}{(\alpha_{j})!(t^{-1}-1)^{\alpha_{j}}}$ (26)
-for $t=2^{-1/(\alpha+\lambda)}$ and
-$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\prod_{j\in u}\frac{1}{(\alpha_{j})!(2^{1/(\alpha+\lambda)}-1)^{\alpha_{j}}}$
-$=$ $A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|},$
-where $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are defined in equation (21). By a similar calculation,
-$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
-$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}.$
-Putting the above bounds into equation (25), we finally get
-$|K_{u}(T)|\leqslant$ $2^{T/(\alpha+\lambda)}\Big{(}B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}A_{\alpha,\lambda}^{|v|}B^{|u|-|v|}_{\alpha,\lambda}\Big{)}$ (27)
-$\leqslant$ $2^{T/(\alpha+\lambda)}\Big{(}B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}\max\big{(}T/(\alpha+\lambda),1\big{)}^{|v|}A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|}\Big{)}$
-$=$ $2^{T/(\alpha+\lambda)}\Big{(}B_{\alpha,\lambda}+\max\big{(}T/(\alpha+\lambda),1\big{)}A_{\alpha,\lambda}\Big{)}^{|u|}.\qed$ (28)
+See Appendix C. ∎
 Corollary 1.
@@
 Proof.
-Let $\delta=1/16$ in Lemma 2 and Corollary 1. By Lemma 2, equation (5) and $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for $\boldsymbol{k}\neq\boldsymbol{0}$, it suffices to show
+Let $\delta=1/16$ in Lemma 2 and Corollary 1. By Lemma 2 and $\Pr(Z(\boldsymbol{k})=1)=2^{-m}$ for $\boldsymbol{k}\neq\boldsymbol{0}$, it suffices to show
 $2^{-m+4}\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}|\hat{f}(\boldsymbol{k})|^{2}\leqslant\Gamma_{\alpha,\lambda,\gamma_{u},\theta}\|f\|^{2}_{\alpha,\lambda}\frac{\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2\theta(\alpha+\lambda)}}{2^{(2\theta(\alpha+\lambda)+1)m}}$
@@
 Remark 2.
-Given a failure probability tolerance $\varepsilon>0$, we can set $r=\lceil\log_{2}(\varepsilon^{-1})\rceil$ to ensure $2^{-r}\leqslant\varepsilon$.
+Given a failure probability tolerance $\bar{\delta}>0$, we can set $r=\lceil\log_{2}(\bar{\delta}^{-1})\rceil$ to ensure $2^{-r}\leqslant\bar{\delta}$.
 For this choice of $r$, the median estimator $\hat{\mu}^{(r)}_{\infty}$ asymptotically achieves a convergence rate $O(N^{-\alpha-\lambda-1/2+\eta})$ for any $\eta>0$, where $N=r2^{m}$ is the total number of function evaluations.
-To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$, so $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$.
-Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}\leqslant 1$,
-$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big{)}\leqslant 2^{-r}\leqslant\varepsilon,$ (33)
-where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$. This convergence rate is almost optimal [25, Subsection 3.2].
+To justify this, observe that $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)^{2(\alpha+\lambda)}=O(2^{\eta m})$ asymptotically for any $\eta>0$ because $\Gamma_{\alpha,\lambda,\gamma_{u}}(m)$ is a polynomial of $m$.
+Setting $\theta=1-\eta/(2(\alpha+\lambda))$, we derive for all $f$ with $\|f\|_{\alpha,\lambda}<\infty$ that
+$\Pr\Big(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}\|f\|^{2}_{\alpha,\lambda}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big)\leqslant 2^{-r}\leqslant\bar{\delta},$ (24)
+where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$.
+This convergence rate is nearly optimal, as it aligns with the lower bound established in [bakh:1959] for one-dimensional $f$ with $\lambda$-Hölder continuous $f^{(\alpha)}$, a subclass of functions with finite $\|f\|_{\alpha,\lambda}$.
 Remark 3.
-In settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$, dimension-independent error bounds hold if there exist weights $\{\gamma_{j},j\in 1{:}s\}$ satisfying
-$\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}\quad\forall u\subseteq 1{:}s\quad\text{ and }\quad\sum_{j=1}^{s}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\leqslant C_{\infty}$ (34)
-for a constant $C_{\infty}$ independent of $s$. To see this, we first bound
+In settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$, dimension-independent error bounds hold if there exist weights $\{\gamma_{j},j\in\mathbb{N}\}$ satisfying
+$\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}\quad\forall u\subseteq\mathbb{N}\quad\text{ and }\quad C_{\infty}=\sum_{j=1}^{\infty}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}<\infty.$ (25)
+To see this, we first bound
 $\frac{\Gamma_{\alpha,\lambda,\gamma_{u},\theta}}{16^{2\theta(\alpha+\lambda)+1}}\leqslant\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}C_{\alpha,\lambda,\theta}^{|u|}\leqslant\prod_{j=1}^{s}(1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}C_{\alpha,\lambda,\theta})\leqslant\exp(C_{\infty}C_{\alpha,\lambda,\theta}),$
@@
 $\leqslant$ $\prod_{j=1}^{s}\Big(1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})\Big),$
-which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [11, Lemma 3].
+which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [HICKERNELL2003286, Lemma 3].
 Putting the above bounds into equation (22), we conclude that equation (24) holds with a constant $C_{\eta}$ independent of $s$.
-Notably, equation (34) resembles the tractability conditions assumed by [9, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
+Notably, equation (25) resembles the tractability conditions assumed by [Goda2024, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
 5 Tractability under more general randomization
 While the completely random design offers theoretical advantages in near-optimal convergence rates and tractability, the random linear scrambling often outperforms it in practical settings due to tighter error bounds on low-dimensional ANOVA components.
-For example, [12] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets.
+For example, [joe:kuo:2008] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets.
 In this section, we extend the analysis of Section 4 to a broader class of randomization schemes, with the random linear scrambling as a key example.
@@
 $\Pr(Z(\boldsymbol{k})=1)\leqslant 2^{-m+\tilde{t}_{u}}\boldsymbol{1}\Big\{\sum_{j\in u}\|\kappa_{j}\|_{(d)}>d(m-\tilde{t}_{u})\Big\}.$
-From [9, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [9, Section 2].
-We frame this assumption to align with broader applications in higher-order digital nets [4], where the parameter $d$ (the order of digital nets) may exceed $1$.
+From [Goda2024, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [Goda2024, Section 2].
+We frame this assumption to align with broader applications in higher-order digital nets [dick:2011], where the parameter $d$ (the order of digital nets) may exceed $1$.
 This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
 $D_{\alpha,\lambda,d,\theta^{\prime}}=\frac{1}{d!(2^{(1-\theta^{\prime})/\beta}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(2^{1/\beta}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(2^{(1-\theta^{\prime})/\beta}-1)^{\alpha^{\prime}}},$ (26)
-with the sum from $1$ to $\alpha-d$ interpreted as 00 when $\alpha=d$.
+with the sum from 11 to $\alpha-d$ interpreted as 0 when $\alpha=d$.
 Proof.
-We adopt the notation used in the proof of Lemma 6.
-For $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$, we again consider the one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}$ for $j\in v$.
-Because $d\in\mathbb{N}$ and $d<\alpha+\lambda$, we must have $d\leqslant\alpha$.
-By $|\kappa^{+}_{j}|=\alpha\geqslant d$ for $j\in v$,
-$\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|_{(d)}=\sum_{j\in v}\|\kappa^{+}_{j}\|_{(d)}-d\ell_{j}=\Big{(}\sum_{j\in v}\|\kappa_{j}\|_{(d)}\Big{)}-d\|\boldsymbol{\ell}_{v}\|_{1}$
-and
-$\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|_{(d)}+\sum_{j\in u\setminus v}\|\kappa_{j}\|_{(d)}=\Big{(}\sum_{j\in u}\|\kappa_{j}\|_{(d)}\Big{)}-d\|\boldsymbol{\ell}_{v}\|_{1}.$
-Let
-$\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})=\Big{\{}(k_{j},j\in u)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})\Big{|}\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N^{\prime}\Big{\}}.$
-Similar to equation (23), we can write
-$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$ (36)
-$=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}-d\|\boldsymbol{\ell}_{v}\|_{1}<N^{\prime}\leqslant N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$
-$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|+$
-$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
-The sum over $\boldsymbol{\ell}_{v}$ is empty if $(T^{\prime}-N^{\prime})/d\geqslant(T-N)/(\alpha+\lambda)$.
-Because $T^{\prime}/d\geqslant T/(\alpha+\lambda)$ and $N\geqslant N^{\prime}$, we further infer the sum is empty unless $N>N_{*}$ for
-$N_{*}=\frac{(\alpha+\lambda)T^{\prime}-dT}{\alpha+\lambda-d}.$
-Therefore,
-$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
-$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}\sum_{N^{\prime}=0}^{N}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
-$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1},$
-where we have applied equation (24) in the last step. Letting $T_{*}=(T-N_{*})/(\alpha+\lambda)$, we put the above bound into equation (36) and get
-$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (37)
-$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
-$\leqslant$ $2^{T_{*}}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T_{*}\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
-$\leqslant$ $2^{T_{*}}\Big{(}B_{\alpha,\lambda}+\max(T_{*},1)A_{\alpha,\lambda}\Big{)}^{|u|},$
-where the last step follows from a calculation similar to equation (25) and (27).
-Next, we notice
-$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant t^{-T}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}t^{N^{\prime}}\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (38)
-for $t=2^{-1/(\alpha+\lambda-d)}$.
-To bound the sum over $N$, we let $w=\{j\in u\mid\alpha_{j}>d\}$ for $\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and define $\boldsymbol{\alpha}^{-}_{u}=(\min(\alpha_{j},d),j\in u)$ and $\boldsymbol{\alpha}^{+}_{u}=(\alpha_{j}-d,j\in w)$ if $w\neq\varnothing$.
-For $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$, we further let $\kappa^{-}_{j,d}=\lceil\kappa_{j}\rceil_{1{:}\min(\alpha_{j},d)}$ for $j\in u$ and $\kappa^{+}_{j,d}=\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}d}$ for $j\in w$.
-Because for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$,
-$\sum_{j\in w}\|\kappa^{+}_{j,d}\|=\sum_{j\in w}\|\kappa_{j}\|-\sum_{j\in w}\|\kappa^{-}_{j,d}\|=\sum_{j\in u}\|\kappa_{j}\|-\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N-N^{\prime},$
-the number of possible $(\kappa^{+}_{j,d},j\in w)$ for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$ with a given set of $(\kappa^{-}_{j,d},j\in u)\in\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})$ is bounded by the number of $(\kappa^{+}_{j,d},j\in w)$ with $\sum_{j\in w}\|\kappa^{+}_{j,d}\|=N-N^{\prime}$, which is $|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ by definition.
-This implies $|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\times|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|$ and
-$\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=N^{\prime}}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|t^{N-N^{\prime}}$
-$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{w}(N,\boldsymbol{\alpha}^{+}_{u})|t^{N}$
-$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}},$
-where we have applied equation (26) in the last step. Putting the above bound into equation (38), we get for any $\theta^{\prime}\in(0,1)$,
-$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$
-$\leqslant$ $t^{-T}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{N^{\prime}}$
-$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\sum_{N^{\prime}=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{(1-\theta^{\prime})N^{\prime}}$
-$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
-where we have applied equation (26) once more. Our conclusion finally follows from equation (37), $t=2^{-1/(\alpha+\lambda-d)}$ and
-$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\Big{(}\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big{)}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}$
-$=$ $\sum_{w\subseteq u}\Big{(}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{-}_{u}\in\mathbb{Z}^{|u|}_{\leqslant d}\\ \alpha_{j}=d\ \forall j\in w\\ \alpha^{-}_{j}>0\ \forall j\in u\setminus w\end{subarray}}\Big{(}\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big{)}^{|u|}$
-$=$ $\sum_{w\subseteq u}\Big{(}\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big{)}^{|w|}\Big{(}\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big{)}^{|u|-|w|}$
-$=$ $D_{\alpha,\lambda,d,\theta^{\prime}}^{|u|},$
-where $D_{\alpha,\lambda,d,\theta^{\prime}}$ is defined in equation (35). ∎
+See Appendix D. ∎
 Corollary 2.
 For $m\in\mathbb{N}$ and $\gamma_{u}$ defined by equation (16), let
-$\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$ (39)
+[…]
 Then for any $\delta\in(0,1)$,
@@
 $|K_{u}(T_{u,m})\cap K^{\prime}_{u}(T^{\prime}_{u,m})|$
-$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}2^{m-\tilde{t}_{u}}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
+[…]
 A union bound over all $\boldsymbol{k}\in K_{u}(T_{u,m})\cap K^{\prime}_{u}(T^{\prime}_{u,m})$ implies
 $\Pr(Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in K_{u}(T_{u,m}))$
-$\leqslant$ $\frac{\delta\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}}{\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}.$
+[…]
 On the other hand, when $T_{u,m}<T^{\prime}_{u,m}$, every $\boldsymbol{k}\in K_{u}(T_{u,m})$ satisfies
@@
 $\leqslant$ $\|f\|^{2}_{\alpha,\lambda}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}2^{-m+\tilde{t}_{u}}\frac{\gamma^{2}_{u}}{4^{\theta T_{u,m}}}C_{\alpha,\lambda,\theta}^{|u|}.$
-Our conclusion follows from Lemma 2 and equation (5) after plugging in $T_{u,m}$ from equation (40). ∎
+Our conclusion follows from Lemma 2 after plugging in $T_{u,m}$ from equation (28). ∎
 Remark 5.
@@
 Consider again settings where $s\to\infty$ while $\|f\|_{\alpha,\lambda}\leqslant 1$.
-We assume equation (34) holds for weights $\{\gamma_{j},j\in 1{:}s\}$.
+We assume equation (25) holds for weights $\{\gamma_{j},j\in\mathbb{N}\}$.
 Setting $\theta^{\prime}=1-\eta/d$ for $\eta>0$, the analysis in Remark 3 implies $\Gamma_{\alpha,\lambda,d,\theta^{\prime},\gamma_{u}}(m)\leqslant C_{\eta}2^{\eta m/(\alpha+\lambda-d)}$ for a constant $C_{\eta}$ independent of $s$.
 Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
-Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [31] that the base-2 Niederreiter sequence [22] satisfies
-$t_{u}\leqslant\sum_{j\in u}\log_{2}(j)+\log_{2}\log_{2}(j+2)+2,$ (42)
-while the Sobol’ sequence [29] satisfies
-$t_{u}\leqslant\sum_{j\in u}\log_{2}(j)+\log_{2}\log_{2}(j+1)+\log_{2}\log_{2}\log_{2}(j+3)+C$
+Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [c812a4a5-75ac-3614-936d-5782e44941d0] that the base-2 Niederreiter sequence [NIEDERREITER198851] satisfies
+$t_{u}\leqslant\sum_{j\in u}\Big(\log_{2}(j)+\log_{2}\log_{2}(j+2)+2\Big),$ (30)
+while the Sobol’ sequence [sobol67] satisfies
+$t_{u}\leqslant\sum_{j\in u}\Big(\log_{2}(j)+\log_{2}\log_{2}(j+1)+\log_{2}\log_{2}\log_{2}(j+3)+C\Big)$
 for an absolute constant $C$.
-For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for nonnegative numbers $\{\tilde{t}_{j},j\in 1{:}s\}$. We can then bound
+For simplicity, we continue our discussion with the assumption $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for nonnegative numbers $\{\tilde{t}_{j},j\in\mathbb{N}\}$. We can then bound
 $\frac{\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}}{16^{2\theta(\alpha+\lambda-d)+1}}\leqslant$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\prod_{j\in u}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}C_{\alpha,\lambda,\theta}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}$
@@
 Therefore, $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ has a dimension-independent bound if
-$\lim_{s\to\infty}\sum_{j=1}^{s}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
+$\sum_{j=1}^{\infty}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
 Remark 7.
 Viewing $\Gamma_{\alpha,\lambda,\gamma_{u},\tilde{t}_{u},\theta}$ from Theorem 3 as a special case of $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ with $d=\alpha+\lambda$, the argument in Remark 6 yields dimension-independent error bounds for the $d\geqslant\alpha+\lambda$ case if $\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}$, $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ and
-$\lim_{s\to\infty}\sum_{j=1}^{s}\gamma_{j}^{2}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
-While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [4], and unbounded increases in $d$ ultimately violate the tractability condition above.
+$\sum_{j=1}^{\infty}\gamma_{j}^{2}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
+While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [dick:2011], and unbounded increases in $d$ ultimately violate the tractability condition above.
 Thus, prudent selection of $d$ is critical for practical implementations.
@@
 Suppose $\|f\|_{\alpha,\lambda}<\infty$ for $\alpha\in\mathbb{N}_{0}$ and $\lambda\in(0,1]$.
-Given a randomization scheme satisfying Assumption 1 with $d<\alpha+\lambda$, we further assume $\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}$ for $\{\gamma_{j},j\in 1{:}s\}$, $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for $\{\tilde{t}_{j},j\in 1{:}s\}$ and
-$\lim_{s\to\infty}\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big{)}^{\phi}<\infty\quad\text{for}\quad\phi=\frac{\alpha+\lambda+1/2}{\alpha+\lambda-d+1/2}.$ (43)
+Given a randomization scheme satisfying Assumption 1 with $d<\alpha+\lambda$, we further assume $\gamma_{u}\leqslant\prod_{j\in u}\gamma_{j}$ for $\{\gamma_{j},j\in\mathbb{N}\}$, $\tilde{t}_{u}\leqslant\sum_{j\in u}\tilde{t}_{j}$ for $\{\tilde{t}_{j},j\in\mathbb{N}\}$ and
+$\sum_{j=1}^{\infty}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big)^{\phi}<\infty\quad\text{for}\quad\phi=\frac{\alpha+\lambda+1/2}{\alpha+\lambda-d+1/2}.$ (31)
 Then under the given randomization scheme, we can find a constant $C_{\eta}$ independent of $s$ for any $\eta>0$ such that for all $m\in\mathbb{N}$ and $r\in\mathbb{N}$,
-$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big{)}\leqslant 2^{-r}.$
+$\Pr\Big(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>C_{\eta}\|f\|^{2}_{\alpha,\lambda}2^{-(2\alpha+2\lambda+1-2\eta)m}\Big)\leqslant 2^{-r}.$
 Proof.
@@
 for a constant $\Lambda$ independent of $s$. Let $\delta^{\prime}=\min(\Lambda^{-\phi},1/16)$, $\theta^{\prime}=1-\eta/d,$
-$\Gamma(m)=1+\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)}$
+[…]
 and
@@
 $\Pr\Big(Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in\bigcup_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}K_{u}(T_{u,m})\Big)$
-$\leqslant$ $\frac{\delta^{\prime}}{\Gamma(m)}\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big{)}^{\phi}\Big{(}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})^{|u|}+2^{\frac{(1-\theta^{\prime})d}{\alpha+\lambda-d}m}D^{|u|}_{\alpha,\lambda,d,\theta^{\prime}}\Big{)},$
+[…]
 which is bounded by $\delta^{\prime}$ due to the definition of $\Gamma(m)$. Equation (29) with the above choice of $T_{u,m}$ then implies
@@
 $=$ $\sum_{\begin{subarray}{c}u\subseteq 1{:}s\\ u\neq\varnothing\end{subarray}}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{u}2^{\tilde{t}_{u}}\Big)^{\phi}C^{|u|}_{\alpha,\lambda,\theta}$,
-which is bounded as $s\to\infty$ because of equation (43). Our conclusion then follows from Lemma 2 with $\delta=1/16$ and equation (5). ∎
+which is bounded as $s\to\infty$ because of equation (31). Our conclusion then follows from Lemma 2 with $\delta=1/16$. ∎
 Remark 8.
 By Hölder’s inequality,
-$\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big{)}^{\phi}$
-$\leqslant$ $\Big{(}\sum_{j=1}^{s}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\Big{)}^{1-\frac{\phi}{1+2(\alpha+\lambda)}}\Big{(}\sum_{j=1}^{s}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}\Big{)}^{\frac{\phi}{1+2(\alpha+\lambda)}}.$
+$\sum_{j=1}^{\infty}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}}\Big)^{\phi}$
+$\leqslant$ $\Big(\sum_{j=1}^{\infty}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}\Big)^{1-\frac{\phi}{1+2(\alpha+\lambda)}}\Big(\sum_{j=1}^{\infty}\gamma^{\frac{1+2d}{\alpha+\lambda+1/2}}_{j}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}\Big)^{\frac{\phi}{1+2(\alpha+\lambda)}}.$
 Therefore, equation (31) holds under the assumptions in Remark 6.
 Remark 9.
-For the random linear scrambling using base-2 Niederreiter sequences, substituting the bound (42) into equation (43) yields the simplified tractability condition:
-$\lim_{s\to\infty}\sum_{j=1}^{s}\Big{(}\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}j\log_{2}(j)\Big{)}^{\phi}<\infty.$
-Since $\phi>1$, this condition is strictly weaker than the requirements in [9, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
-In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, $\mu^{(r)}_{\infty}$ attains dimension-independent convergence if
+For base-2 Niederreiter sequences, substituting the bound (30) into equation (31) yields the simplified tractability condition:
+$\sum_{j=1}^{\infty}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}j\log_{2}(j)\Big)^{\phi}<\infty.$
+Since $\phi>1$, this condition is strictly weaker than the requirements in [Goda2024, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates under the random linear scrambling.
+In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent convergence if
 $\phi(\frac{p}{\alpha+\lambda+1/2}-q)=\frac{p-(\alpha+\lambda+1/2)q}{\alpha+\lambda-d+1/2}>1.$ (32)
@@
 6 Numerical experiments
-In this section, we test our theory across integrands with varying dimensions and decay rates on $\gamma_{u}$. The test integrands have the common form
+This section tests our theoretical results on integrands with varying dimensions and decay rates of $\gamma_{u}$.
+We also reference [pan2024automatic], which provides further experiments demonstrating near-optimal convergence rates of median RQMC and comparisons to higher-order scrambled digital nets.
+The test integrands have the common form
 $f_{s,\gamma,\alpha}(\boldsymbol{x})=\prod_{j=1}^{s}\Big(1+\frac{c_{\alpha}}{j^{\gamma}}(x_{j}\exp(x_{j})-1)\Big),$
@@
 For each integrand $f_{s,\gamma,\alpha}$, we evaluate the performance of the median estimator $\hat{\mu}^{(r)}_{E}$ for $m=1,2,\dots,16$ under two randomization schemes: the completely random design (CRD) and the random linear scrambling (RLS).
-The estimator $\hat{\mu}^{(r)}_{E}$ is computed as the sample median of $12r-$ 1 independent realizations of $\hat{\mu}_{E}$, where $\hat{\mu}_{E}$ is defined as in equation (2) with precision $E=64$.
+The estimator $\hat{\mu}^{(r)}_{E}$ is computed as the sample median of 2r12r-1 independent realizations of $\hat{\mu}_{E}$, where $\hat{\mu}_{E}$ is defined as in equation (2) with precision $E=64$.
 The observed integration errors in our tested range of $m$ is much larger than $2^{-E}$, ensuring that $\hat{\mu}^{(r)}_{E}$ closely approximates $\hat{\mu}^{(r)}_{\infty}$.
-Consequently, the error bounds in Theorem 1, 2 and 3 apply to $\hat{\mu}^{(r)}_{E}$ with probability at least $1-2^{-r}$.
+Consequently, the error bounds in Theorem 2, 3 and 4 apply to $\hat{\mu}^{(r)}_{E}$ with probability at least $1-2^{-r}$.
 We set $r=10$, limiting the failure probability per data point to below $0.1\%$.
-As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the $12r-$ 1 $\hat{\mu}_{E}$ realizations used by RLS.
-The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [12].
+As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the 2r12r-1 $\hat{\mu}_{E}$ realizations used by RLS.
+The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [joe:kuo:2008].
 Refer to caption Refer to caption Refer to caption
@@
 We also note that STD exhibits convergence rates comparable to those of RLS and substantially surpasses the theoretical $O(2^{-(3/2)m})$ rate derived from the root mean squared error when $\gamma=3$ or $4$.
-This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 2 and 3.
+This discrepancy arises because each estimator $\hat{\mu}_{\infty}$ in STD is a special case of $\hat{\mu}^{(r)}_{\infty}$ with $r=1$, which satisfies the rapidly converging probabilistic error bounds in Theorems 3 and 4.
 However, averaging multiple replicates in STD compromises these bounds, as even a single outlier among the replicates can dominate the error and inflate its magnitude.
@@
 By developing a theory for non-smooth integrands, our results could be extended to real-world applications in finance and related fields.
-For instance, [10] showcases an infinite-dimensional integration problem where all ANOVA components are smooth despite discontinuities in the integrand’s derivatives, suggesting that non-smoothness may not fundamentally compromise convergence rates in high-dimensional settings.
+For instance, [griebel2017anova] showcases an infinite-dimensional integration problem where all ANOVA components are smooth despite discontinuities in the integrand’s derivatives, suggesting that non-smoothness may not fundamentally compromise convergence rates in high-dimensional settings.
+Appendix A: Proof of Lemma 4
+Let $\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)$.
+Because $|\kappa^{\prime}_{j}|\geqslant|\kappa_{j}|=\alpha+1$ for $j\in v$ and $|\kappa^{\prime}_{j}|=|\kappa_{j}|=\alpha_{j}$ for $j\in u\setminus v$, we can apply Lemma 3 to get
+$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=$
+$\Big|\int_{[0,1]^{|u|}}\Big(f_{u}^{(\boldsymbol{\alpha}_{u})}(\boldsymbol{x})\prod_{j\in u\setminus v}W_{\kappa_{j}}(x_{j})\Big)\Big(\prod_{j\in v}{\mathrm{wal}_{\kappa^{\prime}_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}\alpha}}}(x_{j})W_{\lceil\kappa_{j}\rceil_{1{:}\alpha}}(x_{j})\Big)\,\mathrm{d}\boldsymbol{x}_{u}\Big|^{2},$
+Let $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$, $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ and $\kappa^{-}_{j}=\kappa^{\prime}_{j}\setminus\kappa_{j}$ for $j\in v$.
+We can rewrite the above equation as
+$|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}=\Bigl|\int_{[0,1]^{|v|}}g(\boldsymbol{x}_{v})\prod_{j\in v}{\mathrm{wal}_{\kappa^{-}_{j}}}(x_{j})(-1)^{\vec{x}_{j}(\ell_{j})}W_{\kappa^{+}_{j}}(x_{j})\,\mathrm{d}\boldsymbol{x}_{v}\Bigr|^{2},$
+where
+$g(\boldsymbol{x}_{v})=I_{v}\Big(f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}W_{\kappa_{j}}\Big).$
+In the special case v=1:sv=1{:}s, [pan2024automatic, page 14 line 14] shows
+$\sum_{\kappa^{-}_{1},\dots,\kappa^{-}_{s}}\Bigl|\int_{[0,1]^{s}}g(\boldsymbol{x})\prod_{j\in 1{:}s}{\mathrm{wal}_{\kappa^{-}_{j}}}(x_{j})(-1)^{\vec{x}_{j}(\ell_{j})}W_{\kappa^{+}_{j}}(x_{j})\,\mathrm{d}\boldsymbol{x}\Bigr|^{2}\\ \leqslant 2^{-s}\Bigl(V^{(s)}_{\lambda}(g)\Bigr)^{2}\Bigl(\prod_{j=1}^{s}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}\Bigr),$
+where the summation $\sum_{\kappa^{-}_{1},\dots,\kappa^{-}_{s}}$ means each $\kappa^{-}_{j}$ runs over all subsets of $\mathbb{N}$ satisfying $\lceil\kappa^{-}_{j}\rceil_{1}<\ell_{j}$.
+After mapping $v$ to 1:|v|1{:}|v| and compactly rewriting the summation as $\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}$, we obtain
+$\sum_{\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)}|\hat{f}_{u}(\boldsymbol{k}^{\prime})|^{2}\leqslant 2^{-|v|}\Big(V^{(|v|)}_{\lambda}(g)\Big)^{2}\Bigl(\prod_{j\in v}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}\Bigr).$ (34)
+Letting $\rho_{j}(x_{j})=W_{\kappa_{j}}(x_{j})/\|W_{\kappa_{j}}\|_{\infty}$ for $j\in u\setminus v$, equation (6) implies
+$V^{(|v|)}_{\lambda}(g)=$ $\Big(\prod_{j\in u\setminus v}2\prod_{\ell^{\prime}\in\kappa_{j}}2^{-\ell^{\prime}-1}\Big)V^{(|v|)}_{\lambda}\Big(I_{u}(f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}\rho_{j})\Big)$
+$\leqslant$ $2^{|u|-|v|}\Big(\prod_{j\in u\setminus v}\prod_{\ell^{\prime}\in\kappa_{j}}2^{-\ell^{\prime}-1}\Big)V^{v}_{\lambda}(f^{(\boldsymbol{\alpha}_{u})}_{u})$
+because $\prod_{j\in u\setminus v}\rho_{j}\in D_{v^{c}}$ from equation (11). Our conclusion follows once we put the above bound into equation (34) and plug in
+$\prod_{j\in v}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}=4^{|v|}\prod_{j\in v}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
+and $\lceil\kappa_{j}\rceil_{\alpha+1}=0$ for $j\in u\setminus v$.
+Appendix B: Proof of Lemma 5
+By equation (8), $\hat{f}_{u}(\boldsymbol{k})=0$ if $\boldsymbol{s}(\boldsymbol{k})\neq u$. Letting
+$\mathbb{N}^{u}_{\alpha}=\{(k_{j},j\in u)\in\mathbb{N}^{|u|}\mid|\kappa_{j}|\leqslant\alpha+1\text{ for }j\in u\},$
+we can apply Lemma 4 and equation (15) to get
+$\sum_{\boldsymbol{k}\in\mathbb{N}_{*}^{s}\setminus K_{u}(T)}|\hat{f}_{u}(\boldsymbol{k})|^{2}$ (35)
+$\leqslant$ $4^{|u|}\|f\|^{2}_{u,\alpha,\lambda}\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\boldsymbol{1}\Big\{\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)\Big\}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}.$
+Because $\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)$ implies
+$(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|>T,$
+we have for any $\theta\in(0,1)$ and $\rho_{\theta}=4^{-1+\theta}$
+$\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\boldsymbol{1}\Big\{\boldsymbol{k}_{u}\notin\bigcup_{v\subseteq u}\mathcal{K}_{u,v}(T)\Big\}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
+$\leqslant$ $\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}4^{-\theta T+\theta(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\theta\sum_{j\in u}\|\kappa_{j}\|}\prod_{j\in u}4^{(1-\lambda)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-\ell^{\prime}-1}$
+$=$ $4^{-\theta T}\sum_{\boldsymbol{k}_{u}\in\mathbb{N}^{u}_{\alpha}}\prod_{j\in u}4^{-(1-\theta)(\lambda-1)\lceil\kappa_{j}\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa_{j}}4^{-(1-\theta)\ell^{\prime}-1}$
+$=$ $4^{-\theta T}\Big(\sum_{\alpha^{\prime}=0}^{\alpha}4^{-\alpha^{\prime}}\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\quad+\quad 4^{-\alpha-1}\sum_{\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})}\rho_{\theta}^{(\lambda-1)\lceil\kappa\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}\Big)^{|u|},$
+where $\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})=\{\kappa\subseteq\mathbb{N}\mid|\kappa|=\alpha^{\prime}\}$. To bound the above sum, consider the generating function
+$g(t)=\Big(\sum_{N=1}^{\infty}t^{N}\Big)^{\alpha^{\prime}},$
+and let $t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [flaj:sedg:2009] for background knowledge on generating functions).
+It follows that $t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$.
+Because each $\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})$ is repeated $(\alpha^{\prime})!$ times in the above counting, we have
+$\sum_{\kappa\in\mathcal{S}_{\alpha^{\prime}}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa}t^{\ell^{\prime}}\leqslant\frac{1}{(\alpha^{\prime})!}\Big(\sum_{N=1}^{\infty}t^{N}\Big)^{\alpha^{\prime}}=\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}.$ (36)
+with $t=\rho_{\theta}$.
+For $\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})$, we let $\kappa^{+}=\lceil\kappa\rceil_{1{:}\alpha}$ and $\ell=\lceil\kappa\rceil_{\alpha+1}$.
+By the one-to-one correspondence between $\kappa^{+}$ and $\kappa^{+}-\ell=\{\ell^{\prime}-\ell\mid\ell^{\prime}\in\kappa\}\in\mathcal{S}_{\alpha}(\mathbb{N})$,
+$\sum_{\kappa\in\mathcal{S}_{\alpha+1}(\mathbb{N})}\rho_{\theta}^{(\lambda-1)\lceil\kappa\rceil_{\alpha+1}}\prod_{\ell^{\prime}\in\kappa}\rho_{\theta}^{\ell^{\prime}}=$ $\sum_{\ell=1}^{\infty}\rho_{\theta}^{(\alpha+\lambda)\ell}\sum_{\kappa^{+}-\ell\in\mathcal{S}_{\alpha}(\mathbb{N})}\prod_{\ell^{\prime}\in\kappa^{+}-\ell}\rho_{\theta}^{\ell^{\prime}}$
+$=$ $\sum_{\ell=1}^{\infty}\rho_{\theta}^{(\alpha+\lambda)\ell}\frac{1}{\alpha!(\rho_{\theta}^{-1}-1)^{\alpha}}$
+$=$ $\frac{1}{\alpha!(\rho^{-\alpha-\lambda}_{\theta}-1)(\rho_{\theta}^{-1}-1)^{\alpha}}.$
+Our conclusion follows once we put the above bounds into equation (35).
+Appendix C: Proof of Lemma 6
+The conclusion is trivial when $T<0$, so we assume $T\geqslant 0$.
+As in the proof Lemma 4, for each $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ we write $\kappa^{+}_{j}=\lceil\kappa_{j}\rceil_{1{:}\alpha}$ and $\ell_{j}=\lceil\kappa_{j}\rceil_{\alpha+1}$ for $j\in v$.
+By equation (14), each $\boldsymbol{k}^{\prime}=(k^{\prime}_{1},\dots,k^{\prime}_{s})\in B(\boldsymbol{k}_{u},v)$ must have the same $\kappa^{\prime}_{j}$ for $j\in 1{:}s\setminus v$.
+For $j\in v$, however, $\kappa^{\prime}_{j}\setminus\kappa_{j}$ can be any subset of $1{:}(\ell_{j}-1)$ if $\ell_{j}\geqslant 2$ and $\kappa^{\prime}_{j}=\kappa_{j}$ if $\ell_{j}=1$.
+Thus, $|B(\boldsymbol{k}_{u},v)|=\prod_{j\in v}2^{\ell_{j}-1}$.
+Because $B(\boldsymbol{k}_{u},v)$ are mutually disjoint for distinct $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$,
+$|K_{u}(T)|=\sum_{v\subseteq u}\sum_{\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)}\prod_{j\in v}2^{\ell_{j}-1}.$ (37)
+Next, we notice
+$(\lambda-1)\sum_{j\in u}\lceil\kappa_{j}\rceil_{\alpha+1}+\sum_{j\in u}\|\kappa_{j}\|=(\alpha+\lambda)\sum_{j\in v}\ell_{j}+\sum_{j\in v}\sum_{\ell^{\prime}\in\kappa^{+}_{j}}(\ell^{\prime}-\ell_{j})+\sum_{j\in u\setminus v}\|\kappa_{j}\|$
+and there is a one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}=\{\ell^{\prime}-\ell_{j}\mid\ell^{\prime}\in\kappa^{+}_{j}\}$ for $j\in v$.
+Furthermore, $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$ implies
+$\sum_{j\in v}\sum_{\ell^{\prime}\in\kappa^{+}_{j}}(\ell^{\prime}-\ell_{j})+\sum_{j\in u\setminus v}\|\kappa_{j}\|=\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|+\sum_{j\in u\setminus v}\|\kappa_{j}\|\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}.$
+For $\boldsymbol{\alpha}_{u}=(\alpha_{j},j\in u)\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$, let
+$\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})=\Big\{(k_{j},j\in u)\in\mathbb{N}_{0}^{|u|}\Big||\kappa_{j}|=\alpha_{j}\ \forall j\in u,\sum_{j\in u}\|\kappa_{j}\|=N\Big\}.$
+We can decompose $\mathcal{K}_{u,v}(T)$ into subsets with the same $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ and $(\kappa^{+}_{j}-\ell_{j},j\in v;\kappa_{j},j\in u\setminus v)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ for $N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}$, and rewrite equation (37) as
+$|K_{u}(T)|=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$ (38)
+$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|+$
+$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}\hskip-10.00002pt2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
+Let $\overline{N}=\lfloor(T-N)/(\alpha+\lambda)\rfloor$.
+For nonempty $v\subseteq u$ and integer $n\geqslant|v|$, there are $n-1\choose|v|-1$} number of $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfying $\|\boldsymbol{\ell}_{v}\|_{1}=n$.
+Hence, when $\overline{N}\geqslant|v|$,
+$\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}\hskip-10.00002pt2^{\|\boldsymbol{\ell}_{v}\|_{1}}=$ $\sum_{n=|v|}^{\overline{N}}{n-1\choose|v|-1}2^{n}\leqslant\sum_{n=|v|}^{\overline{N}}{\overline{N}-1\choose|v|-1}2^{n}\leqslant{\overline{N}-1\choose|v|-1}2^{\overline{N}+1}.$ (39)
+When $\overline{N}<|v|$, no $\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}$ satisfies $\|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)$ and the above sum is trivially 0.
+Putting the above bound into equation (38), we get
+$|K_{u}(T)|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|$ (40)
+$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
+$\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T/(\alpha+\lambda)\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}.$
+To bound the sum over $N$, consider the generating function
+$g(t)=\prod_{j\in u}\Big(\sum_{N=1}^{\infty}t^{N}\Big)^{\alpha_{j}}.$
+Because each element of $\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})$ is counted $\prod_{j\in u}(\alpha_{j})!$ times by $t^{N}[g(t)]$ , an argument similar to equation (36) gives
+$\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|t^{N}\leqslant\prod_{j\in u}\frac{1}{(\alpha_{j})!}\Big(\sum_{N=1}^{\infty}t^{N}\Big)^{\alpha_{j}}=\prod_{j\in u}\frac{1}{(\alpha_{j})!(t^{-1}-1)^{\alpha_{j}}}$ (41)
+for $t=2^{-1/(\alpha+\lambda)}$ and
+$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}\leqslant$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\prod_{j\in u}\frac{1}{(\alpha_{j})!(2^{1/(\alpha+\lambda)}-1)^{\alpha_{j}}}$
+$=$ $A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|},$
+where $A_{\alpha,\lambda},B_{\alpha,\lambda}$ are defined in equation (19). By a similar calculation,
+$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|\leqslant$ $2^{T/(\alpha+\lambda)}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
+$\leqslant$ $2^{T/(\alpha+\lambda)}B^{|u|}_{\alpha,\lambda}$.
+Putting the above bounds into equation (40), we finally get
+[…]
+$\leqslant$ $2^{T/(\alpha+\lambda)}\Big(B^{|u|}_{\alpha,\lambda}+\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}\max\big(T/(\alpha+\lambda),1\big)^{|v|}A_{\alpha,\lambda}^{|v|}B_{\alpha,\lambda}^{|u|-|v|}\Big)$
+$=$ $2^{T/(\alpha+\lambda)}\Big(B_{\alpha,\lambda}+\max\big(T/(\alpha+\lambda),1\big)A_{\alpha,\lambda}\Big)^{|u|}.$
+Appendix D: Proof of Lemma 7
+We adopt the notation used in Appendix Appendix C: Proof of Lemma 6.
+For $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$, we again consider the one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}$ for $j\in v$.
+Because $d\in\mathbb{N}$ and $d<\alpha+\lambda$, we must have $d\leqslant\alpha$.
+By $|\kappa^{+}_{j}|=\alpha\geqslant d$ for $j\in v$,
+$\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|_{(d)}=\sum_{j\in v}\|\kappa^{+}_{j}\|_{(d)}-d\ell_{j}=\Big(\sum_{j\in v}\|\kappa_{j}\|_{(d)}\Big)-d\|\boldsymbol{\ell}_{v}\|_{1}$
+and
+$\sum_{j\in v}\|\kappa^{+}_{j}-\ell_{j}\|_{(d)}+\sum_{j\in u\setminus v}\|\kappa_{j}\|_{(d)}=\Big(\sum_{j\in u}\|\kappa_{j}\|_{(d)}\Big)-d\|\boldsymbol{\ell}_{v}\|_{1}.$
+Let
+$\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})=\Big\{(k_{j},j\in u)\in\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})\Big|\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N^{\prime}\Big\}.$
+Similar to equation (38), we can write
+$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|$ (43)
+$=$ $\sum_{v\subseteq u}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}-d\|\boldsymbol{\ell}_{v}\|_{1}<N^{\prime}\leqslant N\leqslant T-(\alpha+\lambda)\|\boldsymbol{\ell}_{v}\|_{1}\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\prod_{j\in v}2^{\ell_{j}-1}$
+$=$ $\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|+$
+$\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}.$
+The sum over $\boldsymbol{\ell}_{v}$ is empty if $(T^{\prime}-N^{\prime})/d\geqslant(T-N)/(\alpha+\lambda)$.
+Because $T^{\prime}/d\geqslant T/(\alpha+\lambda)$ and $N\geqslant N^{\prime}$, we further infer the sum is empty unless $N>N_{*}$ for
+$N_{*}=\frac{(\alpha+\lambda)T^{\prime}-dT}{\alpha+\lambda-d}.$
+Therefore,
+$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}>(T^{\prime}-N^{\prime})/d,\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
+$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}\sum_{N^{\prime}=0}^{N}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\sum_{\begin{subarray}{c}\boldsymbol{\ell}_{v}\in\mathbb{N}^{|v|}\\ \|\boldsymbol{\ell}_{v}\|_{1}\leqslant(T-N)/(\alpha+\lambda)\end{subarray}}2^{\|\boldsymbol{\ell}_{v}\|_{1}}$
+$\leqslant$ $\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor-1\choose|v|-1}2^{\lfloor(T-N)/(\alpha+\lambda)\rfloor+1},$
+where we have applied equation (39) in the last step. Letting $T_{*}=(T-N_{*})/(\alpha+\lambda)$, we put the above bound into equation (43) and get
+$|K_{u}(T)\cap K^{\prime}_{u}(T^{\prime})|-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (44)
+$\leqslant$ $\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0}\\ N_{*}<N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|{\lfloor(T-N)/(\alpha+\lambda)\rfloor\choose|v|-1}2^{(T-N)/(\alpha+\lambda)+1}$
+$\leqslant$ $2^{T_{*}}\sum_{\begin{subarray}{c}v\subseteq u\\ v\neq\varnothing\end{subarray}}2^{-|v|+1}{\lfloor T_{*}\rfloor\choose|v|-1}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}=\alpha\ \forall j\in v,\\ \alpha_{j}>0\ \forall j\in u\setminus v\end{subarray}}\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N,\boldsymbol{\alpha}_{u})|2^{-N/(\alpha+\lambda)}$
+$\leqslant$ $2^{T_{*}}\Big(B_{\alpha,\lambda}+\max(T_{*},1)A_{\alpha,\lambda}\Big)^{|u|},$
+where the last step follows from a calculation similar to equation (40) and (42).
+Next, we notice
+$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant t^{-T}\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}t^{N^{\prime}}\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$ (45)
+for $t=2^{-1/(\alpha+\lambda-d)}$.
+To bound the sum over $N$, we let $w=\{j\in u\mid\alpha_{j}>d\}$ for $\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}$ and define $\boldsymbol{\alpha}^{-}_{u}=(\min(\alpha_{j},d),j\in u)$ and $\boldsymbol{\alpha}^{+}_{u}=(\alpha_{j}-d,j\in w)$ if $w\neq\varnothing$.
+For $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$, we further let $\kappa^{-}_{j,d}=\lceil\kappa_{j}\rceil_{1{:}\min(\alpha_{j},d)}$ for $j\in u$ and $\kappa^{+}_{j,d}=\kappa_{j}\setminus\lceil\kappa_{j}\rceil_{1{:}d}$ for $j\in w$.
+Because for $\boldsymbol{k}_{u}\in\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})$,
+$\sum_{j\in w}\|\kappa^{+}_{j,d}\|=\sum_{j\in w}\|\kappa_{j}\|-\sum_{j\in w}\|\kappa^{-}_{j,d}\|=\sum_{j\in u}\|\kappa_{j}\|-\sum_{j\in u}\|\kappa_{j}\|_{(d)}=N-N^{\prime},$
+[…]
+$\sum_{N=N^{\prime}}^{\lfloor T\rfloor}t^{N-N^{\prime}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=N^{\prime}}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{w}(N-N^{\prime},\boldsymbol{\alpha}^{+}_{u})|t^{N-N^{\prime}}$
+$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\sum_{N=0}^{\infty}|\tilde{\mathcal{K}}_{w}(N,\boldsymbol{\alpha}^{+}_{u})|t^{N}$
+$\leqslant$ $|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}$,
+where we have applied equation (41) in the last step. Putting the above bound into equation (45), we get for any $\theta^{\prime}\in(0,1)$,
+$\sum_{\begin{subarray}{c}N\in\mathbb{N}_{0},N^{\prime}\in\mathbb{N}_{0}\\ T^{\prime}<N^{\prime}\leqslant N\leqslant T\end{subarray}}|\tilde{\mathcal{K}}_{u,d}(N,N^{\prime},\boldsymbol{\alpha}_{u})|$
+$\leqslant$ $t^{-T}\Big(\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big)\sum_{N^{\prime}=\lceil T^{\prime}\rceil}^{\lfloor T\rfloor}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{N^{\prime}}$
+$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big(\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big)\sum_{N^{\prime}=0}^{\infty}|\tilde{\mathcal{K}}_{u}(N^{\prime},\boldsymbol{\alpha}^{-}_{u})|t^{(1-\theta^{\prime})N^{\prime}}$
+$\leqslant$ $t^{-T+\theta^{\prime}T^{\prime}}\Big(\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big)\Big(\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big)$
+where we have applied equation (41) once more. Our conclusion finally follows from equation (44), $t=2^{-1/(\alpha+\lambda-d)}$ and
+$\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{u}\in\mathbb{Z}_{\leqslant\alpha}^{|u|}\\ \alpha_{j}>0\ \forall j\in u\end{subarray}}\Big(\prod_{j\in w}\frac{1}{(\alpha^{+}_{j})!(t^{-1}-1)^{\alpha^{+}_{j}}}\Big)\Big(\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big)$
+$=$ $\sum_{w\subseteq u}\Big(\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big)^{|w|}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{-}_{u}\in\mathbb{Z}^{|u|}_{\leqslant d}\\ \alpha_{j}=d\ \forall j\in w\\ \alpha^{-}_{j}>0\ \forall j\in u\setminus w\end{subarray}}\Big(\prod_{j\in u}\frac{1}{(\alpha^{-}_{j})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{-}_{j}}}\Big)^{|u|}$
+$=$ $\sum_{w\subseteq u}\Big(\frac{1}{d!(t^{-1+\theta^{\prime}}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(t^{-1}-1)^{\alpha^{\prime}}}\Big)^{|w|}\Big(\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(t^{-1+\theta^{\prime}}-1)^{\alpha^{\prime}}}\Big)^{|u|-|w|}$
+$=$ $D_{\alpha,\lambda,d,\theta^{\prime}}^{|u|},$
+where $D_{\alpha,\lambda,d,\theta^{\prime}}$ is defined in equation (26).
 Acknowledgments
@@
 For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
+The author thanks David Krieg and Art Owen for valuable discussions and two anonymous reviewers for many helpful comments.

### v4

@@
 This structured approach allows QMC to avoid the exponential cost growth with dimension of classical quadrature rules while achieving faster error convergence than MC.
-These advantages have made QMC particularly valuable in fields like computational finance [joy1996quasi, l2009quasi] and physics [kuo:nuye:2016], where high-dimensional integrals are pervasive.
+These advantages have made QMC particularly valuable in fields like computational finance [19, 25] and physics [23], where high-dimensional integrals are pervasive.
 Building on these strengths, we focus on randomized quasi-Monte Carlo (RQMC) methods based on linearly scrambled base-2 digital nets.
-Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [dick:pill:2010].
-The random linear scrambling [mato:1998:2], a computationally efficient alternative to Owen’s scrambling [rtms], preserves the structure of digital nets while introducing controlled randomness.
+Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [4].
+The random linear scrambling [28], a computationally efficient alternative to Owen’s scrambling [31], preserves the structure of digital nets while introducing controlled randomness.
 This randomization enhances the root mean squared error convergence rate to nearly $O(n^{-3/2})$ under smoothness conditions and enables statistical inference.
-However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [superpolyone].
-To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [chen2025randomintegrationalgorithmhighdimensional, goda2024simpleuniversalalgorithmhighdimensional, goda:lecu:2022, Goda2024, ye2025medianqmcmethodunbounded].
-The median of linearly scrambled digital net estimators was first analyzed in [superpolyone], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$.
-Subsequent work in [superpolymulti] extended this result to higher dimensions.
-For finitely differentiable integrands, [pan2024automatic] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
+However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [34].
+To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [3, 10, 11, 12, 44].
+The median of linearly scrambled digital net estimators was first analyzed in [34], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$.
+Subsequent work in [35] extended this result to higher dimensions.
+For finitely differentiable integrands, [36] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
 A critical gap in these prior studies is the role of dimensionality.
-We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [kuo:sloa:wasi:wozn:2010].
-Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [nova:wozn:2008], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3).
-This notion of relative variation aligns with the effective dimension framework introduced in [cafmowen], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions.
-The connection to weighted function spaces [sloa:wozn:1998] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
-Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [Goda2024].
+We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [24].
+Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [30], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3).
+This notion of relative variation aligns with the effective dimension framework introduced in [2], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions.
+The connection to weighted function spaces [38] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
+Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [12].
 A key distinction lies in the error criterion: their analysis employs the worst-case integration error over the unit ball in a weighted Sobolev space, whereas we demonstrate that median estimators achieve near-optimal convergence rates on a fixed integrand with high probability.
-Our probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [kunsch2019solvable, kunsch2019optimal].
+Our probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [21, 22].
 Within this framework, we prove, informally stated, the following result:
 Theorem 1.
-Given $\eta>0$, $k\in\mathbb{N}$ and $f\in C^{(k,\dots,k)}([0,1]^{s})$, it holds with probability $1-2^{-r}$ that the median estimator $\hat{\mu}^{(r)}_{\infty}$ from Subsection 2.3 satisfies
-$|\hat{\mu}^{(r)}_{\infty}-\mu|\leqslant C_{\eta,k,f}n^{-k-1/2+\eta}.$
+Given $\eta>0$, $k\in\mathbb{N}$, $r\in\mathbb{N}$ and $f\in C^{(k,\dots,k)}([0,1]^{s})$, it holds with probability $1-2^{-r}$ that the median estimator $\hat{\mu}^{(r)}_{\infty}$ from Subsection 2.3 satisfies
+$\left|\hat{\mu}^{(r)}_{\infty}-\int_{[0,1]^{s}}f(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}\right|\leqslant C_{\eta,k,f}n^{-k-1/2+\eta}.$
+Here, $C^{(k,\dots,k)}([0,1]^{s})$ denotes functions with continuous dominating mixed partial derivatives of order $k$ (see Subsection 2.4).
 The constant $C_{\eta,k,f}$ is independent of $s$ under the random linear scrambling if the relative variation $\gamma_{u}$ satisfies the tractability conditions in Theorem 5.
-These nearly $O(n^{-k-1/2})$ integrand-specific error rates complement the nearly $O(n^{-k})$ worst-case rates established in [Goda2024].
-Furthermore, Theorem 5 requires strictly weaker assumptions than those assumed in [Goda2024] for the random linear scrambling, demonstrating that dimension-independent convergence is more easily achieved under our error criterion.
-As a final remark, we note that while several algorithms match lower complexity bounds for integration in specific spaces—such as higher-order digital nets [goda2016explicit, goda2017optimal, goda2018optimal] and Frolov lattice rules [frolov1976upper, krieg2017universal, ullrich2016role]—our method offers distinct practical advantages.
+These nearly $O(n^{-k-1/2})$ integrand-specific error rates complement the nearly $O(n^{-k})$ worst-case rates established in [12].
+Furthermore, Theorem 5 requires strictly weaker assumptions than those assumed in [12] for the random linear scrambling, demonstrating that dimension-independent convergence is more easily achieved under our error criterion.
+As a final remark, we note that while several algorithms match lower complexity bounds for integration in specific spaces—such as higher-order digital nets [13, 14, 15], Frolov lattice rules [8, 20, 42] and Smolyak algorithms [9, 37, 39]—our method offers distinct practical advantages.
 It requires no smoothness parameter specification, yet automatically matches theoretical lower bounds up to an arbitrarily small $\eta>0$ in the convergence rate, making it both efficient and immune to parameter misspecification.
 Combined with its straightforward implementation and the dimension-independent convergence established in this paper, the median estimator presents a highly practical and robust alternative for applied problems.
-Furthermore, recent work has successfully extended median RQMC to $L^{2}$-approximation [pan2025universal, pan2025l_2], demonstrating its broader utility beyond numerical integration.
+Furthermore, recent work has successfully extended median RQMC to $L^{2}$-approximation [32, 33], demonstrating its broader utility beyond numerical integration.
 The remainder of this paper is structured as follows.
@@
 The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli(0.5)(0.5) entries.
-A common way to randomize $C_{j}$ is the random linear scrambling [mato:1998:2]:
+A common way to randomize $C_{j}$ is the random linear scrambling [28]:
 $C_{j}=M_{j}\mathcal{C}_{j}\ \ \mathsf{mod}\ 2,$
-where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli(0.5)(0.5)), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [dick:pill:2010, Chapter 4.4] for details).
+where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli(0.5)(0.5)), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [4, Chapter 4.4] for details).
 While effective, the reliance of random linear scrambling on carefully designed generating matrices complicates theoretical analysis.
 We address this in Section 5.
-For initial insights, Section 4 adopts the complete random design [pan2024automatic], where all entries of $C_{j}$ are drawn independently from Bernoulli(0.5)(0.5).
+For initial insights, Section 4 adopts the complete random design [36], where all entries of $C_{j}$ are drawn independently from Bernoulli(0.5)(0.5).
 This approach retains the asymptotic convergence rate of random linear scrambling while circumventing the need for pre-optimized generator matrices.
@@
 In practice, $E$ is constrained by computational precision.
-Lemma 1 of [superpolymulti] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$.
+Lemma 1 of [35] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$.
 Thus, $\hat{\mu}_{E}$ closely approximates $\hat{\mu}_{\infty}$ when $\omega_{f}(\sqrt{s}2^{-E})\ll|\hat{\mu}_{E}-\mu|$.
@@
 $\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})=\prod_{j=1}^{s}\mathrm{wal}_{k_{j}}(x_{j})=(-1)^{\sum_{j=1}^{s}\vec{k}_{j}^{\mathsf{T}}\vec{x}_{j}},$
-where $\boldsymbol{k}=(k_{1},\dots,k_{s})\in\mathbb{N}^{s}_{0}$.
-These functions constitute a complete orthonormal basis for $L^{2}([0,1]^{s})$ [dick:pill:2010], enabling the Walsh series decomposition:
+where $\boldsymbol{k}=(k_{1},\dots,k_{s})\in\mathbb{N}^{s}_{0}$. These functions constitute a complete orthonormal basis for $L^{2}([0,1]^{s})$ [4], enabling the Walsh series decomposition:
 $f(\boldsymbol{x})=\sum_{\boldsymbol{k}\in\mathbb{N}_{0}^{s}}\hat{f}(\boldsymbol{k})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x}),\quad\text{where}\quad\hat{f}(\boldsymbol{k})=\int_{[0,1]^{s}}f(\boldsymbol{x})\mathrm{wal}_{\boldsymbol{k}}(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}.$ (3)
-The equality holds in the $L^{2}$-sense. Using this decomposition, [pan2024automatic, Theorem 1] establishes a key error structure for QMC estimators:
+The equality holds in the $L^{2}$-sense. Using this decomposition, [36, Theorem 1] establishes a key error structure for QMC estimators:
 Lemma 1.
@@
 Proof.
-First note that the random signs $S(\boldsymbol{k})$ in equation (4) are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [superpolyone, Lemma 4] for the one-dimensional case).
+First note that the random signs $S(\boldsymbol{k})$ in equation (4) are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [34, Lemma 4] for the one-dimensional case).
 Since $\mathcal{A}$ is independent of $S(\boldsymbol{k})$, we have $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and
@@
 Proof.
-Equation (5) comes from [SUZUKI20161, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [SUZUKI20161, Section 3]. ∎
+Equation (5) comes from [41, Theorem 2.5], while properties of $W_{\kappa}(x)$ are proven in [41, Section 3]. ∎
 For $\alpha\in\mathbb{N}_{0}$, we say $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ if the mixed partial derivative $f^{(\boldsymbol{\alpha})}$ exists and is continuous on $0,1]^{s}$ for all $\boldsymbol{\alpha}\in\mathbb{Z}^{s}_{\leqslant\alpha}$.
 In this work, we assume $f\in C^{(\alpha,\dots,\alpha)}([0,1]^{s})$ in order to directly apply Lemma 3.
-This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (5), as discussed in [pan2024automatic, Remark 1].
+This regularity condition can be relaxed to the existence of weak derivatives $f^{(\boldsymbol{\alpha})}\in L^{1}([0,1]^{s})$ satisfying the integral identity in equation (5), as discussed in [36, Remark 1].
 3 ANOVA decomposition and effective dimensionality
-To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [meandim]:
+To characterize the effective dimensionality of integrands $f\in L^{2}([0,1]^{s})$, we employ the ANOVA decomposition [26]:
 $f=\sum_{u\subseteq 1{:}s}f_{u},$
@@
 if $j\in\boldsymbol{s}(\boldsymbol{k})$. Thus, $\hat{f}(\boldsymbol{k})=\hat{f}_{u}(\boldsymbol{k})$ if $\boldsymbol{s}(\boldsymbol{k})=u$.
-To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [dick:2008], which generalizes the usual Sobolev norms.
+To quantify the variation of ANOVA components $f_{u}$, we adopt the fractional Vitali variation with $\mathfrak{p}=2$ from [5], which generalizes the usual Sobolev norms.
 For a function $f:[0,1]^{d}\to\mathbb{R}$ and a subcube $J=\prod_{j=1}^{d}[a_{j},b_{j})$, we define the alternating sum over $J$:
@@
 $V^{(d)}_{1}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{2}},$ (10)
-with equality attained when $f\in C^{(1,\dots,1)}([0,1]^{s})$ [dick:2008, equation (3.4)].
-More generally, [LIU2025101935] shows $V^{(d)}_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$.
-Discontinuous $f$ can also have finite $V^{(d)}_{\lambda}(f)$ (see [pan2024automatic, Section 4.1] for an example).
+with equality attained when $f\in C^{(1,\dots,1)}([0,1]^{s})$ [5, equation (3.4)].
+More generally, [27] shows $V^{(d)}_{\lambda}(f)\leqslant\|f^{(1,\dots,1)}\|_{L^{p}}$ for $p=2/(3-2\lambda),\lambda\in[1/2,1]$.
+Discontinuous $f$ can also have finite $V^{(d)}_{\lambda}(f)$ (see [36, Section 4.1] for an example).
 Next, for $v\subseteq 1{:}d$, we define the partial integral operator $I_{v}:L^{1}([0,1]^{d})\to L^{1}([0,1]^{|v|})$ as:
@@
 $V^{v}_{\lambda}(f)=\sup_{\rho_{v^{c}}\in D_{v^{c}}}V^{(|v|)}_{\lambda}\Big(I_{v}(f\rho_{v^{c}})\Big).$ (12)
-By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [pan2024automatic, Section 3.2],
+By convention, $V^{1{:}d}_{\lambda}(f)=V^{(d)}_{\lambda}(f)$. From [36, Section 3.2],
 $\Bigl(V^{v}_{\lambda}(f)\Bigl)^{2}\leqslant\int_{[0,1]^{d-|v|}}\Bigl(V^{(|v|)}_{\lambda}\Bigl(f(\text{ }\cdot\text{ };\boldsymbol{x}_{v^{c}})\Bigr)\Bigr)^{2}\,\mathrm{d}\boldsymbol{x}_{v^{c}},$
@@
 $A_{\alpha,\lambda}=\frac{1}{\alpha!(2^{1/(\alpha+\lambda)}-1)^{\alpha}}\quad\text{and}\quad B_{\alpha,\lambda}=\sum_{\alpha^{\prime}=1}^{\alpha}\frac{1}{(\alpha^{\prime})!(2^{1/(\alpha+\lambda)}-1)^{\alpha^{\prime}}},$ (19)
-with the sum from 11 to $\alpha$ interpreted as 0 when $\alpha=0$.
+with the sum from $1$ to $\alpha$ interpreted as 0 when $\alpha=0$.
 Proof.
@@
 where $C_{\eta}$ is a constant depending on $\alpha,\lambda,\gamma_{u}$ and $\eta$.
-This convergence rate is nearly optimal, as it aligns with the lower bound established in [bakh:1959] for one-dimensional $f$ with $\lambda$-Hölder continuous $f^{(\alpha)}$, a subclass of functions with finite $\|f\|_{\alpha,\lambda}$.
+This convergence rate is nearly optimal, as it aligns with the lower bound established in [1] for one-dimensional $f$ with $\lambda$-Hölder continuous $f^{(\alpha)}$, a subclass of functions with finite $\|f\|_{\alpha,\lambda}$.
 Remark 3.
@@
 $\leqslant$ $\prod_{j=1}^{s}\Big(1+\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}(A_{\alpha,\lambda}m+B_{\alpha,\lambda})\Big),$
-which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [HICKERNELL2003286, Lemma 3].
+which can be further bounded by $C^{\prime}_{\eta}2^{\eta m}$ for any $\eta>0$ with a constant $C^{\prime}_{\eta}$ independent of $s$ [17, Lemma 3].
 Putting the above bounds into equation (22), we conclude that equation (24) holds with a constant $C_{\eta}$ independent of $s$.
-Notably, equation (25) resembles the tractability conditions assumed by [Goda2024, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
+Notably, equation (25) resembles the tractability conditions assumed by [12, Corollary 3.13] for polynomial lattices to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates.
 5 Tractability under more general randomization
 While the completely random design offers theoretical advantages in near-optimal convergence rates and tractability, the random linear scrambling often outperforms it in practical settings due to tighter error bounds on low-dimensional ANOVA components.
-For example, [joe:kuo:2008] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets.
+For example, [18] designs the generating matrices $\mathcal{C}_{j}$ to explicitly optimize the discrepancy of all two-dimensional projections of the resulting digital nets.
 In this section, we extend the analysis of Section 4 to a broader class of randomization schemes, with the random linear scrambling as a key example.
@@
 $\Pr(Z(\boldsymbol{k})=1)\leqslant 2^{-m+\tilde{t}_{u}}\boldsymbol{1}\Big\{\sum_{j\in u}\|\kappa_{j}\|_{(d)}>d(m-\tilde{t}_{u})\Big\}.$
-From [Goda2024, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [Goda2024, Section 2].
-We frame this assumption to align with broader applications in higher-order digital nets [dick:2011], where the parameter $d$ (the order of digital nets) may exceed $1$.
+From [12, Lemma 2.17], the random linear scrambling satisfies the above assumption with parameters $d=1$ and $\tilde{t}_{u}=t_{u}+|u|-1$, where $t_{u}$ denotes the projection-dependent $t$-value defined in [12, Section 2].
+We frame this assumption to align with broader applications in higher-order digital nets [6], where the parameter $d$ (the order of digital nets) may exceed $1$.
 This formulation ensures compatibility with both classical $d=1$ constructions and advanced higher-order constructions.
@@
 $D_{\alpha,\lambda,d,\theta^{\prime}}=\frac{1}{d!(2^{(1-\theta^{\prime})/\beta}-1)^{d}}\sum_{\alpha^{\prime}=1}^{\alpha-d}\frac{1}{(\alpha^{\prime})!(2^{1/\beta}-1)^{\alpha^{\prime}}}+\sum_{\alpha^{\prime}=1}^{d}\frac{1}{(\alpha^{\prime})!(2^{(1-\theta^{\prime})/\beta}-1)^{\alpha^{\prime}}},$ (26)
-with the sum from 11 to $\alpha-d$ interpreted as 0 when $\alpha=d$.
+with the sum from $1$ to $\alpha-d$ interpreted as 0 when $\alpha=d$.
 Proof.
@@
 Therefore, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent error bounds if $\Gamma_{\alpha,\lambda,d,\gamma_{u},\tilde{t}_{u},\theta}$ remains bounded as $s\to\infty$, which depends on growth of $\tilde{t}_{u}$.
-Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [c812a4a5-75ac-3614-936d-5782e44941d0] that the base-2 Niederreiter sequence [NIEDERREITER198851] satisfies
+Recall that $\tilde{t}_{u}=t_{u}+|u|-1$. It is shown in [43] that the base-2 Niederreiter sequence [29] satisfies
 $t_{u}\leqslant\sum_{j\in u}\Big(\log_{2}(j)+\log_{2}\log_{2}(j+2)+2\Big),$ (30)
-while the Sobol’ sequence [sobol67] satisfies
+while the Sobol’ sequence [40] satisfies
 $t_{u}\leqslant\sum_{j\in u}\Big(\log_{2}(j)+\log_{2}\log_{2}(j+1)+\log_{2}\log_{2}\log_{2}(j+3)+C\Big)$
@@
 $\sum_{j=1}^{\infty}\gamma_{j}^{2}2^{\tilde{t}_{j}(2(\alpha+\lambda)+1)}<\infty.$
-While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [dick:2011], and unbounded increases in $d$ ultimately violate the tractability condition above.
+While increasing $d$ may appear to enhance tractability, parameters $\tilde{t}_{u}$ and $\tilde{t}_{j}$ typically scale with $d$, as in the interlacing construction [6], and unbounded increases in $d$ ultimately violate the tractability condition above.
 Thus, prudent selection of $d$ is critical for practical implementations.
@@
 $\sum_{j=1}^{\infty}\Big(\gamma^{\frac{1}{\alpha+\lambda+1/2}}_{j}j\log_{2}(j)\Big)^{\phi}<\infty.$
-Since $\phi>1$, this condition is strictly weaker than the requirements in [Goda2024, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates under the random linear scrambling.
+Since $\phi>1$, this condition is strictly weaker than the requirements in [12, Corollary 3.11] for such digital nets to achieve near $O(2^{-(\alpha+\lambda+1/2)m})$ convergence rates under the random linear scrambling.
 In general, when $\gamma_{j}=O(j^{-p})$ and $2^{\tilde{t}_{j}}=O(j^{q})$ for $p\geqslant 0$ and $q\geqslant 0$, $\hat{\mu}^{(r)}_{\infty}$ attains dimension-independent convergence if
@@
 This section tests our theoretical results on integrands with varying dimensions and decay rates of $\gamma_{u}$.
-We also reference [pan2024automatic], which provides further experiments demonstrating near-optimal convergence rates of median RQMC and comparisons to higher-order scrambled digital nets.
+We also reference [36], which provides further experiments demonstrating near-optimal convergence rates of median RQMC and comparisons to higher-order scrambled digital nets.
 The test integrands have the common form
@@
 As a benchmark, we compute the standard RQMC estimator (STD), defined as the average of the 2r12r-1 $\hat{\mu}_{E}$ realizations used by RLS.
-The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [joe:kuo:2008].
+The generating matrices $\mathcal{C}_{j}$ used by RLS and STD come from [18].
 Refer to caption Refer to caption Refer to caption
@@
 By developing a theory for non-smooth integrands, our results could be extended to real-world applications in finance and related fields.
-For instance, [griebel2017anova] showcases an infinite-dimensional integration problem where all ANOVA components are smooth despite discontinuities in the integrand’s derivatives, suggesting that non-smoothness may not fundamentally compromise convergence rates in high-dimensional settings.
-Appendix A: Proof of Lemma 4
+For instance, [16] showcases an infinite-dimensional integration problem where all ANOVA components are smooth despite discontinuities in the integrand’s derivatives, suggesting that non-smoothness may not fundamentally compromise convergence rates in high-dimensional settings.
+Appendix A Proof of Lemma 4
 Let $\boldsymbol{k}^{\prime}\in B(\boldsymbol{k}_{u},v)$.
@@
 $g(\boldsymbol{x}_{v})=I_{v}\Big(f_{u}^{(\boldsymbol{\alpha}_{u})}\prod_{j\in u\setminus v}W_{\kappa_{j}}\Big).$
-In the special case v=1:sv=1{:}s, [pan2024automatic, page 14 line 14] shows
+In the special case v=1:sv=1{:}s, [36, page 14 line 14] shows
 $\sum_{\kappa^{-}_{1},\dots,\kappa^{-}_{s}}\Bigl|\int_{[0,1]^{s}}g(\boldsymbol{x})\prod_{j\in 1{:}s}{\mathrm{wal}_{\kappa^{-}_{j}}}(x_{j})(-1)^{\vec{x}_{j}(\ell_{j})}W_{\kappa^{+}_{j}}(x_{j})\,\mathrm{d}\boldsymbol{x}\Bigr|^{2}\\ \leqslant 2^{-s}\Bigl(V^{(s)}_{\lambda}(g)\Bigr)^{2}\Bigl(\prod_{j=1}^{s}4^{-\lambda\ell_{j}}\prod_{\ell^{\prime}\in\kappa^{+}_{j}}4^{-\ell^{\prime}-1}\Bigr),$
@@
 and $\lceil\kappa_{j}\rceil_{\alpha+1}=0$ for $j\in u\setminus v$.
-Appendix B: Proof of Lemma 5
+Appendix B Proof of Lemma 5
 By equation (8), $\hat{f}_{u}(\boldsymbol{k})=0$ if $\boldsymbol{s}(\boldsymbol{k})\neq u$. Letting
@@
 $g(t)=\Big(\sum_{N=1}^{\infty}t^{N}\Big)^{\alpha^{\prime}},$
-and let $t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [flaj:sedg:2009] for background knowledge on generating functions).
+and let $t^{N}]g(t)$ denote the coefficient of $t^{N}$ in $g(t)$ (see [7] for background knowledge on generating functions).
 It follows that $t^{N}]g(t)$ gives the number of ways to choose an ordered array of length $\alpha^{\prime}$ from $\mathbb{N}$ while ensuring the elements in the array sum to $N$.
@@
 Our conclusion follows once we put the above bounds into equation (35).
-Appendix C: Proof of Lemma 6
+Appendix C Proof of Lemma 6
 The conclusion is trivial when $T<0$, so we assume $T\geqslant 0$.
@@
 $=$ $2^{T/(\alpha+\lambda)}\Big(B_{\alpha,\lambda}+\max\big(T/(\alpha+\lambda),1\big)A_{\alpha,\lambda}\Big)^{|u|}.$
-Appendix D: Proof of Lemma 7
-We adopt the notation used in Appendix Appendix C: Proof of Lemma 6.
+Appendix D Proof of Lemma 7
+We adopt the notation used in Appendix C.
 For $\boldsymbol{k}_{u}\in\mathcal{K}_{u,v}(T)$, we again consider the one-to-one correspondence between $\kappa^{+}_{j}$ and $\kappa^{+}_{j}-\ell_{j}$ for $j\in v$.